### v1

# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

Abstract
As the potential for autonomous vehicles to be integrated on a large scale into modern traffic systems continues to grow, ensuring safe navigation in dynamic environments is crucial for smooth integration. To guarantee safety and prevent collisions, autonomous vehicles must be capable of accurately predicting the trajectories of surrounding traffic agents. Over the past decade, significant efforts from both academia and industry have been dedicated to designing solutions for precise trajectory forecasting. These efforts have produced a diverse range of approaches, raising questions about the differences between these methods and whether trajectory prediction challenges have been fully addressed. This paper reviews a substantial portion of recent trajectory prediction methods and devises a taxonomy to classify existing solutions. A general overview of the prediction pipeline is also provided, covering input and output modalities, modeling features, and prediction paradigms discussed in the literature. In addition, the paper discusses active research areas within trajectory prediction, addresses the posed research questions, and highlights the remaining research gaps and challenges.
keywords:
Autonomous Driving, Survey, Trajectory Prediction, Trajectory Forecasting
journal: Information Fusion
[…]
1 Introduction
The era of autonomous driving is on the verge of revolutionizing the modern transportation system. Automotive companies like Tesla, Waymo and Cruise are pioneers in today’s self-driving vehicle technology. The practical, innovative solutions designed by these companies, in conjunction with academic research, have led to rapid development in perception, prediction, planning, and control algorithms for autonomous driving. Predicting the future trajectories of traffic participants is crucial for ensuring safety in autonomous driving. The predicted trajectories are passed as input to planning algorithms to exclude paths leading to potential collisions and to select a safe course of action. Due to the diversity of traffic scenarios, the complex interactions between traffic participants, various maneuvers, and computational time requirements, the trajectory prediction problem poses a range of challenges (please refer to Figure 1 for factors impacting trajectory forecasting).
To accurately predict trajectories, the prediction model must account for the motion of vehicles, cyclists, pedestrians — all traffic participants. For vehicle motion, autonomous vehicles (AV) should be able to track surrounding vehicles and their potential maneuvers. As human drivers have an intuition of inter-dependency of surrounding vehicles, e.g., it can be anticipated if a neighboring car is cut off, there is a high chance that this vehicle will be compelled to perform a risky maneuver and potentially interfere with the ego-vehicle’s lane to avoid collision or suddenly hit the breaks. To equip autonomous vehicles with interaction-aware intuition, the prediction module needs to build a representation of each target agent’s surroundings and accounting for inter-dependency of traffic participants’ decisions. Regarding the human motion, beyond the complexity of interaction and inter-dependency, predicting human trajectories involves accounting for complex human behavior. Pedestrians’ motion is influenced not only by road topology and actions of surrounding agents but also by social rules, norms, and relations between agents. To accurately predict pedestrian trajectories, many factors need to be inferred from contextual and general knowledge information, e.g., a child and an adult exhibit different motion patterns. Additionally, prediction models that rely on traffic rules benefit from a reduced set of possible future trajectories. However, human drivers often violate traffic rules by over-speeding, changing lanes through a solid line, turning from a non-designated lane, ignoring stop signs, or even passing through a red light. Prediction models need to account for the risk of such violations to ensure robust and reliable predictions in real-world scenarios. Lastly, the trajectory prediction task has a multimodal nature, i.e., the same past trajectory can unfold into several possible future trajectories, e.g., a vehicle may slow down for a traffic signal or speed through if it turns green. To account for multimodality, the output of the prediction model must contain the distribution of possible trajectories with probabilities.
Figure 1: Factors impacting Trajectory Prediction: 1) cars waiting at a red light, demonstrating adherence to traffic rules, 2) a pedestrian waiting to cross the road, indicative of heterogeneous traffic, as the pedestrian waits for the car to pass, 3) interaction between vehicles, such as a car potentially overtaking a bus if travelling at a higher speed, since the bus limits its way, 4) detecting intention, like a car planning to take a right shoulder and turn right, 5) multi-modal trajectories, exemplified by a car having several options at a given point, such as going straight or changing to the left lane.
To address these challenges, the research community has focused on designing models capable of capturing complex patterns essential for accurate prediction. Among the top-performing solutions is Trajectron++ [1], which integrates environmental context through a graph-structured recurrent architecture to model interactions between agents. The Spatio-Temporal Transformer Network (S2TNet) [2] leverages transformer layers to capture both spatial and temporal dependencies. Furthermore, from the intersection between academia and industry, the MultiPath++ model [3], released by the Waymo research group, employs multi-context gating fusion and dynamically learns reference points (anchors) to support goal-oriented predictions. Collectively, these models combine diverse techniques, such as graph-based interaction modeling, transformer architectures, and dynamic goal sampling, leading to more accurate and robust trajectory predictions. On the other side, the commercial research prioritizes real-time execution, cost-effectiveness, and handling uncertain situations. This leads each company to use its own combination of sensors, compatible with the set of algorithms it employs. Tesla’s prediction algorithm is an end-to-end vision-based deep learning model [^1] https://www.tesla.com/support/autopilot, forecasting trajectories based on the visual sequences and aggregated contextual information over time. The model continuously learns from the vast amount of data collected by Tesla’s global fleet of vehicles. In contrast, Waymo utilizes a combination of high-fidelity sensors (LiDAR, cameras, and radar). The sensor data is processed by the perception module and then passed to motion prediction and planning components. Waymo adopts a modular structure, ensuring that each component can be thoroughly tested independently and then integrated into a single pipeline [^2] https://waymo.com/research/. It is challenging to determine the specific algorithms employed in real operations, but Waymo has a strong R&D focus, actively publishing its latest research findings. Aurora company focuses on autonomous solutions for trucks serving logistics purposes. Its prediction module [^3] https://blog.aurora.tech/ performs joint object detection and trajectory forecasting, combining learning-based methods with rule-based constraints to align predictions with driving rules. To ensure safety across a wide range of scenarios, from highway driving delivering cargo between cities to urban logistics, Aurora’s system is tested extensively through a combination of virtual testing suites and real-world operations. Each company’s strategy reflects a balance between predictive accuracy and operational constraints, tailored to the specific purpose of their vehicles.
For evaluation and testing, two robust open-source platforms, Autoware [^4] https://github.com/autowarefoundation/autoware and Apollo [^5] https://github.com/ApolloAuto/apollo, provide environments to test perception modules. Autoware’s perception stack includes functionalities such as object recognition (2D/3D detection using LiDAR, cameras, and radar), obstacle segmentation (distinguishing drivable areas via point cloud processing), occupancy grid mapping (representing occupied and free spaces), and traffic light recognition (detecting light states and directions). However, it lacks an integrated prediction module. Starting from release 5.0, Apollo incorporates an integrated prediction module [^6] https://github.com/ApolloAuto/apollo/blob/master/modules/prediction/README.md, supporting a range of predictors: Free Move Predictor for non-lane-constrained obstacles, Single Lane Predictor for vehicles on a single lane, and Lane Sequence Predictor for multi-lane and behavior-based trajectory predictions. Later versions, such as 6.0.0 and 7.0.0, introduced advanced prediction models, including a semantic map-based pedestrian prediction model and the Inter-TNT prediction model, which leverage interactive prediction and planning evaluation. While Apollo provides a comprehensive platform for testing prediction algorithms within the full autonomous driving stack, there is a clear need for an open, lightweight platform that enables faster experimentation with prediction modules. Such a platform should integrate a perception component to provide realistic inputs for testing, while remaining flexible enough to support subsequent integration into full-stack systems.
For researchers entering the field of trajectory prediction, the volume of conducted works and proposed methods might be overwhelming, raising questions about how to categorize and classify existing approaches and which areas still offer room for improvement. For professionals experienced in the realm, keeping track of new emerging designs that lead to significant performance improvements, considering how rapidly the field is evolving, might be challenging. To assist both groups, this review covers a wide range of topics relevant to the prediction field, provides comprehensive overviews, and lists relevant references for readers to fetch additional information, which is either out of scope or briefly mentioned in the paper. In this review, we pose the following research questions:
  • 1.
    RQ1: What are the existing solutions for trajectory prediction? This question forms the foundation of the survey of modeling approaches. It seeks to identify the techniques and methodologies developed by the research community for trajectory prediction, covering the main modeling approaches used in recent years.
  • 2.
    RQ2: What are the emerging research trends in the trajectory forecasting realm? This question aims to discern the current focus of researchers and which pressing challenges they are addressing.
  • 3.
    RQ3: Is trajectory prediction problem solved or there is a room for improvement? To answer this question, we discuss the effectiveness of current state-of-the-art (SOTA) prediction modules are, what are current limitations and which research gaps are still existing.
To address the posed research questions, this study surveys the latest advancements in prediction models and offers a structured framework for understanding the trajectory forecasting landscape. We comprehensively review key factors that influence the performance and robustness of prediction models. This analysis aims to assist researchers in identifying gaps in the current literature and guiding future developments in the field. Our contributions are as follows:
  • 1.
    Provide an extensive overview of the prediction pipeline, which includes discussions of input modalities, modeling features and approaches, output modalities, evaluation metrics, and prediction paradigms.
  • 2.
    Devise a taxonomy and categorization of modeling approaches, offering a structured classification of methods. Each category includes a comparative table summarizing key models and their characteristics for quick reference.
  • 3.
    Provide an exhaustive literature review of active research areas in the trajectory prediction domain, covering topics such as uncertainty estimation, integration of driving knowledge, planning-conditioned prediction, vision-language model (VLM)-guided prediction models, and collaboration between vehicles.
  • 4.
    Provide an open repository [^7] https://github.com/AV-Lab/Trajectory_Prediction_Models_Zoo featuring two prediction pipelines to facilitate experimentation, benchmarking, and the development of new methods. The first pipeline includes multiple learning-based models that can be trained on SOTA datasets or custom datasets, enabling flexibility for diverse research needs. The second pipeline implements an integrated detect-track-predict approach, designed to process inputs from specified sensor configurations (e.g., frontal camera, stereo camera, LiDAR) for seamless trajectory forecasting in real-world scenarios.
Further, we discuss the related surveys conducted for trajectory prediction task and elaborate on the scope and structure of this study.
1.1 Related Surveys
Table 1: Comparison with Existing Review Papers: RP1 - Liu et al. [4], RP2 - Teeti et al. [5], RP3 - Leon and Gavrilescu [6], RP4 - Mozaffari et al. [7], RP5 - Huang et al. [8], RP6 - Ding and Zhao [9], RP7 - Ghorai et al. [10], RP8 - Rudenko et al. [11], RP9 - Sighencea et al. [12], RP10 - Korbmacher and Tordeux [13], RP11 - Zhang and Berger [14], RP12 - Gomes and Wolf [15], RP13 - Bharilya and Kumar [16]
Discussed Topics RP1 RP2 RP3 RP4 RP5 RP6 RP7 RP8 RP9 RP10 RP11 RP12 RP13 Our Review Trajectory prediction problem formulation Overview of prediction pipeline Overview of input and output modalities Summary of prediction datasets Survey of physical modeling approaches Survey of machine learning modeling approaches Survey of deep learning modeling approaches Survey of reinforcement learning modeling approaches Discussion of prediction paradigms Discussion on intention-aware approaches Discussion on interaction-aware approaches Discussion of active research topics Discussion of research gaps
Figure 2: Structure of the review paper: the study comprises seven main sections. The first two sections - an overview of the prediction pipeline and a survey of modeling approaches, provide general knowledge of the trajectory prediction domain. Next two sections are dedicated to incorporating driving knowledge and addressing uncertainty focusing enhancing robustness and generalization of prediction models. Last three sections cover planning-conditioned prediction, VLM-guided prediction, and collaboration between intelligent vehicles for trajectory prediction.
This section reviews key survey papers related to the trajectory prediction domain. Each paper is briefly discussed, emphasizing its major focus points. Table 1 provides a comparative overview of the topics covered in these surveys, contrasting them with the focus areas of this study. This comparison delineates the scope and layout of this review.
Liu et al. [4] review learning-based methods for trajectory predictions, considering aspects of representation, modeling, and learning. For representation, the authors examine agent state and scene context. In terms of modeling, the survey categorizes methods into feature encoding, interaction modeling, and generative models. Teeti et al. [5] focus their survey on vision-based intention and trajectory prediction approaches. All methods are summarized while detailing their architectures for input encoding, temporal and social encoding, decoding and loss function. The survey also discusses goal-conditioned methods and prediction in stochastic environments. As prospective research directions, it outlines real-time inference and the shortage of self-supervised solutions. Leon and Gavrilescu [6] conducted a survey on tracking and motion prediction methods. The authors argue that for effective trajectory prediction, a system should rely on physical and traffic rules constraints to narrow down the set of possible trajectories and select a reasonable subset. To classify prediction methods, it categorizes the existing approaches into neural networks-based, methods using stochastic techniques, and mixed methods that combine stochastic techniques with neural networks.
Next, Mozaffari et al. [7] review deep learning-based prediction methods, categorizing based on input representation, prediction method, and output type. For input representation, the paper identifies four classes: the past trajectory of the target vehicle, the past trajectories of the surrounding vehicles, a simplified bird’s-eye view (BEV) image with information about the road geometry, and raw sensor data. As research gaps, the paper highlights running time evaluation of prediction algorithms, assumption of full observability of BEV images, lack of integration of traffic rules, and generalizability of the prediction models across diverse scenarios. Huang et al. [8] base their classification on the prediction method, modeling of contextual factors, and generated output. For contextual factors, the paper distinguishes physics-related, road-related, and interaction-related factors. For prediction methods, it categorizes into physics-based, classic machine learning, deep learning, and reinforcement learning. The suggested potential research directions are inclusion of more interaction-related information and improving model robustness towards uncertainties and perturbations caused by tracking, location, and map errors. Ding and Zhao [9] review deep learning (DL) based models with a focus on the inclusion of driving knowledge. The study reviews how SOTA models incorporate the driving knowledge to enhance performance, explainability, and interpretability. Ghorai et al. [10] focus their survey on motion prediction and state estimation, through thoroughly analyzing and discussing sensory perception capabilities. The authors divide the discussion of perception capabilities into two categories: ego-vehicle perception and cooperative perception.
The next line of work focuses on pedestrian trajectory prediction. This task introduces an additional layer of complexity as pedestrian motion is less constrained compared to that of vehicles and is influenced by a broader range of factors. These include an individual’s own goal intent, actions of surrounding agents, the topology of the environment, social relations between agents, and social rules and norms. Some of these impacting factors are not directly observable and must be inferred or modeled from contextual information. Rudenko et al. [11] conducted a survey on pedestrian trajectory prediction and proposed to categorize prediction methods based on the motion modeling approaches and contextual cues. The authors divided modeling approaches into “sense-predict” (physics-based approaches), “sense-learn-predict” (pattern-based approaches) and “sense-reason-predict” (planning-based approaches). The contextual cues are categorized into cues of the static environment and dynamic agents. Sighencea et al. [12] concentrate their review on the types of sensors used for acquiring data for pedestrian trajectory prediction and the deep learning-based approaches employed. The authors offer an overview of LiDAR, RADAR, and camera sensors, discussing their comparative features and the associated processing methodologies to transform acquired data into predicted trajectories. In terms of prediction approaches, the paper distinguishes CNN-based solutions, sequential prediction algorithms (based on RNN or LSTM), and generative-based approaches. Next, the survey conducted by Korbmacher and Tordeux [13] compares knowledge-based approaches, used to simulate crowd dynamics, with deep learning-based methods. They explore the relevance of using knowledge-based methods for prediction, given the high accuracy of data-driven algorithms and their ability to learn collective dynamics. Zhang and Berger [14] conducted a review of deep learning-based approaches for predicting pedestrian trajectory and intention in urban scenarios. They categorize existing works by prediction tasks (trajectory, intention, and joint prediction), types of pedestrian behavior features considered by prediction models, and the network structures of these models (sequential and non-sequential). As the research gaps, the authors mention integration of appearance and skeleton behavioral features into prediction models and the lack of models predicting from raw sensor data without assumptions of perfect detector and tracker. Lastly, Gomes and Wolf [15] conducted a survey focused on intention and interaction-aware solutions. Within the realm of deep learning solutions, the survey highlights methods that utilize attention mechanisms, transformers-based solutions, and continual learning approaches.
The current survey aims to fill the gap by extensively analyzing recent research on trajectory prediction solutions. Given its comprehensive scope, structured approach, and focus on emerging topics, we believe this survey is timely and will serve a broad audience in the field.
1.2 Scope and Structure
The scope of this review is defined by the objective to offer a comprehensive overview of the trajectory prediction field for novice researchers and already experienced practitioners. Figure 2 illustrates the structure of our review, detailing the topics and main points discussed.
The first part consists of two sections aimed at providing general knowledge of the trajectory prediction realm. Section 2 provides an overview of a trajectory prediction pipeline and covers problem formulation and prediction paradigms. Section 3 lays out the taxonomy and categorization of the modeling approaches and elaborates on the models available in the literature from each category. The next part is dedicated to the discussion of factors impacting model robustness and generalization. This part also comprises two sections. Section 4 reviews the inclusion of driving knowledge in prediction models, covering different types of driving knowledge and the ways they are incorporated into prediction models. Section 5 is dedicated to uncertainty estimation and mitigation in prediction models. It elaborates on the sources of uncertainty and approaches to mitigate the risks associated with them. The last part depicts active research areas and is presented by three sections. Section 6 discusses planning-prediction-coupled frameworks, where for generating predictions the ego-vehicle models internal planning intentions of surrounding agents. Next, Section 7 discusses the application of Vision-Language models in scene understanding and trajectory prediction. Further, Section 8 discusses the concept of collaborative perception, elaborating on different types of fusion and different types of challenges appearing because of collaborating systems. Lastly, Section 9 provides a discussion on the posed research questions, outlines the research gaps, and highlights future research directions.
For a comprehensive survey of trajectory prediction methods, we recommend Section 3. Practitioners may find Section 5 particularly valuable, where sources of uncertainty are discussed. Researchers interested in specific topics within the domain of trajectory prediction should refer to Sections 6, 7 and 8.
2 Overview of Prediction Pipeline
Figure 3: Overview of trajectory prediction pipeline: including input modalities (past trajectories, sensor data, and HD maps), key modeling features (target agent, interaction, and environment features), and modeling approaches (conventional, deep learning-based, and reinforcement learning-based methods). The outputs of forecasting models can be unimodal and multimodal predictions, in the form of intentions, and occupancy maps.
This section aims to provide a comprehensive overview of a prediction pipeline. Figure 3 illustrates the general prediction pipeline, highlighting its main components. The overview begins with the problem formulation of trajectory prediction task, elaborates on input modalities, provides a discussion of modelling features and approaches, along with the classification of output modalities. The section continues by elaborating on metrics for performance evaluation and listing benchmark datasets with leader-board solutions. It concludes with a discussion of trajectory prediction paradigms.
2.1 Problem Formulation
A traffic scenario is characterized by an ego vehicle $ev$ and a set of agents $A=\{a_{1},a_{2},...,a_{n}\}$. Each agent $a_{i}$ is defined by a state vector $\tau_{i}$ $\in$ $\mathbb{R}^{t\>\times\>m}$, where $t$ indicates the length of observation and $m$ is an arbitrary dimension and in the most reduced form $(m=2)$. The state vector $(\tau)\in\mathbb{R}^{t\>\times\>2}$ for each agent is defined by its position $(x,y)$ over time:
$\tau_{i}=[(x_{i}^{1},y_{i}^{1}),(x_{i}^{2},y_{i}^{2}),...,(x^{t-1}_{i},y^{t-1}_{i}),(x_{i}^{t},y_{i}^{t})],\>\>\>\>t\in\mathbb{N}$
Depending on the type of sensors observations, the state vector ($\tau$) can additionally include orientation and speed of the agent. The prediction task is to estimate trajectory sequences $\zeta_{i}^{t+1:t+\Delta t}$ for each traffic agent $a_{i}$ and specified time horizon $\Delta t$ given the previous states:
$\zeta_{i}^{t+1:t+\Delta t}=\{(x^{t+1}_{i},y^{t+1}_{i}),(x^{t+2}_{i},y^{t+2}_{i})...,(x^{t+\Delta t}_{i},y^{t+\Delta t}_{i})|\tau_{i}\}$
in which $t$ is the current time frame and $(x_{i}^{k},y_{i}^{k})$ is the prediction state of the i-th agent. The goal of trajectory predictor ($\mathcal{P}$) is to estimate trajectory sequences $\zeta_{i}^{t+1:t+\Delta t}$ for $i\in\{1,2,...,n-1,n\}$ given the historical state of the traffic agents (T):
$\mathcal{P}(T)=\{\zeta_{1},\zeta_{2},...,\zeta_{n-1},\zeta_{n}\},\>\>\>\>n\in\mathbb{N}$
In the given formulation, the predictor ($\mathcal{P}$) is a function of a single parameter, past trajectories ($\mathcal{T}$). This can be expanded to make $\mathcal{P}$ a function of multiple parameters if additional inputs are available, e.g., road geometrical features ($G$) or interaction modeling features ($I$).
2.2 Input Modalities
In this subsection, we discuss the classification of input modalities and review studies in the literature for each category. We classify input modalities into three groups: past trajectories of agents, sensor data, and HD maps.
2.2.1 Past Trajectories
Past trajectory is a fundamental input in trajectory prediction models. Predicting future trajectories solely based on past trajectories [17, 18, 19, 20] forces the models to learn patterns by relying purely on spatio-temporal dependencies. This approach, which excludes additional contextual or environmental information, is efficient in terms of prediction speed and computational cost. However, it lacks the capacity to capture knowledge about the environment or interactions between agents, making it less robust in complex scenarios. One way to implicitly encode interactions is to include the past trajectories of surrounding agents in the prediction process. Studies [21, 22, 23, 24, 25, 26, 27] have demonstrated that including the trajectories of neighboring agents allows the model to infer interdependencies and interactions among them. This enables predictions that are more reflective of real-world behaviors, such as reactions to nearby vehicles or pedestrians. However, the main limitation of these models is reliance on error-free observations during training, using ground-truth past trajectories. In practical applications, the past trajectories available to prediction models are often noisy or incomplete (when we just start tracking the object, the missing observations will be zeros, until we accumulate the past trajectory of the required length), and robust handling of these inaccuracies is critical for practical deployment. When discussing past trajectories, they are typically represented as $(x,y)$ coordinates over multiple timestamps. However, additional motion states, such as heading and speed, can complement this information. These inputs can either be included as part of the agent’s state across multiple past timestamps or represent characteristics of the agent’s current state. Studies [28, 29, 30] have highlighted the importance of including velocity, acceleration, and orientation as input features to enhance prediction accuracy.
Models relying solely on past trajectories assume error-free perception, as they are trained on ground-truth data. In reality, perception systems often produce noisy or incomplete inputs. To address this, either the perception pipeline must be highly robust, or prediction models should incorporate additional input modalities, from which environmental context or agent interactions can be extracted, to improve accuracy and compensate for perception errors.
2.2.2 Sensors Data
Figure 4: Point clouds captured using a 64-layered Ouster LiDAR and 64-layered Kinetic LiDAR: (a) in a parking area, (b) on a campus, (c) during an Autonomous Racing Competition (A2RL) with the LiDAR mounted on a Formula 1 car, and (d) at A2RL in a top-down view.
Sensor data, provided by exteroceptive sensors such as LiDAR, radar, and cameras, enables prediction models to capture environmental context and integrate it into predictions. This section first examines each sensor individually and then explores approaches that leverage multi-modal inputs for enhanced prediction performance.
LiDAR. LiDAR sensors enable environmental perception through laser-based distance measurements, making them indispensable for autonomous driving systems. Two types of LiDAR are commonly employed: perimeter LiDAR and $360^{\circ}$ LiDAR. Perimeter LiDAR, mounted on the vehicle’s edges or corners, is optimized for monitoring close-range objects, blind spots, and collision risks, particularly in scenarios like parking or navigating tight spaces. In contrast, $360^{\circ}$ LiDAR, typically roof-mounted, provides a complete $360^{\circ}$ field of view, supporting long-range detection and situational awareness. Current SOTA datasets predominantly include data from $360^{\circ}$ LiDARs (please refer to Figure 4 for examples), while perimeter LiDARs are used by automotive manufacturers as complementary sensors to enhance overall perception. One of LiDAR’s key strengths is its ability to function reliably in all lighting conditions, including complete darkness and bright sunlight, due to its light-invariant nature. However, its performance degrades in adverse weather, such as heavy rain, or snow, where laser beam scattering reduces detection range and accuracy along with highly reflective surfaces such as glass, potentially creating blind spots. Despite these challenges, LiDAR remains highly effective for detecting obstacles [31], vehicles [32, 33, 34], lane markers [35, 36, 37, 38], and forecasting [39].
Luo et al. [40] proposed a deep-learning model for 3D detection, tracking, and motion prediction using LiDAR data as input. The model processes point clouds with 3D convolutions to extract spatial and temporal information, completing all tasks within 30 milliseconds on a custom dataset recorded with LiDAR-equipped vehicles. Further, Völz et al. [41] developed a LiDAR-based model for classifying pedestrian intentions at urban crosswalks. This approach projects LiDAR point clouds into a 2D representation based on angular coordinates. The resulting 2D images are passed to CNNs to extract spatial and motion-related features, which are further combined with pedestrian velocity and distance to the crosswalk for intention classification.
Two significant limitations hinder the widespread adoption of LiDAR technology are its relatively high cost and its accuracy in detecting pedestrians. Although LiDAR technology is advancing and prices are gradually declining, it remains expensive and impractical for consumer-level mass-market adoption. The second limitation is its lower accuracy in pedestrian detection. For example, on the KITTI leaderboard, the top pedestrian detection score is only 56% [42], compared to a 93% vehicle detection score achieved by [43]. This disparity stems from LiDAR’s lack of rich semantic information, which is crucial for accurate pedestrian detection. As a result, LiDAR systems are ideally complemented by cameras, which provide the necessary semantic context to enhance pedestrian detection performance.
Camera. Cameras play a crucial role in perception by scanning the environment and detecting traffic participants, road signs, and traffic lights. These sensors are characterized by their resolution, field of view, and configuration. RGB cameras are the standard visual sensors for AVs and are widely used to collect data for current SOTA datasets. The vision community for RGB videos is highly advanced, providing efficient algorithms for object detection [44], semantic segmentation [45], and lane recognition [46]. Camera setups in AV systems are generally categorized as single frontal cameras [47] or multi-view configurations [48]. Single frontal camera setups primarily rely on 2D perception and are mostly used in driving assistance systems. In contrast, multi-view setups arrange cameras to ensure complementary, non-overlapping fields of view, providing 360-degree observability. This configuration enables 3D perception [49] and close-range monitoring, especially in urban and dynamic driving conditions. Additionally, camera systems can employ either monocular or stereo setups. Monocular cameras [44, 45, 46, 49] are lightweight and cost-effective but restricted to 2D image data, relying on advanced algorithms for depth estimation. In contrast, stereo cameras utilize dual lenses to calculate depth through stereo triangulation, enabling accurate 3D perception [50], which is essential for tasks like depth estimation and obstacle avoidance.
Bhattacharyya et al. [51] proposed a prediction model that uses data from an onboard monocular camera. Pedestrian bounding boxes detected in the camera feed are combined with the ego-vehicle’s odometry information and processed by a two-stream Bayesian RNN encoder-decoder architecture. The encoded bounding boxes and odometry data are used to predict pedestrian trajectories over a one-second horizon. Similarly, Zhong et al. [50] designed a system utilizing a stereo camera setup to estimate 3D human poses and predict pedestrian trajectories. A twin network processes the video streams from each camera independently to estimate 2D poses. These 2D poses are then aligned and fused using a stereo consistency constraint to produce accurate 3D pose estimates, which are subsequently processed by an adapted SocialGAN framework to generate 3D pedestrian trajectories.
RGB cameras, however, significantly degrade in performance under low-light or adverse weather conditions. To address these limitations, thermal and infrared cameras can serve as complementary sensors. Thermal cameras [^8] https://www.flir.com/oem/adas/adas-dataset-form/, which detect heat signatures, are highly effective for detecting pedestrians and animals in complete darkness or adverse weather conditions. However, their shorter range and lack of fine detail limit their utility for tasks like detecting road signs and require specialized deep learning solutions. Infrared cameras enhance visibility in low-light conditions and are suitable for detecting road signs, detecting and tracking pedestrians and vehicles [52]. Nonetheless, they are sensitive to bright light sources, which can cause glare and blooming artifacts. This issue can be mitigated with filters designed to reduce glare and enhance image quality in challenging scenarios. Despite their individual limitations, combining thermal or infrared cameras with RGB cameras creates a more robust sensor suite, ensuring reliable operation across diverse environments.
Radar. Radar sensors are deployed to estimate the relative velocity, distance, and direction of targets by analyzing reflected electromagnetic waves [53]. Modern automotive radar systems provide coverage with angles up to 150° and distances of up to 250 meters [54], making them suitable for detecting objects in diverse driving scenarios. Frequency-Modulated Continuous Wave (FMCW) radar is a key technology in these systems, enabling precise measurements. By comparing transmitted and reflected signals, the distance of a target is calculated from the beat frequency, while the relative velocity is determined through Doppler shift analysis. For direction, FMCW radar pairs with digital beam-forming [55] or multi-antenna setups to estimate the angle of arrival of reflected waves.
Kim et al. [56] presented a model to classify agent types (pedestrian, cyclist, or vehicle) using radar data. Radar signals are transformed into range-velocity (RV) images that capture an object’s distance and motion. These RV images are processed by a convolutional LSTM to learn spatial and temporal features, followed by a fully connected layer to map these features to target classes. Steinhauser et al. [57] proposed a system focused on individual limb movements for pedestrian behavior prediction. Their method employs a deconvolution algorithm combined with a clustering technique to extract and analyze reflections from pedestrian body parts, decomposing the range-Doppler spectrum to identify the torso, arms, and legs. Despite all advancements, radar remains primarily a supplementary sensor in SOTA models, complementing cameras and LiDAR by providing reliable velocity and heading information.
Rasterization. Rasterization converts raw sensor data into structured, grid-like representations, enabling prediction models to effectively process spatial and temporal information. For example, LiDAR point clouds can be rasterized into occupancy grids, where each cell indicates whether it is occupied or free. Radar data adds dynamic attributes such as velocity and direction [58], which can be encoded either through color coding (e.g., higher speeds represented by brighter colors) [59] or as separate velocity channels, storing velocity values in an additional data layer for each grid cell. Similarly, camera images can be rasterized to map rich semantic information, such as object classes and bounding boxes. This unified rasterized format facilitates the fusion of multimodal sensor data and enables machine learning models to process environmental context with high precision.
The most common type of rasterization in autonomous vehicles is Bird’s Eye View (BEV) rasterization [60, 61, 62, 63, 64, 65], where sensor data is transformed into a top-down perspective of the environment. In BEV rasterization, dynamic parameters such as positions and velocities are mapped into a unified plane aligned with the ego vehicle. This format often employs color-coded semantics to differentiate objects and incorporates temporal information through stacked frames or brightness variations to represent motion over time. Simplified BEV representations extend this concept by depicting static and dynamic elements, e.g., road lanes, vehicles, and pedestrians, using polygons and lines.
2.2.3 HD Maps
Figure 5: BEV representation aligning dynamic and static elements: raw point cloud data, HD map, and ground truth 3D bounding boxes.
Latest datasets supplement collected data with High Definition (HD) maps, which provide detailed representations of static road features, such as lane markers, lane edges, pedestrian crossings, road layouts, and stop signs. Typically, the map is presented as a single large file and requires pre-processing to extract locally relevant environmental information. A Region of Interest (ROI), usually encompassing 50–100 meters around the vehicle’s position, is extracted to capture nearby lanes, crosswalks, traffic signs, and other relevant features. Some SOTA datasets, such as NuScenes [66] and Argoverse [67, 68], provide SDKs to query map data and efficiently extract ROIs. This ROI is then transformed into a BEV image, where semantic elements like drivable areas and lane boundaries are encoded using distinct colors or layers. Prediction models can process this BEV image independently, combining it with past trajectories for improved trajectory forecasting [69, 70, 64], or use it in conjunction with sensor data to create a combined rasterized representation [71, 1] (please refer to Figure 5 for illustration).
Incorporating information about the driving environment from HD maps significantly enhances the accuracy of prediction models by constraining predictions to plausible outcomes. However, this approach requires precise localization of the ego-vehicle and is sensitive to localization errors. Additionally, constructing HD maps involves substantial costs and ongoing efforts to update them. To alleviate the human labor associated with constructing HD maps, the research community is actively exploring methods for automatically constructing them from raw sensor data [72, 73, 74, 75, 76]. Combining HD maps with sensor data offers a promising solution to mitigate localization errors and handle changes in the road environment, such as road closures or modifications to traffic signs. However, this approach requires additional computational pre-processing, which impacts the inference time of the prediction pipeline.
2.3 Modelling Features
Broadly, modelling features can be categorized into static and dynamic. For static elements, the actions of traffic agents depend on the physical layout of the environment. Methods that do not account for the environment’s topology assume agents operate in an open-space environment. For the dynamic elements, prediction models extract features related to the target agents and their interactions. Further, we elaborate on each category.
2.3.1 Environment Features
Traffic agents rely on the environment’s topology to plan their potential routes, and this geometric information can significantly enhance prediction accuracy [77, 78, 79, 80, 81]. One approach to integrating environmental features is through explicit modelling. Ma and et al. [29] incorporated the distance to the boundaries of crossroads into their prediction model, while Yang and et al. [82] included the presence of traffic lights, stop signs, and zebra crossings. Similarly, Pool et al. [80] incorporated static features for cyclist path prediction by estimating the distance of cyclists to the nearest intersection. This estimation was passed as a contextual feature to a GRU-based RNN, enabling the model to account for the impact of intersections on trajectory predictions.
Rehder et al. [81] presented a more advanced approach by constructing a topological map that organizes semantic features, such as roads, sidewalks, and obstacles, into an occupancy grid. A recurrent Mixture Density Network (rMDN) predicts potential pedestrian destinations, which are integrated into the topological map to encode spatial relationships and constraints. This map serves as the basis for trajectory prediction, ensuring paths align with environmental features.
Another method is implicit modelling, where environmental features are integrated indirectly, focusing on their influence on agent behavior without explicitly extracting or representing them. Bera et al. [83] employs implicit modelling through the Reciprocal Velocity Obstacle (RVO) framework, which adjusts pedestrian trajectories to account for static obstacles like buildings or parked cars without explicitly modelling these elements.
2.3.2 Target Agents
As discussed in Section 2.2, prediction methods generate forecasts based on the past trajectories of target agents [84, 85, 86, 87], and/or incorporate velocity history [88, 41, 89], as well as heading and orientation [90, 91], to predict future trajectories. In heterogeneous models, including the category of the object (e.g., person, truck, cyclist) enables prediction models [92, 93] to learn category-specific dynamics, although these methods may be prone to errors caused by the incorrect target agent classification by the detector. For vehicles, learning individual behavioral patterns and motion dynamics from historical data [94] can improve prediction accuracy by capturing unique driving styles. Similarly, in pedestrian trajectory prediction, appearance and posture offer valuable insights into pedestrians’ intentions and planned trajectories. Studies [95, 96, 97] have focused on extracting visual cues to capture behavioral patterns and improve prediction accuracy. To estimate posture, two main approaches are employed: using skeleton key-points [50, 98, 99, 100] or leveraging pretrained CNNs [101, 102, 98].
Incorporating additional semantic attributes requires either integrating a subnetwork within the model to extract these attributes from the input modalities or utilizing networks pre-trained for these tasks, e.g., estimating head direction and gestures in pedestrian behavior prediction. In the first case, the availability of annotated data containing such attributes is often limited, and errors in predicting these attributes can impact overall accuracy. In the second case, the trade-off between the accuracy boost from including these attributes and the added time complexity to the inference pipeline must be carefully evaluated.
2.3.3 Interaction Features
The integration of autonomous vehicles into human-driven traffic requires a deep understanding of socially compatible communication among traffic agents and the ability to interpret these interactions. To ensure smooth navigation and safe arrival at destinations, trajectory prediction models must account for interaction-related factors [103, 104, 105, 106]. This is particularly critical in dense, interaction-intensive environments, where avoiding collisions depends on accurately modelling the interplay between traffic agents. A simple approach to integrating interactions involves using hand-crafted features, such as relative position and speed of neighboring agents, the distance between vehicles and pedestrians, and time to collision [29, 107, 82, 108]. While straightforward, this method often fails to capture the complexity and variability of interaction patterns in real-world scenarios. More sophisticated techniques for modelling interactions include game theory, social field models, and graph-based approaches.
This section discusses Computational Cognition and Social Pooling methods, while GNN-based modelling of interactions between traffic agents is addressed in Section 3.4. Social Pooling is a widely adopted learning-based method in the literature. For a comprehensive review of other approaches and a detailed study on interaction modelling for autonomous driving, readers are referred to the most extensive review available on this topic [109].
Computational cognition methods are based on the premise that interactions between agents are underpinned by cognitive mechanisms, suggesting that a deeper understanding of cognitive science can guide the development of socially-compatible autonomous driving systems [110]. A key theory in this context is Theory of Mind (ToM), which posits that humans can infer the intentions of others by reasoning from their perspective. Cuzzolin et al. [111] argue that integrating ToM into intelligent vehicles is essential for enabling them to share the road with humans effectively, as naive pattern recognition approaches often fail to capture the nuances of spontaneous human behavior. Leveraging knowledge from traffic psychology, particularly about humans’ ability to infer others’ beliefs and intents without explicit communication, can formalize machine ToM for autonomous vehicles [112].
To implement machine ToM, Inverse Reinforcement Learning (IRL) provides a framework to model how agents’ actions are influenced by the presence of others. Sun et al. [113] applied hierarchical IRL to learn distributions of future trajectories from human demonstrations, capturing multi-agent interactions. Ozkan and Ma [114] introduced a control design for autonomous vehicles that accounts for socially desirable behavior in car-following scenarios, embedding the influence of one agent on another. Tian et al. [115] extended this approach by proposing an IRL framework to infer humans’ latent intelligence levels, demonstrating how agents reason about mutual decision-making in game-theoretic scenarios.
Social pooling is one of the most common learning-based approaches for modelling interactions between agents. It involves extracting relevant features from neighboring agents and aggregating them into a pooled representation, typically using max or average pooling. Social-LSTM [116] is a pioneering model in this type of interaction modelling, introducing a social pooling layer that integrates the hidden states of individual LSTMs (for each pedestrian in a scene) to account for social interactions. Subsequent studies have refined this concept by incorporating attention mechanisms to model interactions more effectively [117, 118]. Xu et al. [119] introduced a weighted spatial affinity function, where weights are computed based on spatial features to represent interaction strength. Similarly, SoPhie model [120] prioritized attention to closer objects, reflecting the idea that agents are more influenced by their immediate neighbors. These advancements highlight the evolution of social pooling methods, moving from uniform aggregation to more context-aware interaction modelling. Amirian et al. [121] included attention-based social pooling block, which encodes geometric features such distance, bearing angle into interaction features via a fully connected layer. The attention weights are computed through scalar products with LSTM-encoded hidden states.
On more advanced pooling mechanism architectures, Gupta et al. [122] use relative positions of agents and process them through a Multi-Layer Perceptron (MLP) followed by max pooling to create a compact and globally aware representation, capturing both local and distant interactions. Convolutional Social Pooling (CSP), introduced by Deo and Trivedi [62], leverages spatial grid representations and convolutional operations to model interactions. Each vehicle is represented by an LSTM, and the interaction and spatial features are captured through a CNN that processes a spatial grid relative to the target vehicle. This grid encodes the states of surrounding vehicles’ LSTMs, enabling the pooling block to hierarchically extract global and local interactions. The CNN output, combined with the target vehicle’s LSTM encoding, forms the input to a decoder that predicts both future trajectories and maneuvers.
2.4 Modelling Approach
The choice of a modelling approach depends on the application requirements, platform characteristics, and environmental complexity, such as crowded scenarios or off-road autonomy [123, 124]. Traditional model-based methods, like the Constant-Velocity model or Interactive Multiple Model, explicitly encode motion constraints and are efficient for short-horizon predictions. Physics-based methods are discussed in more details in Section 3.2. For long-horizon predictions, data-driven approaches, which can incorporate spatial and temporal contexts, are more suitable. One of the categories are sampling methods [125, 126, 127, 128, 129, 130, 131, 132], which generate plausible trajectories by modelling the distribution of future trajectories conditioned on the past. These methods estimate the parameters of the distribution from observed data and sample future trajectories given the past trajectory. Depending on the modelling approach, sampling can be parametric, such as Gaussian or Gaussian Mixture Models, or learning-based, like GANs or Autoencoders. For parametric methods, Makansi et al. [125] introduced a two-stage framework that employs a modified Evolving Winner-Takes-All (EWTA) loss during sampling to produce diverse and representative hypotheses from Gaussian distributions. A second network in the framework is used to refine the drawn samples to better align them with the true data distribution. In learning-based methods [126, 127, 128, 129, 130, 131, 132], models are trained to map the relationship between past and future trajectories into a latent space. At inference, given a latent variable and past trajectory, the model generates plausible forecasts. Ma et al. [128] proposed a likelihood-based approach that optimizes for high-likelihood predictions while incorporating diversity-aware losses to represent multiple future behaviors. Gilles et al. [129] introduced a heatmap-driven method that encodes spatial reasoning and agent interactions, producing interpretable likelihood maps for socially-aware multi-agent sampling. We elaborate further on learning-based methods with latent spaces in Sections 3.4.3 and 3.4.4.
On more recent approaches, diffusion models [133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143] have been actively utilized in trajectory forecasting applications. These models involve a two-phase process: a forward step where Gaussian noise is progressively added to the data, transforming it into a noisy latent representation aligned with a simple prior distribution, and a reverse step where the model iteratively learns to denoise this representation conditioned on past trajectories and scene context. During inference, random Gaussian noise serves as the starting point, and the denoising process progressively refines it into a plausible future trajectory. By conditioning on contextual information such as past motion and environmental features, diffusion models generate diverse, multimodal, and context-aware trajectory predictions. Jiang et al. [137] proposed MotionDiffuser, which leverages a permutation-invariant transformer-based denoiser with cross-attention mechanisms to integrate scene context, past trajectories, and agent interactions. It employs PCA to compress trajectories into a latent space for computational efficiency and supports controlled sampling via differentiable cost functions, enabling constraints such as collision avoidance. Barquero et al. [138] introduced the BeLFusion model, a diffusion framework for behavior-driven human motion prediction, which uses a behavior coupler to conditionally sample from a behavioral latent space, ensuring realistic and behaviorally coherent trajectories. Yuan et al. [136] developed PhysDiff, integrating a motion projection module with a physics simulator to iteratively guide the denoising process, eliminating artifacts like floating and ground penetration, and producing physically accurate trajectories.
Continuing with data-driven models, learning-based prediction methods [144, 145] learn motion dynamics from driving data and are well-suited for complex scenarios. However, their generalization and robustness heavily depend on data variability, often requiring additional mechanisms that increase architectural complexity and computational demands. Among Machine Learning (ML) methods, techniques Gaussian Processes, Hidden Markov Models, Dynamic Bayesian Networks, and Gaussian Mixture Models are discussed in Section 3.3. While ML methods are effective in simpler environments with limited interactions, Deep Learning (DL) methods excel at long-term predictions in dynamic environments due to their ability to learn hierarchical and non-linear representations. Section 3.4 discusses key DL approaches, including Sequential Models, Convolutional Neural Networks, Graph Neural Networks, Transformers, Autoencoders, and Generative Adversarial Networks. Another approach, Reinforcement Learning (RL), assumes agents aim to achieve goals by optimizing policies that maximize rewards or minimize costs. RL is particularly suited for trajectory prediction tasks involving interactions and decision-making, as it estimates underlying cost functions to identify optimal policies. RL approaches are further elaborated in Section 3.5.
Modelling interactions between road agents has significantly enhanced the performance of learning-based methods. Interaction-aware approaches [146, 147, 148, 149, 150, 151, 152] explicitly model each agent’s surroundings to capture inter-dependencies between their actions while adhering to social norms. In mixed traffic scenarios, where autonomous vehicles operate alongside human-driven vehicles, achieving safe and efficient navigation requires understanding and interpreting the implicit and explicit ways road users communicate. Incorporating interaction-related factors into trajectory prediction models is essential for enabling autonomous vehicles to avoid collisions, particularly in dense and interaction-intensive environments. This survey discusses interaction modelling across various approaches, emphasizing its importance in developing socially-compatible and reliable trajectory prediction systems. Another line of approaches focuses on intention-aware models. These methods [153, 154, 155, 156, 157, 158, 159] incorporate the vehicle’s maneuver intentions to predict its future states more accurately. Maneuvers are defined as short-term, goal-oriented decisions comprising a sequence of continuous states that collectively lead to achieving a global objective. They can be broadly categorized into two types: lateral maneuvers (e.g., maintaining the current lane, turning left or right, changing lanes) and longitudinal maneuvers (e.g., maintaining speed, accelerating, or decelerating). Integrating intention-awareness introduces an intermediate layer into prediction models, where future forecasts are informed by planning-based reasoning rather than being treated as random or purely reactive paths.
2.5 Output Modalities
The primary classification of trajectory prediction outputs is based on whether they are unimodal or multimodal. Unimodal outputs are deterministic trajectory points or parameters of a unimodal trajectory distribution. Several studies [22, 92, 27, 24] focus on generating unimodal predictions. However, a significant limitation of unimodal approaches is their inability to capture the inherently multimodal nature of driving behaviour. For instance, in a lane with arrows indicating both straight and left turns, a vehicle could either proceed straight or turn left. Until the driver signals their intent (e.g., by activating an indicator or beginning to incline left), both options should be considered as equally plausible. Ignoring these multimodal characteristics makes it difficult to ensure collision avoidance in complex traffic scenarios, such as lane merging or navigating intersections.
Multimodal outputs address this limitation by generating predictions as a multimodal probability distribution [19, 20, 64]. These outputs better align with real-world traffic scenarios, though they present greater challenges in training. A key advantage of multimodal distributions is their ability to associate probabilities with each possible trajectory. This capability allows models to reflect prediction confidence. For example, in scenarios where an object is partially occluded or just entering the field of view, prediction certainty naturally decreases. For objects, that just entered the scene, robust predictions require past trajectory data of at least 7–10 frames; before this is achieved, models often pad past trajectories with zeros to the desired length. This lack of historical data should be accounted for and reflected in the output probabilities. Furthermore, by explicitly modelling uncertainty, multimodal predictions improve downstream planning by enabling risk-aware decision-making, such as adjusting speed during ambiguous situations or prioritizing safer maneuvers.
A portion of studies focuses on predicting manoeuvre intentions as output [18, 17, 21, 26, 60]. Common manoeuvre categories include changing lanes, staying in the current lane, going straight, and turning left or right. For the joint prediction of trajectories and manoeuvres, two main approaches exist. The first approach involves multi-task models [160], where manoeuvre intentions and trajectories are predicted as separate outputs but share a common feature extraction module. The second approach [161, 162, 163, 101, 102] predicts manoeuvres and trajectories separately, with the predicted manoeuvre used to guide trajectory prediction.
Lastly, occupancy grid and occupancy flow are alternative output formats for prediction models. These methods [164, 165, 166] estimate the future occupancy of each cell in a BEV map of the driving area. Occupancy grids provide a spatial representation by predicting the probability of each grid cell being occupied at future time steps [167, 168] and directly estimating spatial probabilities for the entire scene. However, the accuracy of these predictions is constrained by the resolution of the grid cells. Increasing the number of cells improves precision but significantly raises computational demands. A potential solution is to use a flexible grid resolution that adapts to the environment: finer grids for dense urban traffic and coarser grids for highways, optimizing the balance between precision and computational efficiency. Occupancy flow extends the concept of occupancy grids by not only estimating the future occupancy of grid cells but also capturing the motion dynamics within these cells [169, 170, 171]. While occupancy grids encode only spatial information, occupancy flow incorporates temporal changes by predicting both the future occupancy and the flow, reflecting how the objects will move. By modelling motion dynamics at the grid-cell level, occupancy flow eliminates the need for explicitly tracking individual past trajectories, as it encodes both spatial and temporal information directly from the input data. These properties, make both these formats suited for complex traffic scenarios and smooth integration with planning module.
2.6 Performance Evaluation
The variety of datasets requires selecting suitable evaluation metrics to fairly compare and evaluate prediction methods. As discussed in the Section 2.5, unimodal methods predict a single deterministic trajectory or the parameters of a unimodal trajectory distribution. In contrast, multimodal methods generate multiple possible trajectories or a multimodal probability distribution. Below, we review evaluation metrics for both unimodal and multimodal methods, where $\zeta_{i,t}$ and $\hat{\zeta}_{i,t}$ denote the predicted and ground truth trajectories for the i-th agent at time step $t$, respectively.
  • 1.
    Root Mean Squared Error (RMSE) computes the average of the squared error between predicted $\zeta_{i,t}$ and ground truth $\hat{\zeta}_{i,t}$ trajectories at time step $t$ for $n$ agents:
    $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}|\zeta_{i,t}-\hat{\zeta}_{i,t}|^{2}}$
  • 2.
    Average Displacement Error (ADE) estimates the average L2 distance between predicted and ground truth trajectories over the entire prediction horizon $\Delta t$:
[…]
    For multimodal predictions, the metric MinADE_k computes the L2 distance between the ground truth trajectory and the closest prediction out of $k$ possible trajectories:
[…]
  • 3.
    Final Displacement Error (FDE) estimates the average L2 distance between predicted and ground truth trajectories at the final time step $\Delta t$:
[…]
    For multimodal predictions, FDE can be extended to MinFDE_k:
[…]
  • 4.
    Negative Log Likelihood (NLL) evaluates the accuracy of probabilistic prediction algorithms by measuring the likelihood of the ground truth trajectory under the predicted distribution $q(\hat{\zeta}_{i,1:\Delta t}|\Theta)$, where $\Theta$ represents the model parameters:
    $NLL=\sum_{i}-\log(q(\hat{\zeta}_{i,1:\Delta t}|\Theta))$
  • 5.
    Miss Rate (MR) measures the ratio of cases where the predicted trajectory deviates significantly from the ground truth, based on criteria such as a final displacement threshold (e.g., greater than 2 meters).
Lastly, measuring and reporting inference time is crucial for a thorough evaluation of prediction algorithms, as it reflects their suitability for real-time operation. To compute the reported metrics accurately, alignment between predicted and ground truth trajectories is required. In the absence of explicit object identifiers, this alignment can be performed using proximity-based matching techniques, such as IoU thresholds or nearest centroid matching. Once the alignment is completed, the reported metrics can be calculated to assess prediction accuracy.
Table 2: Benchmark Datasets for Trajectory Prediction in Autonomous Driving
Dataset Modality Size Description Leaderboard KITTI [172], prediction split from [173] RGB Camera,   LiDAR,     GPS/IMU 1.5 h Includes 6-second segments. Training set: 8,613 top-view trajectories, test set: 2,907 samples. MANTRA [172] (ADE@4s:0.74, FDE@4s:1.68) Daimler [174] Stereo RGB, Camera 19,612 images Short sequences less than 2 seconds, contains 19,612 stereo image pairs. - Ko-PER [175] RGB Camera, LiDAR 4 chunks, 6.28 min each Seq. 1 has 4 chunks capturing several agents crossing the intersection, Seq. 2 and Seq. 3 featuring single vehicle at the intersection. - ApolloScape [176] RGB Camera, LiDAR,    GPS/IMU 2 h Each frame annotations contain objects type, position in world coordinates, dimensions, and heading indicating the steering angle. AI-TP [177] (WSADE:1.15, WSFDE:2.12)          S2TNet [2] (WSADE:1.16, WSFDE:2.17) PIE [101] RGB Camera 6 h Collected over 6 hours of driving. Contains intention estimation and features to study pedestrians’ behaviour. SGNet [178] (MSE@0.5:34, MSE@1.5:442)        Bitrap-D [179] (MSE@0.5:41, MSE@1.5:511)        PIE_traj [101] (MSE@0.5:58, MSE@1.5:636) Argoverse I [67] RGB Camera, LiDAR,     GPS/IMU 324,557 trajectories Collection of 324,557 scenarios, each 5 seconds long featuring 2D BEV trajectories. The dataset is accompanied with high-definition maps. HPNet [180] (MinADE:1.59, MinFDE:3.46)        iDLab-SEPT [181] (MinADE:1.44, MinFDE:3.17)        QCNet-AV1 [182] (MinADE:1.52, MinFDE:3.34) Argoverse II [68] [183] RGB Camera, LiDAR,     GPS/IMU 250,000 trajectories Collection of 250,000 scenarios, each 11 seconds long featuring 2D BEV trajectories and headings. iDLab-SEPT++ (MinADE:1.41, MinFDE:3.51) iDLab-SEPT [181] (MinADE:1.48, MinFDE:3.7)     GACRND-XLAB (MinADE:1.56, MinFDE:3.91) nuScenes [66] RGB Camera, LiDAR,     GPS/IMU, Radar 6 h 1,000 urban scenes, including LiDAR, radar, and cameras data, paired with high-definition maps. UniTraj [184] (MinADE@5:1.18, MinFDE@1:6.6) FRM [185] (MSE@0.5:41, MSE@1.5:511)    LaPred++ [186] (MinADE@5:1.24, MinFDE@1:7.58) WAYMO [187] RGB Camera, LiDAR,     GPS/IMU, Radar 570 h 570 hours of unique data across 1,750 km of roadways. Over 100,000 scenes collected. MTR_v3 [188] (mAP:0.49, minADE:0.55)                ModeSeq (mAP:0.47, minADE:0.56)               RMP_Ensemble (mAP:0.47, minADE:0.55) Lyft Level 5 [189] RGB Camera, LiDAR,    GPS/IMU, Radar 1,118 h Collection of 170,000 scenes, each 25 seconds long, including vehicles, cyclists, and pedestrians. -
2.7 Datasets and Leaderboard
Assessing the effectiveness of trajectory prediction models involves not only selecting accuracy metrics but also designing diverse test scenarios to evaluate the models’ performance under various factors, such as the number of interacting agents and the degree of maneuvering in the data. The test scenarios can vary widely in complexity, location, and traffic density, including intersections, highways, mixed urban streets, crowded public spaces, residential areas, and more. Each scenario presents unique challenges and variability that the algorithm must handle effectively to demonstrate its practicality and potential for real-world autonomous driving applications. Table 2 lists trajectory prediction datasets that can be utilized to test prediction models. It provides details on the modalities used to collect the datasets, along with short descriptions and leaderboard solutions. For some datasets, certain details are omitted, encouraging readers to refer to the datasets’ official websites for the most up-to-date information.
The KITTI dataset [172] is a well-known resource in the autonomous driving community. The official dataset does not provide an official split for trajectory prediction. The lack of a standardized data split causes inconsistencies in baseline comparisons across different models. Marchetti et al. introduced a well-defined split of the KITTI data. Samples are collected in 6-second segments with 2 seconds of past data and 4 seconds of future data. The dataset comprises 8,613 top-view trajectories for training and 2,907 for testing, supplied with top-view maps. The recent addition by the authors is KITTI-360 [190], collected over a distance of 70 kilometers and geo-localized with OpenStreetMap. The dataset serves as a benchmark for scene understanding and is a step forward toward map-less navigation, encouraging models to use only the global map without relying on the granularity of HD maps. The Daimler dataset [174] was introduced for predicting pedestrian paths over short time horizons (less than 2 seconds). The dataset includes 19,612 stereo image pairs, where 12,485 images are manually labeled with pedestrian bounding boxes, and 9,366 images contain measurements from a pedestrian detector. The Ko-PER Intersection dataset [175] provides data for intersection analysis. The dataset contains raw LiDAR measurements and undistorted monochrome camera images. It comprises 3 sequences: sequence 1 captures different agents crossing the intersection, sequence 2 contains data of a single car performing a right turn, and sequence 3 features a car crossing the intersection straight ahead. The dataset comes with a viewer and sample code to ease inspection and data transformation for researchers.
Next, the ApolloScape dataset [176] contains camera-based images and LiDAR point clouds of trajectory data collected in Beijing, China. The dataset was captured under various lighting conditions and features complex traffic flows. Each training file is a sequence 1 minute long. Annotations contain object types, positions, dimensions, and headings. The evaluation protocol defines prediction over the next 3 seconds. The PIE [101] dataset was collected using a monocular dashboard camera over 6 hours of driving in Toronto, Canada. The dataset captures various pedestrian behaviors at crosswalks under different urban settings and weather conditions. It is annotated with bounding boxes, occlusion flags, pedestrian actions and intentions, and ego-vehicle information such as speed, heading angle, and GPS coordinates. The NuScenes dataset [66] provides data for trajectory prediction of various traffic participants, such as vehicles, pedestrians, and cyclists. The data is collected from 1,000 urban scenes and includes sensor data from LiDAR, radar, and cameras, paired with high-definition maps.
The WAYMO Motion Prediction Challenge [187] introduced a diverse interactive motion dataset. This dataset includes around 100,000 scenes, each 20 seconds long, and features detailed 3D bounding boxes and high-definition maps. The Argoverse dataset is broadly used by trajectory prediction community. It is released in two versions. Version I [67] is a collection of 324,557 scenarios, each 5 seconds long, featuring 2D BEV trajectories of tracked objects. It includes complex driving scenarios such as intersections and lane changes. The dataset is accompanied by high-definition maps, aiding in the precise prediction of object motion. Version II [68, 183] was released two years later. It consists of 250,000 scenarios, each 11 seconds long, capturing the 2D trajectories and headings of tracked objects. The dataset includes complex interactions, such as vehicles yielding to pedestrians and cyclists navigating city streets. Compared to the first version, this collection has longer sequences and more diverse scenarios. Lastly, the Lyft Level 5 dataset [189] comprises over 1,000 hours of data collected by 20 autonomous vehicles along a set route in California over four months. The dataset includes 170,000 scenes, each 25 seconds long, detailing the precise movements of nearby vehicles, cyclists, and pedestrians, and features a high-definition semantic map and aerial view.
2.8 Prediction Paradigms
Figure 6: Overview of prediction paradigms in autonomous driving: (a) Detect-Track-Predict: a modular pipeline where detections are used to initialize and update active tracklets, which are subsequently provided as past trajectories to the prediction module; (b) End-to-End Detection-Prediction: an integrated approach where detections are pooled and directly utilized by the prediction module; (c) Query-Based Prediction: a query-driven framework where predictions are refined based on object queries applied to encoded sensor data.
Prediction models do not exist in isolation; it is an integral part of the autonomous driving stack and must be designed with its role in perception and planning in mind. To ensure feasibility for real-world deployment, prediction models should seamlessly integrate into the broader system architecture. This section discusses three existing prediction paradigms: Detect-Track-Predict, End-to-End Approaches, and Query-Based Approaches, which are graphically illustrated in Figure 6. For each paradigm, we outline the general concept and highlight its advantages and limitations for integration.
2.8.1 Detect-Track-Predict
The Detect-Track-Predict paradigm is a modular cascade approach consisting of detection, tracking, and prediction modules. First, the detection module localizes objects from sensor data. Next, the tracking module initializes tracklets for the detected objects and maintains their trajectories consistently over time. Finally, the prediction module takes the active tracklets and forecasts their future positions. This approach is one of the most common implementations in autonomous driving and is widely used in the literature. Its modular design makes it easier to debug, as most datasets provide ground truth annotations for each of the modules. Additionally, the extensive research output for detectors [33] and trackers [191], covering both 2D and 3D modalities, further supports its seamless development.
The latest detection approaches leverage data fusion from cameras and LiDAR sensors to enhance detection and tracking accuracy as shown in Figure 7. Early Fusion integrates raw sensor data, such as images and point clouds. A common implementation involves detecting objects in the image, projecting the point cloud onto the 2D image plane, and matching the projected points with the detected objects to classify them and measure distances accurately. Guidolini et al. [192] proposed such a pipeline for combining the strengths of both modalities. In contrast, Late Fusion operates at the detection stage, combining results from LiDAR and camera detections. This can be implemented in two ways: (1) merging 3D bounding boxes detected independently from LiDAR and camera, or (2) projecting LiDAR’s 3D bounding boxes onto the 2D image plane and aligning them with the 2D detections from the camera. This strategy provides flexibility, allowing fusion to occur in either the two-dimensional or three-dimensional space, depending on the accuracy of the detection techniques for both modalities. Another fusion strategy, Augmenting Point Clouds, enhances LiDAR data by projecting point clouds onto images and enriching each point with matching image features. This augmented representation is then fed into a LiDAR-only detection method. Vora et al. [193] proposed PointPainting, which projects LiDAR points onto semantic segmentation outputs from the camera and appends the corresponding class scores to each point. Camera-to-LiDAR projections can result in a loss of rich semantic information. Bird’s-eye view (BEV) Fusion mitigates this issue by combining features from different modalities within a unified BEV representation, providing a more comprehensive spatial understanding. For a detailed overview, we refer the reader to the multi-sensor fusion framework presented in [194].
Figure 7: Different fusion approaches for LiDAR and camera data: early fusion, late fusion and BEV fusion.
Kalman Filter-based [195, 196, 197], leverage predicted object states to mitigate detection inaccuracies caused by sensor noise or missed detections. Advanced trackers [198] address partial occlusions, ensuring robust trajectory maintenance and minimizing false positives that lead to ghost tracklets. Prediction models rely on the precision of trajectories over the past 10–15 frames, making robust tracklets essential for accurate forecasting. Excessive ghost tracklets result in non-existent predictions being passed to the planner, potentially degrading overall system performance.
While the Detect-Track-Predict paradigm is modular and intuitive, its sequential nature causes errors to propagate across modules, with detection inaccuracies leading to unreliable tracklets and degraded predictions. Additionally, the modular independence of its components often results in suboptimal performance due to the lack of end-to-end optimization. These limitations motivated the need for alternative paradigms that integrate perception and prediction more cohesively.
2.8.2 End-to-End Models
End-to-End models [170, 199, 200, 201, 100] are purposed to address the limitations of modular approaches by jointly optimizing object detection and trajectory prediction within a single model. This approach assume no explicit tracker and capturing temporal dependencies directly from raw sensor inputs, e.g., sequential LiDAR or camera frames. As shown in Figure 6, a pool module can be integrated to bridge the gap between detection and prediction. The pool aggregates and encodes object features from detection outputs into a unified representation optimized for trajectory prediction. Additionally, it organizes and streamlines inputs to the predictor, dynamically filtering or prioritizing detections based on their relevance to reduce redundant or irrelevant information.
By sharing high-dimensional scene features and utilizing components like the pool, end-to-end models retain richer contextual information, enabling the learning of intermediate representations that benefit both detection and prediction tasks. This paradigm avoids the information bottlenecks of cascade approaches by eliminating the tracking module and jointly optimizing detection and prediction. However, despite these advantages, errors in the shared features from the detection stage can propagate through the system, affecting overall performance.
2.8.3 Query-Based Prediction Models
Query-Based prediction models represent a cutting-edge approach to trajectory prediction by leveraging learnable queries within a unified framework. Rather than splitting detection and prediction into separate stages, query-based models directly focus on objects of interest, iteratively refining their trajectory predictions. These learnable object queries encapsulate spatial and temporal parameters, guiding the model in producing accurate and consistent predictions. By integrating contextual data over multiple refinement steps, query-based models capture fine-grained details and interaction dynamics. A key advantage of this approach is the elimination of error propagation across separate subsystems, as predictions are generated holistically within a single framework. In the following discussion, we highlight several query-based methods proposed in the literature.
Zhou et al. [182] proposed a query-centric paradigm for trajectory prediction, encoding scene information to reuse encodings efficiently. The model employs anchor-free and anchor-based queries to iteratively refine trajectory proposals, enabling parallel multi-agent trajectory decoding. Next, Hu et al. [202] introduced UniAD, a unified framework integrating perception, prediction, and planning with a query-based design. Modules such as TrackFormer, MapFormer, MotionFormer, and OccFormer perform tasks like detection, tracking, road segmentation, future movement prediction, and occupancy forecasting. The final planner predicts waypoints while avoiding collisions and optimizing trajectories, ensuring coordinated decision-making. Its scene-centric approach considers agent interactions and environmental context. Casas et al. [203] presented DeTra, a query-based framework unifying detection and trajectory prediction. Temporal and spatial object queries are used to process LiDAR point clouds and HD maps, iteratively refining object poses. Lastly, Agro et al. [169] proposed a query-based occupancy flow prediction model where query points represent BEV spatio-temporal coordinates. These guide the network to predict occupancy probabilities and motion flow dynamically, integrating past and present spatial information.
Query-based methods offer flexibility and holistic scene understanding but may face challenges in real-time deployment in dense scenarios due to the computational overhead of iterative query refinement. Additionally, these methods depend heavily on accurate scene encoding and high-quality input data, such as HD maps and LiDAR, for reliable forecasting.
3 Survey of Modelling Methods
This section describes the proposed taxonomy of prediction approaches. To systematically categorize a broad range of prediction models from the literature, we structure our taxonomy based on the modeling approach, providing a concise classification and clustering of methods. Selecting an appropriate modeling approach based on application requirements and motion modeling aspects is crucial for designing an efficient, robust, and competitive prediction model comparable to SOTA approaches.
3.1 Taxonomy
In the proposed taxonomy, prediction methods are divided based on the way motion is modeled, i.e., physics-based methods explicitly define motion equations, while learning-based approaches learn motion patterns from observations. Learning-based approaches are further subdivided into classical machine learning, deep learning, and reinforcement learning-based methods (please refer to Figure 8). Each category of methods is elaborated upon below.
  1. 1.
    Physics-based methods predict motion by simulating a set of equations describing a physics-based model. Single-Trajectory methods predict trajectory by directly applying dynamic or kinematic equations to the agent’s current state. Kalman Filtering methods account for the agent’s state uncertainty, which increases during prediction steps and reduces after receiving observations. Monte Carlo methods approximate the state distribution by randomly sampling from the input variables and applying a physics model to generate potential future trajectories. Multiple Model methods account for motion-mode uncertainty and combine several motion models to better describe the complexity of the agent’s motion behavior.
  2. 2.
    Learning-based methods, in contrast to physics-based approaches, do not exploit explicitly defined functions governing the motion but instead learn motion dynamics from driving data.
    1. (a)
      Classical Machine Learning (ML) methods learn motion dynamics from driving data by fitting function approximators, e.g., Markov Models or Gaussian Processes. In this paper, we consider four main groups of methods: Gaussian Processes, Hidden Markov Models, Dynamic Bayesian Networks, and Gaussian Mixture Models.
    2. (b)
      Deep Learning (DL) methods also learn motion dynamics from driving data but are more suitable for long-term predictions in complex environments than conventional ML models, due to their capability of learning complex and hierarchical data representations. Among DL methods, the following main groups of models are reviewed: Sequential Models, Convolutional Neural Networks, Graph Neural Networks, Transformers, AutoEncoders, and Generative Adversarial Networks.
    3. (c)
      Reinforcement Learning (RL) methods interpret driving as a high-dimensional complex policy. Despite these methods having direct applications for planning, recent works increasingly utilize RL methods for trajectory prediction through learning via a reward system. Inverse Reinforcement Learning methods learn a reward function based on expert demonstrations. Imitation Learning bypasses the stage of learning reward functions and directly extracts an optimal policy from the data, generating future trajectories similar to expert trajectories.
Figure 8: Overview of the Taxonomy of modelling methods.
3.2 Physics-based Methods
Physics-based methods accomplish motion prediction by simulating a set of dynamics equations describing a physics-based model. These methods apply the laws of physics to estimate steering and acceleration, predicting position, orientation, and speed. To describe vehicle movement, they employ kinematic [204] and dynamic models. Kinematic models, which do not account for the forces governing motion, are widely used due to their simplicity. Commonly used kinematic models include the Constant-Velocity (CV) model, which estimates position and velocity under the assumption of constant velocity; the Constant-Acceleration (CA) model, which estimates position, velocity, and acceleration assuming constant acceleration; and the Constant Turn Rate and Velocity (CTRV) and Constant Turn Rate and Acceleration (CTRA) models, which estimate position and orientation under the assumption of constant yaw rate. Dynamic models, on the other hand, account for forces impacting motion, introducing significant complexity when modeling components such as wheels, engines, and gearboxes. However, the added complexity in dynamic models often results in only marginal improvements in prediction accuracy. Furthermore, the forces governing the motion of other vehicles are not observable through sensors. These limitations make dynamic models less preferable than kinematic models, with a tendency toward simpler dynamic models for practical applications.
The simplest method to predict a trajectory is to directly apply the vehicle’s current state to dynamic [205, 206, 207, 208] or kinematic models [209, 210]. This method assumes no noise in the vehicle’s state estimation, making it suitable for short-term prediction. However, as it does not account for state uncertainty, it becomes unreliable for long-term motion prediction. Kalman Filtering (KF) Methods address state uncertainty by representing the vehicle state with a Gaussian distribution. The state includes position $(x,y)$ and velocity $(v_{x},v_{y})$, and its prediction is computed from the current state using a state transition matrix defined by dynamic or kinematic models. The filter updates the covariance matrix of the state estimation error based on system dynamics, predicts the expected measurement, and computes the error between predicted and actual measurements. The Kalman gain is then calculated to update the state estimate and reduce uncertainty. Variants of Kalman filters, such as the linear Kalman Filter [211], the Extended Kalman Filter (EKF) [212], and the Unscented Kalman Filter (UKF) [213], have been applied to non-linear systems in trajectory prediction.
While KF methods assume Gaussian distributions, Monte Carlo Methods approximate state distributions without such assumptions. These methods randomly sample input variables (e.g., initial conditions, sensor noise) and apply a physics model to generate potential future trajectories [214]. Monte Carlo methods are effective in estimating uncertainty but are computationally intensive and require a well-defined set of parameters and constraints to generate feasible results. Several works extend physics-based methods by embedding map information and incorporating road constraints. Petrich et al. [215] proposed a stochastic filter that uses digital map information to select a set of reasonable trajectories by projecting detected vehicles onto the centerlines of map lanes. Rehder and Kloeden [216] mapped environments to location maps representing pedestrian locations from occupancy grids of sidewalks and obstacles, converting the prediction problem into a planning task by modelling pedestrian destinations as latent variables. Coscia et al. [93] used polar grids centered at predicted agents to construct point-wise circular distributions, combining them into statistically smoothed predictions. Next, Koschi et al. [217] incorporated pedestrian dynamics, traffic rules, road connectivity, and semantic regions to obtain reachable pedestrian occupancy. Batkovic et al. [218] employed graph representations of road geometry to predict pedestrian motion, assuming motion along graph edges representing sidewalks and crosswalks, with a unicycle model ensuring trajectory adherence to edges.
In real traffic, agent motion behavior is complex and cannot be described by a single model, often involving linear movements, turns, and acceleration/deceleration modes. Since motion modes of other vehicles are not directly observable, Multiple-Model (MM) methods address this uncertainty by using a set of possible motion models. These methods employ a bank of elemental filters, each based on a model from the set, to represent complex motion. At each time step, candidate models are evaluated, and the filters generate estimates, which can be combined or fused. For further details on handling discrete and continuous-based uncertainties, the review on MM methods [219] provides comprehensive insights. The Interacting Multiple Model (IMM) estimator [220, 221] is a hybrid filter with nearly linear computational complexity, offering a balance between complexity and performance. Kaempchen et al. [222] proposed an IMM approach utilizing constant speed, constant acceleration, and acceleration change models to compute vehicle states as weighted state estimates. Other IMM designs accounting for road constraints are presented in [223, 224]. Recent works, such as Interacting Multiple Model Kalman Filter (IMM-KF) [225], employs intention-based models and optimization-based projections for generating non-colliding predictions using a hierarchical approach for multiple vehicles. Kuhnt et al. [226] proposed a Unified Traffic Situation Estimation Model using IMM with constant and target velocity models, incorporating predefined environmental geometry to estimate future trajectories.
Figure 9: ML approaches for trajectory prediction: (a) - GP modelling observed and predicted trajectories showing uncertainty bounds; (b) - HMM representing latent states and their temporal transitions, with associated observations.
Figure 10: Main DL approaches presented schematically in application to prediction where the input is past trajectory: (a) - RNN/LSTM models, (b) - convolution models, (c) - transformers, (d) - GNNs, (e) - autoencoders and (f) - GANs.
Physics-based models are efficient for short-term predictions, but for long-term predictions, hybrid models [227, 228, 229, 230] have been developed by combining physics-based and learning-based approaches. Kim et al. [227] proposed a hybrid method that uses the Reciprocal Velocity Obstacles (RVO) framework to predict pedestrian trajectories in dynamic environments, enhanced with Ensemble Kalman Filtering (EnKF) and Expectation-Maximization (EM) for real-time learning and parameter refinement. Ma et al. [228] introduced a deep latent variable model built on a Wasserstein auto-encoder, enabling the automatic satisfaction of kinematic constraints. Yue et al. [229] presented Neural Social Physics (NSP), a Neural Differential Equation model that integrates an explicit physics-based framework as an inductive bias, paired with deep neural networks for system parameter estimation and stochastic dynamics modelling.
3.3 Machine Learning-based Methods
Machine Learning-based methods learn motion dynamics from driving data by fitting function approximators. Figure 9 provides an overview of two key ML techniques discussed in this section: Gaussian Processes (GPs) and Hidden Markov Models (HMMs). Additionally, the section discusses prediction models based on Gaussian Mixture Models (GMMs), which combine multiple GPs, and Dynamic Bayesian Networks (DBNs), which effectively capture interactions between agents. Each method is discussed in dedicated subsections, elaborating on the theoretical foundations, their application to prediction tasks, and relevant studies available in the literature.
3.3.1 Gaussian Process
Gaussian Process (GP)-based prediction models [231, 232] treat each position at discrete time steps as a random variable governed by a probabilistic prior. To capture the structure of the trajectory and assumptions about smoothness, the GP uses a kernel function to measure the similarity between inputs. This similarity translates into how correlated the outputs (e.g., predicted positions) are, ensuring that similar inputs result in similar outputs. The kernel also defines the covariance matrix, which connects observed inputs and outputs to predictions for unobserved points. During training, the kernel hyperparameters are optimized by maximizing the likelihood of the observed data, ensuring the GP accurately reflects the underlying motion dynamics. Predictions for future positions are made by conditioning the prior distribution on observed data, yielding a posterior distribution with a mean trajectory (future trajectory) and variance (uncertainty) for each future time step.
Mínguez et al. [233] proposed a model based on Balanced Gaussian Process Dynamical Models (B-GPDMs) for pedestrian path and intention prediction, reducing high-dimensional motion data into a latent space to enable smooth trajectory modelling. Guo et al. [234] utilized a GP velocity field to represent each motion pattern, mapping positions to velocities in a region of interest. The model integrates a Dirichlet Process (DP) prior to avoid assumptions about the number of motion patterns, allowing it to learn them directly from data. Lastly, Ferguson et al. [235] employed Gaussian Process Mixture Models to detect changes in pedestrians’ intent and learn new motion patterns online, enabling the handling of previously unobserved scenarios.
3.3.2 Hidden Markov Models
Prediction models based on Hidden Markov Models (HMMs) [154, 236, 237, 238, 239, 240] represent trajectories as sequences of hidden states evolving over time. Figure 9 (b) illustrates HMM-based trajectory prediction, where hidden states $q_{t},q_{t+1},...,q_{t+\Delta t}$} represent latent factors such as intent or motion patterns, linked by transition probabilities $a_{ij}$ that model temporal dependencies. The observed trajectory $u_{t-k},...,u_{t-1}$} serves as input to infer hidden states using decoding algorithms. These states generate the predicted trajectory $u_{t},u_{t+1},...,u_{t+\Delta t}$} via emission probabilities $b_{i}(u)$, which define the likelihood of observations given the hidden states. By learning transition and emission probabilities from data, HMMs capture sequential dependencies and enable robust trajectory prediction.
HMM-based models are particularly effective in scenarios where the agent’s behavior is governed by discrete modes or intentions. Rathore et al. [241] proposed a hybrid framework, Traj-clusiVAT, combining trajectory clustering and Markov chain modelling. The clustering algorithm partitions trajectories into representative clusters, and Markov models are trained for prediction. Li et al. [242] proposed an enhanced HMM for behavior prediction, with key improvements including random initial parameter generation to ensure reliability, parameter learning through the Baum-Welch algorithm, and prediction via the Viterbi algorithm combined with Bayesian decision theory. On Continuous HMMs, Ren et al. [243] proposed a framework for behavior recognition. The CHMM is employed for early detection of lane-changing behavior by analyzing vehicle motion patterns and contextual traffic features.
3.3.3 Gaussian Mixture Models
Applying Gaussian Mixture Models (GMMs) for trajectory prediction leverages their ability to capture the multimodal nature of future trajectories. The GMM represents the distribution of possible future paths as a weighted sum of Gaussian components:
$p(x)=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(x|\mu_{k},\Sigma_{k}),$
where $\pi_{k}$ is the weight of the $k$-th Gaussian component, $\mu_{k}$ represents the mean, and $\Sigma_{k}$ is the covariance. The mean $\mu_{k}$ corresponds to the central point or the most likely position of a trajectory predicted by the $k$-th Gaussian component. The covariance $\Sigma_{k}$ quantifies the spread and uncertainty of predictions around this mean, capturing both variance along each axis and correlations between them. Together, these parameters enable the model to represent multiple plausible future paths while accounting for prediction uncertainty.
Wiest et al. [244] applied Chebyshev polynomial decomposition to represent trajectories uniformly and used GMMs to provide multimodal predictions by estimating conditional densities. Combining HMM and GMMs, Deo et al. [154] proposed a holistic motion prediction framework that integrates HMM for maneuver recognition, Variational GMM (VGMM) for trajectory prediction, and a Vehicle Interaction Module (VIM) for inter-vehicle interactions. VGMM models the conditional probability of future trajectories using a variational Bayesian approach, improving robustness against overfitting and outliers. In more recent approaches, Mebrahtu et al. [245] utilize a Transformer architecture to extract sequential dependencies from observed pedestrian trajectories, coupled with a Mixture Density Network (MDN) to estimate GMM parameters.
Table 3: Summary of reviewed DL-based Models relying on RNN, GRU and LSTM (part 1)
Reference Name Description Alahi et al. [116] (2016) Social LSTM Individual LSTMs model each pedestrian’s motion and are connected in a social pooling sharing hidden state information with neighbors. Al-Molegi et al. [246] (2016) STF-RNN Space Time Features-based RNN, extracts spatial data from GPS logs and temporal information from timestamps. Zheng et al. [247] (2016) - Hierarchical policy network (HPN), integrates macro-goals for long-term trajectory planning and micro-actions using attention mechanisms. Jain et al. [248] (2016) Structural-RNN Transforms spatio-temporal graphs into a structured RNN-based architecture by employing RNNs for nodes and edges. Bhattacharyya et al. [51] (2017) LSTM-Bayesian First RNN stream with CNN-encoder for odometry prediction, second RNN stream with Bayesian framework to model uncertainty and condition on odometry. Xue et al. [249] (2017) - Uses a bidirectional LSTM to classify trajectories into route classes. Kim et al. [166] (2017) Occupancy-LSTM Individual LSTMs for each surrounding vehicle, generating the output as an occupancy grid map. Varshneya and Srinivasaraghavan [250] (2017) SSCN-LSTM LSTM for short-term dependencies, Spatially Static Context Network (SSCN) for spatial context and polling mechanism for interactions. Bock et al. [84] (2017) Continuous-LSTM Continuous learning with LSTM model, uses Kalman Filter for prediction if new data is not sufficiently covered in dataset. Saleh et al. [85] (2017) - Uses stacked LSTMs for intention prediction. Zyner et al. [17] (2017) - LSTM for driver intention prediction trained on position, heading, and velocity. Phillips et al. [21] (2017) - LSTM for intention prediction at intersection trained on base, history, traffic and rule features. Zyner et al. [19] (2018) - Combines recurrent model with a mixture density network. Deo and Trivedi [251] (2018) - Encodes past trajectories into a context vector, combines it with representations of lateral and longitudinal maneuvers, and decodes into maneuver specific trajectories and probabilities of maneuvers. Zyner et al. [18] (2018) - A single fully connected layer followed by three recurrent layers. Xu et al. [119] (2018) CIDNN Combines LSTMs for individual motion and a multi-layer perceptron for transforming locations into high-dimensional feature space to further predict displacement. Saleh et al. [252] (2018) - Stacked recurrent neural networks for cyclist trajectory prediction. Xue et al. [253] (2018) SS-LSTM Uses three LSTMs: to capture individual information, social and scene scale. Manh and Alaghband [254] (2018) Scene-LSTM Two coupled LSTMs trained simultaneously: first group pedestrian LSTMs modelling movement of agents and second group Scene-LSTMS one per grid-cell. Hasan et al. [255] (2018) MX-LSTM Merges tracklets with short sequences of head pose estimations using a joint optimization. Xin et al. [256] (2018) - Contains two LSTM blocks: one block to decode driver intentions and second block for predicting. Bisagno et al. [257] (2018) Group-LSTM Splits pedestrians into groups by clustering similar motion patterns and social-LSTM for prediction. Bartoli et al. [258] (2018) - Social LSTM with context-aware pooling layer for static objects in the neighborhood of a person. Ivanovic and Pavone [259] (2018) Trajectron Combines recurrent and graph-based models to capture individual behavior and interactions. Pfeiffer et al. [260] (2018) Static-LSTM Encodes pedestrian using 1D grid; accounts for static obstacles and influences of other pedestrians. Huynh and Alaghband [261] (2019) - Combines Scene-LSTM for capturing traveled paths and Pedestrian-LSTM for individual prediction. Huang et al. [262] (2019) STGAT Combines a graph attention for capturing spatial interactions and LSTM for temporal interactions.
3.3.4 Dynamic Bayesian Networks
Dynamic Bayesian Networks (DBNs), applied to trajectory prediction [263, 264], capture temporal dependencies and interactions between multiple agents. A typical DBN architecture comprises three layers: the behavior layer, representing input data for each agent, such as past trajectories, velocities, interaction features, and scene context; the hidden layer, modeling latent variables that capture unobserved, intermediate states influencing motion and their probabilistic transitions over time; and the observation layer, which outputs the predicted future trajectories, incorporating the uncertainty of the prediction process. During training, the forward-backward algorithm estimates the most likely sequence of hidden states and their influence on future observations.
On maneuver prediction, Schreier et al. [265] model driving maneuvers as latent states in a DBN and use probabilistic inference to predict maneuvers, which are subsequently used by maneuver-specific models to generate future trajectories. Li et al. [266] focus on vehicle maneuver prediction in highway scenarios, leveraging multi-dimensional predictive features, including historical vehicle states, road structures, and traffic interactions, to infer maneuver probabilities. Similarly, He et al. [267] propose a framework to predict vehicle intentions in highway driving scenarios using DBN. The approach incorporates contextual inputs, such as relative speed, safety distance, and lane information, to infer driver intentions like lane-keeping or lane-changing. The DBN’s observation layer captures lane deviation and heading angle, which are influenced by the inferred intentions.
Further, focusing on interaction modelling, Liu et al. [268] employ DBN to design a Driver Characteristic and Intention Estimation (DCIE) module. This probabilistic framework estimates driver characteristics, e.g., stable or aggressive behavior, and intentions, by integrating lateral position, velocity, and acceleration. Lastly, Schulz et al. [269] leverage DBN for predicting future trajectories and interactions, comprising three layers: the route intention layer to determine the desired path of each agent based on the road layout and traffic rules; the maneuver intention layer; and the action layer, capturing continuous actions based on current intentions and interactions.
3.4 Deep Learning-based Methods
Deep Learning-based methods excel in learning motion dynamics from driving data and are particularly effective for long-term predictions in complex environments. Figure 10 illustrates the main DL models reviewed in this section. Each approach is detailed in a separate subsection, with methods summarized in tables that highlight the key features of the architectures.
3.4.1 Sequential Models
Table 4: Summary of reviewed DL-based Models relying on RNN, GRU and LSTM (part 2)
Reference Name Description Lisotto et al. [270] (2019) SNS-LSTM Combines three factors using pooling mechanisms: past trajectories patterns, interpersonal interactions, and environment’s semantics. Syed and Morris [271] (2019) SSeg-LSTM Uses SegNet for semantic segmentation and extracting scene features, which are further integrated into SS-LSTM. Chandra et al. [272] (2019) graph-LSTM Two-stream framework: first stream - LSTM encoder-decoder to predict spatial trajectories, second stream predicts eigenvectors of dynamic geometric graphs (DGGs) to infer agent behaviors. Zhang et al. [273] (2019) SR-LSTM Includes intentions of neighbors through message-passing and refines the states of all pedestrians. Chandra et al. [274] (2019) TraPHic A hybrid network, which combines LSTM and CNN for predictions, focusing on the varied interactions among different road users. Min et al. [275] (2019) - Focus on forecasting short and long-term future trajectories, employs ensemble mechanism to enhance robustness. Mangalam et al. [276] (2019) - Predicts by disentangling global (rigid body) using Quasi-Recurrent Neural Networks (QRNNs) and local (joint-specific) motions using Encoder-Recurrent-Decoder framework. Dai et al. [22] (2019) ST-LSTM Integrates spatial interactions within LSTM. Ding et al. [26] (2019) - Uses LSTM to predict vehicle’s driving policy to guide a low-level optimization process. Ma et al. [176] (2019) TrafficPredict Instance layer for individual movements and interactions and category level to capture similarities in a group motion. Bhujel and et al. [277] (2019) - Combines RNN and attention mechanism to capture human-human and human-scene interactions. Zhang et al. [88] (2020) AT-LSTM Inferred a set of variables impacting a crossing intention and trained LSTM with attention mechanism modeling these parameters. Tao et al. [278] (2020) - Two components: Context Module (ICM) for temporal dependencies and Social-aware Context Module (SCM) for spatial interactions. de Brito et al. [279] (2020) Socila-VRNN Feature extraction module for agent states, social interactions, and environmental context, followed by a probabilistic inference module. Lorenzo et al. [280] (2020) - Recurrent model with additional features as pedestrians direction orientation and gaze direction. Wu et al. [163] (2020) - Includes pedestrian intention and behavior to generate predictions. Bertugli et al. [281] (2021) AC-VRNN Integrates GCN and KL-divergence loss with belief maps to capture interaction and context. Li et al. [282] (2020) RAI Contains a temporal attention module to mitigate speed variations and spatial pooling to capture social dynamics and intentions. Xing et al. [283] (2020) - Classifies driving styles using GMM and uses a shared LSTM layer for each driving style. Zhang et al. [284] (2021) - Uses a separate LSTM model for intention prediction and integrates intersection prior trajectories built from analyzing traffic flow. Chiara et al. [285] (2022) - Has a goal-estimation module using a U-Net architecture to predict scene-aware trajectories. Zhang et al. [286] (2023) STS-LSTM Spatial-Temporal-Spectral LSTM uses spatial, temporal, and spectral domain features and contains a transferable model for pedestrian trajectory prediction. Xu et al. [287] (2023) GC-VRNN Integrates a Multi-Space GNN for spatial feature extraction and RNN with a Temporal Decay (TD) module for temporal dependencies and handling incomplete data.
Sequential models operate under the assumption that motion can be described by temporally causal dependencies. To capture complex long-term temporal relationships, recurrent neural networks (RNNs) are widely used for trajectory prediction. RNNs and their variants effectively capture the dynamic properties of motion and the temporal correlations in prediction sequences by incorporating information from previous timestamps and using hidden states to generate future trajectory sequences. Focusing on model design, earlier studies primarily utilized a single recurrent model to map past trajectories to future paths [288, 97, 102]. However, more recent works employ multiple groups of recurrent models to address different aspects of motion [289, 256, 22], as well as stacked LSTMs and convolutional LSTMs. Dai et al. [22] proposed a model, where one group of LSTMs models individual vehicle motions, while another group captures interactions between agents. The architectural approach for stacked LSTMs involves hierarchically organizing recurrent networks into layers, as seen in [290, 251]. Tables 3 and 4 provide a summary of architectures in the literature that rely on recurrent models.
Varshneya and Srinivasaraghavan [250] combined LSTM with an attention mechanism to simultaneously preserve dynamic context and learn patterns from the spatial coordinates of agents. To capture spatial context, the model utilizes a Spatially Static Context Network (SSCN). Ding and Shen [23] introduced a prediction system with a dual-level design. The first level is a policy anticipation network, which employs an RNN-based encoder to analyze past trajectories and predict drivers’ intentions. These predicted intentions are passed to the second level, where a context reasoning module incorporates detailed contextual data. The framework uses non-linear optimization to generate continuous predictions based on structured contextual information. For multimodal predictions, Zyner et al. [19] utilize an encoder-decoder three-layer LSTM to predict parameters of a weighted Gaussian Mixture Model (GMM) to generate trajectories. Park et al. [20] also use encoder-decoder LSTMs, focusing on predicting the probability of occupancy on a grid map. A beam search is then employed to select the $k$ most probable future trajectories from the candidates. Another approach is presented in [251], where six distinct decoder LSTMs are used, each corresponding to a specific highway driving maneuver. An encoder LSTM processes past trajectories, and the hidden state of each decoder LSTM is initialized with a combination of the encoder LSTM’s last hidden state and a one-hot vector representing the maneuver specific to that decoder. These decoder LSTMs predict parameters for a maneuver-conditioned bi-variate Gaussian distribution of future vehicle locations.
Table 5: Summary of reviewed DL-based Models relying on CNNs
Reference Name Description Yi et al. [291] (2016) Behavior CNN Integrates formation of stationary crowd groups and analyzes their interaction with moving pedestrians to generate predictions. Varshneya and Srinivasaraghavan [250] (2017) SSCN Models static artifacts to improve predictions and uses a soft attention mechanism to learn motion patterns of humans using different navigational modes. Lee et al. [60] (2017) - Uses CNN and predictive controller to predict a lane-change intention from frontal camera image. Luo et al. [40] (2018) - Uses 3D convolutions across space and time on LiDAR data for predictions. Casas et al. [292] (2018) - One-stage detector and forecaster from 3D point cloud and dynamic environmental maps. Rehder et al. [81] (2018) RMDN Destination network uses CNN to predict a mixture of possible destinations, topology network uses FCN to predict planning topology for generating predictions. Nikhil and Morris [86] (2019) CNN-TP CNN framework with multiple convolution layers run in parallel. Jain et al. [293] (2019) DRF-Net CNN for long-term human motion prediction, by capturing a wide range of potential future movements. Cui et al. [61] (2019) - Uses CNN to encode contextual surroundings of each agent from raster images. Du et al. [294] (2019) - Two-stream architecture: spatial stream at frame level and temporal stream across frames. Applies TVNet model to generate optical-flow-like features for each frame-pair. Krüger et al. [295] (2020) - Uses potential field method to model navigation of each agent and passes potential field representation to neural network as input. Mo et al. [296] (2020) - Interaction-Aware Trajectory Prediction of Connected Vehicles using CNN-LSTM Networks. Fang et al. [297] (2020) TPNet Employs a two-stage process: first, generates trajectory proposals using polynomial curve fitting; second, refines them, incorporating knowledge from semantic maps and prior physical constraints. Phan-Minh et al. [69] (2020) CoverNet Treats trajectory prediction problem as a classification task over a curated set of possible trajectories. Chou et al. [298] (2020) - Transforms HD maps and agents’ surrounding context into BEV and passes to CNN as the input. Strohbeck et al. [299] (2020) - CNN for raster features and TCN for states and motion history. Marchetti et al. [173] (2020) MANTRA Uses LSTM for encoding, Memory Augmented NN to decode future trajectories and CNN for refining. Gilles et al. [300] (2021) HOME Uses CNN to encode local context from rasterized image of the agent’s environment. Ye et al. [301] (2021) TPCN Contains dual-representation spatial learning module and dynamic temporal learning modules for spatial and temporal features. Wang et al. [302] (2021) MI-CNN Uses multi-input: past trajectories, pose, depth maps, and 2D-3D size data to generate predictions. Zhang et al. [303] (2021) Social-IWSTCNN Social Interaction-Weighted Spatio-Temporal CNN with a Social Interaction Extractor for learning social interaction weights and spatial features. Zamboni et al. [304] (2022) Conv2D Uses CNN with data augmentation. Chen et al. [305] (2024) DSTCNN Deformable Spatial-Temporal CNN models spatial and temporal interactions and uses adaptive kernels.
3.4.2 CNNs
Table 6: Summary of reviewed DL-based Models relying on Autoencoders
Type Reference Description Encoder- Decoder Park et al. [20] (2018) LSTM-based encoder-decoder architecture. Deo and Trivedi [251] (2018) LSTM-based encoder-decoder with maneuver classification branch. Hong et al. [306] (2019) Encodes scene context using a unified spatial grid representation. Chen et al. [307] (2022) End-to-end CNN-LSTM encoder-decoder with an attention mechanism, leveraging semantic segmentation to extract obstacle information. Hui et al. [308] (2022) DNN-based Encoder-Decoder Park et al. [309] (2024) Masked Autoencoder, leveraging masked reconstruction of input features to adapt to distribution shifts. Lv et al. [310] (2024) Learning Autoencoder Diffuser modelling group relationships for pedestrian trajectory prediction. Variational Autoencoder Lee et al. [64] (2017) CVAE-based RNN encoder-decoder for samples generation and ranking the most likely samples. Hu et al. [311] (2018) Two modules: module predicting intentions and module for predicting motion within the interacting scene. Cho et al. [312] (2019) CVAE-based interaction module. Feng et al. [313] (2019) Uses GRU-based encoder-decoder and integrates intention prediction module. Hu et al. [314] (2019) Samples joint trajectories of interacting vehicles, converts joint distribution to conditional + MPC. Hu et al. [315] (2019) Conditioned on agent’s intention and historical scene information. Fei et al. [316] (2020) Uses Wasserstein distance for loss and generates predictions using a GMM. Ivanovic et al. [317] (2020) Encodes interaction history and target agent intentions. Zhang et al. [318] (2020) Stacked auto-encoders to encode state; LSTM encoder for past trajectory and CNN pooling for interactions. Cheng et al. [319] (2020) Integrates a self-attention with a two-stream encoder to encode observed trajectories and agent interactions. Lee et al. [320] (2022) A cascade of Conditional VAEs: macro-stage predicts high-level goals and micro-stage refines predictions into sequential paths. Xu et al. [321] (2022) Uses a backward RNN to capture navigation patterns and attention mechanism for social interaction. Zhong et al. [322] (2022) Employs CVAE with a designed modal-wise sampling strategy to improve prediction stability. Chen et al. [323] (2022) Uses graph partitioning to form cliques of interacting agents to reduce latent space complexity. Zhou et al. [324] (2023) Cascaded CVAE module using past trajectories and earlier predicted locations with a socially-aware regression module for refinement. Qiu et al. [325] (2023) VAE with a diffusion prior. Xu et al. [326] (2023) Uses a dual-attention mechanism to encode environmental and social contexts, generating map-compliant and socially-aware predictions. Xiang et al. [327] (2024) Uses a socially explainable energy map to model social interactions.
Convolution prediction models are categorized into two groups in the literature: those using image-based input and Temporal Convolution Networks (TCNs). Image-based inputs can come from various perspectives, including BEV, vehicle-mounted cameras, or other sensor modalities. BEV images or aerial photographs offer a top-down perspective that effectively captures spatial relationships and road structures. CNN-based models [69, 61, 298] primarily focus on the feature extraction stage to capture road geometries, traffic agent positions, and environmental context. The second category involves Temporal Convolution Networks (TCNs), which apply convolution layers directly to trajectory sequences to model temporal dependencies. Unlike RNNs, TCNs operate in parallel over the entire sequence, incorporating causal convolutions to preserve temporal order and dilated convolutions to extend the receptive field, capturing long-term dependencies. TCNs are designed to maintain stable gradients, avoid information leakage, and achieve computational efficiency for long sequences. One example is the framework proposed by Nikhil and Morris [86], which preserves spatio-temporal consistency through multiple convolution layers while leveraging parallelism for efficiency. Table 5 provides a summary of CNN-based architectures for trajectory prediction.
Marchetti et al. [173] introduced a multimodal prediction model utilizing a Memory Augmented Neural Network. This model encodes previously observed trajectories as keys to access probable future trajectory encodings from memory, which are then decoded into multimodal predictions conditioned on observed data. The surrounding context is processed by a CNN and passed through a refinement module to enhance the decoded predictions. In a separate study, Gilles et al. [300] developed a prediction model that takes a rasterized image of the agent’s environment as input and generates a heatmap indicating potential future positions with associated probabilities. A standard CNN encodes the local context from a rasterized HD map, while past trajectories are encoded using a 1D convolutional layer followed by a recurrent layer. Interactions are modeled using an attention mechanism. Strohbeck et al. [299] designed a model for trajectory prediction from rasterized images, employing a CNN to extract features that capture both static and dynamic elements. These features are merged with current and past states processed through a TCN. The combined data is then mapped to future states through a series of linear layers. Lastly, Ye et al. [301] proposed a Temporal Point Cloud Network (TPCN) for trajectory prediction, featuring two main blocks: a dual-representation spatial learning module and a dynamic temporal learning module. The spatial module consists of two sub-modules: one for learning point-wise features to capture geometric information and interactions, and another for learning voxel features to extract semantic context. These modules are trained jointly.
A series of studies have combined LSTM’s ability to capture temporal dependencies with CNN’s spatial feature extraction to address spatio-temporal patterns in an integrated manner. Convolutional LSTM (ConvLSTM) models have emerged as a unified solution, incorporating convolutional operations directly within recurrent units to enable simultaneous spatial and temporal processing. ConvLSTM-based models [274, 328, 329, 330, 331, 332, 333, 334] are designed to handle tasks requiring spatio-temporal coherence. Chandra et al. [274] proposed TraPHic, which uses convolutional operations for localized spatial interactions and a horizon-based weighting mechanism to rank interactions by proximity. Khakzar et al. [328] utilized ConvLSTM to model spatio-temporal motion patterns and inter-vehicle interactions, processing risk maps based on metrics like Time-to-Collision (TTC). Mukherjee et al. [329] introduced a three-module framework comprising interaction learning for surrounding vehicle effects, temporal learning for past movements, and motion learning for future position predictions.
Table 7: Summary of reviewed DL-based Models relying on GANs
Reference Name Description Fernando et al. [335] (2018) GD-GAN GAN pipeline for trajectory forecasting and group detection. $G$ generates trajectories conditioned on the local neighborhood context. Gupta et al. [122] (2018) Social GAN $G$ consists of LSTM encoder, social pooling and LSTM decoder. $D$ consists of LSTM encoder and MLP for classification. Amirian et al. [121] (2019) Social Ways Uses attention social pooling for modelling interactions. Roy et al. [336] (2019) Interaction-GAN Embeds interactions as social context; tested on dataset with inconsistent lane behavior at intersections. Li et al. [337] (2019) Prob-GAN Both $G$ and $D$ are combination of one recurrent layer and several fully connected layers. Sadeghian et al. [120] (2019) SoPhie LSTM-based GAN takes as the input features from attention module to generate realistic future paths complying with learnt constraints. Zhao et al. [63] (2019) MATF $G$ is conditioned on the static scene context and agents past trajectory. Li et al. [338] (2019) CGNS Architecture consists of a deep feature extractor, conditional encoder to learn latent space, $G$ to sample trajectory hypotheses from the learnt distribution and recurrent $D$. Lai et al. [339] (2020) AEE-GAN Consists of three blocks: Feature Encoder, Visual and Social Attention, and LSTM-based GAN. The attended ecology visual and social embedding is passed to $G$. Hegde et al. [340] (2020) Vehicle-GAN $G$ consists of LSTM-based encoder and decoder, and pooling layer to model interactions. $D$ employs LSTM-based encoder and MLP. Yang et al. [341] (2020) TPPO Includes latent variable predictor encoding environmental context and motion patterns, and Social Attention Pooling module. Huang et al. [342] (2021) STI-GAN Similar architecture to Social GAN. Employs graph attention model instead of pooling layer. Choi et al. [343] (2021) DSA-GAN Contains a driving style recognition network. $G$ generates trajectories conditioned on driving style. Dendorfer et al. [344] (2021) MG-GAN Multi-Generator GAN, each generator specializes in a distinct trajectory mode + Path Mode Network (PM-Net) to predict a categorical probability distribution over these generators based on the scene context. Wang et al. [345] (2022) TS-GAN $G$ consists of encoder and LSTM-based decoder. Encoder includes Social Recurrent module and Social Convolution block. Li et al. [346] (2021) STL-GAN Employs signal temporal logic (STL) to encode rules. Rule features extracted from STL syntax tree are incorporated into $D$. Fang et al. [347] (2022) Atten-GAN $G$ contains an attention pooling module to allocate influence weights among pedestrians and extract interaction information. Jiang et al. [348] (2023) Two-stage GAN $G$ uses LSTM and graph attention networks (GAT) to estimate cost of transitioning between road network nodes and predict next movements, ensuring spatial and temporal coherence. Yang et al. [349] (2024) SAGAN Social self-attention mechanism to refine interaction information by focusing on most relevant agents.
3.4.3 Autoencoders
In trajectory prediction, autoencoders encode past trajectories into compact latent representations, which are then decoded to predict future positions. These models excel at capturing complex temporal dependencies and can function as standalone predictors or feature extractors integrated into larger architectures. Variational Autoencoders (VAEs) extend this approach by modeling uncertainty, generating a latent distribution instead of fixed representations. By sampling from this distribution, VAEs generate diverse and multimodal future trajectories, modelling the inherent uncertainty in trajectory forecasting tasks. Table 6 provides a summary of autoencoder architectures available in the literature, grouped by deterministic or stochastic type.
Deo and Trivedi [251] introduced an LSTM-based encoder-decoder architecture with a maneuver classification branch. The encoder processes the historical trajectories of the target vehicle and its neighbors, condensing this information into a context vector. This vector, combined with maneuver encodings, is further utilized by the decoder to produce maneuver-specific probabilistic distributions of future trajectories. Hong et al. [306] designed a model to map a world representation to the future trajectories of a single target agent. For the surrounding world context, the authors created a 4D tensor comprising the target agent’s state, its surrounding agents’ states, dynamic context, and road network. This tensor is encoded into a latent representation, which the decoder uses to model future state distributions at specified time offsets, enhanced by temporal convolutions.
Among VAEs, Cho et al. [312] proposed a CVAE-based interaction module where past and future trajectories are encoded to generate a stochastic latent variable $z$. The ego vehicle’s state and its neighbors’ states are used to infer rule information, integrating encoded states with $z$ to generate future trajectories. Hu et al. [314] proposed a generic prediction framework to address cases of irrational human behaviour. First, the model samples joint future trajectories of interacting vehicles based on past trajectories. Next, the joint distribution is converted to a conditional distribution, and Model Predictive Control (MPC) is employed to generate the most probable trajectories. Finally, the weights of generated trajectories are re-weighted to align with rational behaviours. Lastly, Zhang et al. [318] proposed a vehicle-descriptor-based LSTM model to capture interactions between the global context and agents’ local position resolutions. The model encodes the agent’s multimodal state information using stacked autoencoders and employs an LSTM-based decoder to generate trajectory distributions based on estimated maneuvers.
Table 8: Summary of reviewed DL-based Models relying on GNN (part 1)
Reference Name Description Lee et al. [350] (2019) - RNN Encoder, GNN Layers to process pairwise features, and to refine node embeddings through edge-type-specific message passing, RNN Decoder. Sun et al. [351] (2019) Graph-VRNN Each agent is a node with a VRNN to capture dynamic uncertainties, and hidden states of these VRNNs interact through a graph interaction network. Yang et al. [352] (2020) - Uses LSTM for motion encoding, social graph attention with obstacle avoidance for interaction modeling, and a pseudo oracle predictor to handle uncertainties via informative latent variables. Wang et al. [353] (2020) GraphTCN Combines edge-feature GATs (EFGAT) for spatial interactions and TCNs for temporal dynamics. Casas et al. [354] (2020) - Models scenes as fully connected graphs, using GNN for inter-agent dependencies, with latent variables encoding stochastic scene dynamics. Zhang et al. [355] (2020) ST-LGSL Constructs spatiotemporal graphs using a latent graph structure, MLP and kNN; employs diffusion graph convolutions and gated temporal convolutions. Zhao et al. [356] (2020) GISNet Encodes each vehicle’s past trajectory for interaction-aware prediction. Li et al. [357] (2020) EvolveGraph Static graph learning for interactions and dynamic graph refinement to model temporal dependencies. Salzmann et al. [1] (2020) Trajectron++ Uses CNN to encode semantics from environment, LSTMs and attention for agent histories and interactions, and CVAE to generate predictions based on GNN-encoded features. Haddad and Lam [358] (2021) STR-GGRNN Uses Gated Graph Recurrent Neighbourhood Network (GGRNN) to construct dynamic graphs with Non-negative Matrix Factorization (NMF) to adapt neighbourhood structures. Zhang et al. [359] (2021) - Dual-scale context fusion with layers for geometric and topological features. Cao et al. [360] (2021) SpecTGNN Captures inter-agent correlations and temporal dependencies in frequency and time domains, using spectral graph and temporal gated convolutions. Zhou et al. [361] (2021) AST-GNN using spatial GNN for interaction modeling and temporal GNN for motion pattern extraction, combined with a time-extrapolator CNN for predictions. Chen et al. [362](2021) SCSG Uses multi-head attention for social interactions and star graph decoder. Fang et al. [363] (2022) - Uses Heterogeneous Risk Graph (HRG) to capture collision risks; and Hierarchical Scene Graph (HSG) for agent-environment constraints. Li et al. [364] (2022) - Combines two GNN modules: one for recognizing interactive events and another for predictions, integrated with LSTM for multi-step predictions. Hasan et al. [365] (2022) MESRNN Leverages meta-path features in spatio-temporal graphs for interactions, improves long-term predictions and social compliance in dense crowds. Xu et al. [366] (2022) GroupNet Multiscale hypergraph neural network: for pair-wise interactions across different groups sizes - multiscale hypergraph topology inference; for group-wise interactions - hypergraph neural message passing. Xu et al. [367] (2022) TGNN Uses a domain-invariant GNN and attention module for features transfer. Gilles et al. [368](2022) GOHOME Represents HD map as graph and uses graph convolutional layers capture spatial and agent-lane interactions. Bae et al. [369] (2022) - Assigns pedestrians to behaviour groups, forming intra- and inter-group interaction graphs; uses group pooling operations to capture collective dynamics. Zhu et al. [370] (2023) Tri-HGNN Triple Policies: extrinsic-level policy uses GNN for interactions, intrinsic-level policy uses GCN to infer intentions, and basic-level policy to integrate features from both policies using TCN. Westny et al. [371] (2023) MTP-GO Uses Graph-GRU units to encode interactions and neural ODEs to capture dynamic motion constraints. Xu et al. [372] (2023) EqMotion Includes a reasoning module to infer interaction graphs and employs equivariant geometric feature learning. Westny et al. [373] (2023) - Combines MTP-GO with differentially-constrained motion models. Mo et al. [374] (2024) - Uses Heterogeneous Graph Social (HGS) pooling to model vehicles and infrastructure in a unified graph for vehicle-infrastructure interactions. Kim et al. [375] (2024) - Uses a collision-aware kernel function to construct adjacency matrices and higher-order graph convolutions to model indirect social influences.
Table 9: Summary of reviewed DL-based Models relying on GNN (part 2, GCNs)
Reference Name Description Li et al. [27] (2019) GRIP Alternates between temporal convolutional layers for motion feature extraction and graph operation layers for feature propagation. Li et al. [376] (2019) GRIP++ Refines interaction modeling with dynamic edge weights and incorporating scene context. Liu et al. [97] (2020) - Constructs pedestrian-centric spatiotemporal graphs, leveraging GCNs to aggregate contextual and temporal features, with GRU-based temporal connections integrating node-level features over time. Liang et al. [377] (2020) LaneGCN Four blocks: ActorNet, encoding past trajectories, MapNet, encoding lane graph features; FusionNet, modeling interactions; and the Prediction Header, generating multimodal predictions. Mohamed et al. [378] (2020) Social-STGCNN Constructs social behaviour graphs with GCNs for group-based interactions, recursive updates, and an LSTM decoder for trajectory prediction. Sun et al. [379] (2020) - Uses GCNs to model group social interactions recursively, combining with BiLSTM-encoded individual features, uses LSTM-based decoder. Xue et al. [380] (2020) - Models interactions using a star graph combined with scene features extracted by a DeepLabv3+ encoder, employs a gating mechanism, a VAE for multimodality, and LSTM encoders/decoders. Liu et al. [381] (2021) AVGCN Integrates learned attention weights from gaze data and visual field constraints to infer neighbor importance. Shi et al. [382] (2021) SGCN Uses sparse graphs with adaptive adjacency matrices, capturing spatial & temporal dynamics via self-attention and asymmetric convolutions. Bae and Jeon [383] (2021) - Constructs multi-relational weighted graphs and employs disentangled multi-scale aggregation and global temporal aggregation. Zhao and Liu [384] (2021) STUGCN Usese kernel-based adjacency matrices to capture social interactions and Time-Extrapolator CNN to decode predictions from these embeddings. Zhou et al. [385] (2021) GroupTron A multi-scale graph neural network - models interactions at individual with LSTMs, group with STGCNs, and scene levels. Lv et al. [386] (2021) SSAGCN Employs social soft attention mechanism for interactions and a sequential scene attention mechanism; combined features are processed by GCNs and TCNs to generate predictions. Sheng et al. [387] (2021) - Models spatial interactions with a GNN applied to proximity-based weighted graphs and temporal features using a CNN-based Temporal Dependency Extractor (TDE). Schmidt et al. [388] (2022) CRAT-Pred Combines crystal graph convolutional neural networks (CGCNNs) and multi-head self-attention to model social interactions among vehicles. Liu et al. [389] (2022) MDST-DGCN Motion encoder for individual motion features, multilevel dynamic spatiotemporal directed graph encoder (MDST-DGEN) for multilevel social interaction modelling and fusion, motion decoder for generating predictions. Su et al. [390] (2022) - Employs directed graph topologies (view, direction, and rate graphs) to model asymmetric interactions between agents. Lu et al. [391] (2023) - Combines individual trajectories via a TCN, scene context using a 2D CNN with temporal attention, and interaction patterns with a Spatio-Temporal Dynamic GCN. Lian et al. [392] (2023) PTP-STGCN Uses a spatial GCN with a crowd interaction attention function to capture spatial interactions and a temporal transformer to extract motion features. Xu et al. [393] (2023) MVHGN Combines multi-view logical correlations and adaptive spatial topology; logical-physical features are extracted using graph convolution modules and regional clustering. Chen et al. [394] (2023) STIGCN Uses spatial and temporal graphs to model correlation between social interactions and pedestrian movement. Wang et al. [395] (2024) - Combines directed graphs with multi-head self-attention and graph convolutions to model spatial interactions. Du et al. [396] (2024) SFEM-GCN Combines semantic, position, and velocity-based social force graphs into a mixed spatio-temporal graph. Liao et al. [397] (2024) MFTraj Four blocks: behaviour-aware module with dynamic geometric graphs, position-aware with LSTMs, interaction-aware module with adaptive GCN, and a residual decoder.
Table 10: Summary of reviewed DL-based Models relying on GNN (part 3, GATs)
Reference Name Description Huang et al. [262] (2019) STGAT Uses graph attention mechanism for spatial interactions and an LSTM to encode temporal correlations. Kosaraju et al. [398] (2019) Social-BiGAT Utilizes GAT for interactions, MLP and LSTM to encode pedestrians’ past trajectories, CNN as scene-level feature extractor and Bicycle-GAN to generate predictions. Zheng et al. [399] (2020) GMAN Employs an encoder-decoder architecture with Spatial-Temporal Attention blocks and gated fusion to capture dynamic spatio-temporal correlations. Li et al. [400] (2020) Social-WaGDAT Uses spatio-temporal graphs with double-attention mechanisms for interactions, and context- and a kinematic constraint-aware decoder. Eiffert et al. [401] (2020) - Uses RNN-based encoder-decoder with a Mixture Density Network (MDN) and Graph Vehicle-Pedestrian Attention Network (GVAT) to model heterogeneous interactions. Carrasco et al. [30] (2021) SCOUT Uses three attention mechanism: self-attention: for temporal dependencies, cross-attention interactions for Models interactions attending relevant features from neighbouring agents and contextual attention for scene context. Monti et al. [402] (2021) DAG-Net Incorporates both past interactions and future goals using two attentive graph neural networks. Mo et al. [403] (2022) HEAT Combines agent-specific dynamics encoders, a directed edge-featured graph for interaction modeling, and an adaptive map selector for road context. Yang et al. [404] (2022) PTPGC Uses a graph attention network to capture spatial interactions, and applies TCN to encode individual dynamics. Li et al. [405] (2022) STG-DAT Uses dual-attention mechanism to capture topological and temporal relationships and a kinematic constraint layer in the decoder. Tang et al. [406] (2022) EvoSTGAT Models recursive interactions across contiguous time points using a dynamic attention mechanism, evolving social influences frame by frame. Zhou et al. [407] (2022) GCHGAT Model intragroup, outgroup, and intergroup interactions, combining a vanilla generative adversarial network for rough predictions with a state-refinement module to incorporate interactions. Cheng et al. [408] (2023) GATraj Uses attention mechanisms for spatial-temporal dynamics, GCN for interactions, and a Laplacian mixture decoder for generating predictions. Li et al. [409] (2024) GRIN Encodes inter-agent latent codes to model social relations and intra-agent latent codes to represent individual intentions, leveraging these in a graph attention network for continuous pair-wise interaction learning. Zhang et al. [177] (2023) AI-TP Integrates GAT for interactions.
3.4.4 GANs
Generative Adversarial Networks (GANs) have emerged as a powerful framework for trajectory prediction, leveraging their ability to model complex, multimodal data distributions. In this context, the generator ($G$) predicts plausible future trajectories conditioned on past trajectories, while the discriminator ($D$) evaluates these predictions against real trajectory data to guide the generator’s learning. This adversarial training setup enables GANs to generate diverse and realistic trajectories that account for the inherent uncertainty in future motion. A key advantage of GAN-based models is their capacity to address the multimodal nature of trajectory prediction, where multiple plausible futures exist for a given past trajectory. This is achieved by conditioning $G$ on contextual information such as social interactions, environmental constraints, or road semantics, allowing it to adapt predictions to dynamic and complex scenarios. Table 7 provides a summary of GAN architectures in the literature, detailing the design of $G$ and $D$.
Gupta et al. [122] presented the Social GAN model, which combines sequence-to-sequence predictions with a pooling mechanism for the generator and a recurrent discriminator. The generator takes as input the past trajectories of all agents in the scene and employs an LSTM encoder to capture motion history. After encoding, the social pooling mechanism aggregates the encoded trajectories to capture interactions between agents. The generated pooled vectors reflect the influence of nearby agents on an individual’s trajectory. These pooled vectors are then combined with the encoded past trajectories and passed to the decoder, while the discriminator classifies plausible trajectories. Amirian et al. [121] proposed the Social Ways model, which aims to generate more diverse samples and avoid mode collapse. To model interactions, the model uses LSTM to encode the influence of other agents and applies an attention weighting process to generate interaction weights. Sadeghian et al. [120] focused on designing a model that generates trajectories plausible under physical and social constraints. The model consists of three modules: a feature extractor with CNN and LSTM encoders, a physical and social attention module, and a GAN network to generate trajectories. While the physical attention module learns spatial constraints, the social attention layer captures interactions between agents. Finally, Zhao et al. [63] introduced a multi-agent tensor fusion model, which employs a conditional generator for predictions conditioned on the static scene context and agents’ past trajectories. Both the generator and discriminator utilize CNN and LSTM-based encoders and an LSTM-based decoder.
3.4.5 GNNs
Graph Neural Networks (GNNs) are increasingly applied in trajectory prediction for their ability to model complex spatial and temporal interactions. GNNs operate on graphs where nodes represent traffic agents, and edges encode pairwise relationships, such as proximity or relative motion. GNNs are scene-centered models, and there are two approaches for constructing graphs: static and dynamic graphs, depending on how node and edge updates are handled over time. In dynamics graphs approach, a new graph $G_{t}$ is constructed at each time step from sensor data [1, 378, 410]. Since the number of agents varies across frames, the graph structure and adjacency matrix change dynamically, allowing for a variable number of nodes per time step. If an agent enters or leaves the scene, the graph updates accordingly without requiring a fixed structure. Static graphs approach maintain a fixed number of nodes throughout the model’s operation, as seen in models like [64, 121]. If fewer agents are present in a frame than the fixed number, dummy nodes are introduced to maintain a consistent structure. Conversely, if the number of agents increases in a later frame, new agents are assigned previously unoccupied dummy node slots. Usually, valid and dummy nodes are handled through masking to exclude inactive nodes from message passing and feature updates. Implementation-wise, many PyTorch-based implementations of GNNs operate on static graphs with a fixed adjacency matrix size for computational efficiency. For dynamic graphs, to process a sequence of $\{G_{t-T},...,G_{t}\}$ and learn temporal and spatial dependencies, Temporal Graph Networks (TGNs) can be employed, available in PyTorch Geometric.
Message passing can employ Graph Convolutional Networks (GCNs), which uniformly aggregate neighbor information, or Graph Attention Networks (GATs), which dynamically compute attention weights to prioritize more influential interactions. However, GCNs and GATs only capture spatial relationships from a single snapshot and do not inherently model temporal dependencies. To aggregate historical trajectory information and learn temporal dependencies, two approaches can be used when static graphs coupled with GCN/GATs are employed. The first option is to store past trajectories inside node embeddings, allowing the network to process temporal dependencies in a sequence-to-sequence manner. The second approach is to use an external temporal mechanism, such as recurrent networks (LSTMs, GRUs) or memory modules, to handle historical information before passing the processed features into the GNN.
Alternatively, dynamic graphs coupled with Temporal Graph Networks (TGNs) provide more flexibility by naturally capturing temporal dependencies across a sequence of graphs. However, TGN relies on consistent object IDs across frames to correctly track agent-specific memory updates. If object IDs are inconsistent, past information cannot be reliably retrieved, leading to incorrect temporal associations. In such cases, alternative techniques like memory-based retrieval or node re-identification should be applied. The final GNN layer outputs refined node embeddings, which are then decoded into future trajectories. We summarize GNN-based methods across three categories: GCN-based models in Table 9, GAT-based networks in Table 10 and the rest of GNN-based models are shown in Table 8.
Focusing on interaction modelling, Wang et al. [353] proposed the GraphTCN architecture, which combines edge-feature GATs to capture spatial information and TCNs for temporal dynamics. EFGAT models spatial interactions using relative spatial relations as edge features, dynamically learning adaptive adjacency matrices via multi-head graph attention layers. TCNs handle temporal dynamics with gated 1D convolutions, enabling efficient parallel processing for long-term dependencies. Li et al. [357] introduced EvolveGraph, which leverages a latent interaction graph parametrized by learnable edge embeddings. This framework employs a two-stage training process: static graph learning initializes stable interactions, while dynamic graph refinement models temporal dependencies. A recurrent decoder generates multi-modal predictions based on the evolving graph structure, with training guided by supervised trajectory loss and relational consistency loss. Trajectron++ [1], a SOTA graph-structured recurrent model for trajectory prediction, represents the environment as a spatiotemporal graph, where nodes represent agents and edges capture inter-agent interactions. Semantic information from the environment, encoded via a CNN, is integrated into the graph as node or edge features. The GNN processes these enriched features, encoding agent histories and interactions using LSTMs and attention mechanisms. Finally, a CVAE framework generates diverse and multimodal predictions based on the GNN-encoded features.
Continuing on SOTA models, the GOHOME framework [368] adopts a graph-based approach for trajectory prediction, representing the HD map as a graph with lane segments as nodes and connectivity as edges. Graph convolutional layers capture spatial and agent-lane interactions, while cross-attention integrates past trajectories and map features. Lane-level heatmaps identify high-probability regions for future trajectories. Among the latest models, MTP-GO is a graph-based probabilistic model designed by Westny et al. [371] and EqMotion is a prediction model with invariant interaction reasoning proposed by Xu et al. [372]. MTP-GO incorporates graph-gated recurrent units (Graph-GRU) to encode inter-agent interactions over time and employs neural ODEs to model dynamic motion constraints. Specifically, neural ODEs solve differential equations parametrized by neural networks to represent motion constraints. EqMotion ensures motion equivariance, guaranteeing predicted trajectories transform consistently under translation and rotation, and interaction invariance, maintaining stable inter-agent relationships. It achieves this through an invariant reasoning module for interaction graph inference and equivariant geometric feature learning to model temporal and spatial dependencies.
Table 11: Summary of reviewed DL-based Models relying on Attention Mechanism
Reference Name Description Mercat et al. [411] (2019) - Multi-head self-attention mechanism applied before and after encoding the input, where multiple heads specialize in distinct interaction patterns. Zhu et al. [412] (2019) ARNP Leverages global latent variables to encode uncertainty and capture fine-grained local dynamics. Pan et al. [413] (2019) - Attention to aggregate information from lanes surrounding vehicle. Bhat et al. [414] (2020) Trajformer Operates within a transformer-based backbone, combining multi-headed self-attention to encode social and spatial dependencies among agents. Luo et al. [415] (2020) - Goal-oriented lane attention mechanism, leveraging dot-product attention to predict the probability of each lane a vehicle is heading toward. Kamra et al. [416] (2021) - Uses a pairwise attention mechanism with fuzzy decisions, focusing on sender-receiver interactions. Messaoud et al. [417] (2021) - Multi-head attention mechanism captures spatio-temporal dependencies, and uses multiple heads to model different-order interactions. Liang and Zhao [418] (2021) NetTraj Local graph attention captures dynamic spatial dependencies with road networks and sliding temporal attention for historical states. Li et al. [419] (2021) RAIN Hybrid attention mechanism combining RL-based hard attention to select key information and soft graph attention to rank selected elements. Shafiee et al. [420] (2021) Introvert Uses 3D spatio-temporal attention mechanism to capture human-human and human-space interaction, focusing on relevant scene elements. Messaoud et al. [421] (2021) - Uses multi-head attention, each head specializing in capturing distinct interactions and predictive modes. Sekhon and Fleming [422] (2021) SCAN Interleaved spatial and temporal attention mechanism. Xu et al. [423] (2021) Tra2Tra Global social spatial-temporal attention mechanism for interactions. Nayakanti et al. [424] (2022) Wayformer Employs multi-axis attention and optimizations like factorized attention and latent query attention to manage spatio-temporal and multimodal data. Lin et al. [425] (2022) STA-LSTM Uses spatial-temporal attention mechanisms. Zhang et al. [426] (2022) - Integrates attention mechanism for temporal and spatial dependencies. Peng et al. [427] (2022) SRAI-LSTM Social Relation Attention-based mechanism focusing on interactions within walking groups, aggregating motion features from neighbouring pedestrians. Li et al. [428] (2022) - Unified spatial-temporal attention mechanism to select key information from surrounding individuals in both spatial and temporal domains. Zhang and Li [429] (2022) MTPT Uses a modified Swin Transformer, attention to identify the most critical spatial and contextual factors. Da and Zhang [430] (2022) PAGA Path-Aware Graph Attention to capture interactions between lanes and agents by learning attention between vertices in a heterogeneous graph. Cao et al. [431] (2022) - Attention with total variation temporal smoothness prior for interactions. Guo et al. [432] (2022) - Dual-attention mechanism integrating temporal and spatial attention. Duan et al. [433] (2022) CAGN Normal and inverse attention to capture frequent and rare patterns. Wu et al. [434] (2023) - Uses self-attention for temporal dependency, sparse self-attention for spatial interactions, and cross-attention for spatio-temporal correlations. Tang et al. [180] (2024) HPNet Triple Factorized Attention module: spatial agent attention for inter-agent interactions, historical prediction attention, and self-attention across different trajectory modes. Li et al. [435] (2024) MFAN Mixing feature attention mechanism to determine the importance of different interaction features. Sun et al. [436] (2024) - Multiplex attention for interactions, by constructing layers in a latent graph, each layer for distinct interaction type.
Table 12: Summary of reviewed DL-based Models relying on Transformers (part 1)
Reference Name Description Yu et al. [437] (2020) STAR Combines temporal Transformers for modelling individual pedestrian dynamics and a graph-based spatial Transformer to capture crowd interactions, with alternating layers to model spatio-temporal dependencies. Giuliari et al. [87] (2021) - Uses encoder-decoder Transformer, encoding past trajectories with positional encodings and self-attention. Yuan et al. [438] (2021) AgentFormer Agent-aware Transformer modelling temporal and social dimensions, with attention mechanism for intra-agent and inter-agent interactions. Quintanar et al. [439] (2021) - Transformer architecture augmented with positional, heading information. Sui et al. [440] (2021) - Incorporates bounding boxes, speeds, orientations, and road semantics, weighted by an attention module, for trajectory and intention prediction. Yin et al. [441] (2021) MTN Models long-range temporal dependencies and uses co-attention mechanism for observed trajectories and ego-vehicle speed. Liu et al. [442] (2021) mmTransformer Combines motion extractor for past trajectories, map aggregator for road topology, and a social constructor to model agent interactions. Huang et al. [443] (2021) - Employs multi-head attention for interactions and multi-modal attention mechanism to capture diverse agent-map relationships. Girgis et al. [444] (2021) - Uses interleaved temporal and social multi-head self-attention. Ngiam et al. [445] (2021) - Uses self-attention layers for inter-agent interactions and integrates road graph information using cross-attention mechanisms. Chen et al. [2] (2021) S2TNet Self-attention sub-layer for interactions and temporal Transformer decoder. Saleh [446] (2020) - Context-Augmented Transformer fusing positional information, agent interactions, and scene semantics. Zhou et al. [447] (2022) GA-STT Combines spatial and temporal Transformers, using cross-attention to fuse spatial-temporal embeddings and capture both individual and group-level motion features. Tsao et al. [448] (2022) Social-SSL Introduces three pretext tasks: interaction type prediction, closeness prediction, and masked cross-sequence to sequence prediction for encoding. Hu and Cham [449] (2022) - Uses current-frame observations as query to suppress cascading errors and integrates spatio-temporal information in the encoder via self-attention and in the decoder via flipped cross-attention. Hu et al. [450] (2022) - Holistic Transformer with sparse multi-head attention for noise reduction, feature selection sparse attention for leveraging prior knowledge, and multi-head attention for utilizing posteriori knowledge. Li et al. [451] (2022) - Graph-Based Spatial Transformer with Memory Replay. Zhou et al. [452] (2022) HiVT Two-stage Transformer: first stage focuses on local context extraction and second stage integrates a global interaction module. Zhu et al. [453] (2022) - Graph-Based Transformer with a memory mechanism to enhance temporal continuity and employing HuberLoss for improved robustness. Su et al. [454] (2022) - Uses cross-modal transformers for cross-relation features between modality pairs, fused with a modality attention module. Achaji et al. [455] (2022) PreTR Uses a factorized spatio-temporal attention module to extract features and non-autoregressive parallel decoding mechanism to mitigate exposure bias. Zhang et al. [456] (2022) - Graph-Based Transformer uses GAT to model agent-agent and agent-infrastructure interactions, combining CNNs for spatial features. Amirloo et al. [457] (2022) LatentFormer Uses hierarchical attention to model interactions among agents and Vision Transformer to capture local and global scene context.
Among GCN models, the GRIP architecture [27] is designed for interaction-aware trajectory prediction and consists of three main components. First, the Input Preprocessing Model converts observed trajectories into a 3D tensor, followed by constructing a spatial graph where edges encode proximity-based interactions. Then, the GCN alternates between temporal convolutional layers to extract motion features and graph operation layers to propagate features across nodes using a spatial adjacency matrix and normalized graph operations. Its successor, GRIP++ [376], enhances the original GRIP by refining interaction modelling with dynamic edge weights and incorporating scene context. GRIP++ introduces advanced GNN architectures with improved attention mechanisms, enforces temporal consistency, and optimizes computational efficiency. Liang et al. [377] proposed LaneGCN, which constructs a lane graph from vectorized HD maps to preserve topology and connectivity. The architecture includes four modules: ActorNet, encoding past trajectories of traffic agents; MapNet, encoding lane graph features using LaneGCN; FusionNet, modelling interactions between agents and lanes; and the Prediction Header, generating multimodal trajectory predictions with confidence scores. By integrating structured map features and dynamic actor-map interactions, LaneGCN effectively addresses motion forecasting in structured driving environments. Lastly, Liao et al. [397] proposed a map-free, behaviour-driven trajectory prediction model - MFTraj. The architecture consists of four components: the behaviour-aware module leveraging dynamic geometric graphs; the position-aware module emphasizing relative positions and spatial dynamics using LSTMs; the interaction-aware module with adaptive GCNs to model spatio-temporal dependencies; and a residual decoder that processes these features into predictions.
Among GAT Networks, Social-BiGAT, designed by Kosaraju et al. [398], is a generative model for pedestrian trajectory forecasting. It employs GAT to capture social interactions and a Bicycle-GAN architecture to generate multimodal trajectories. The model integrates an MLP to embed each pedestrian’s relative displacements, an LSTM to encode temporal pedestrian movements, and a CNN to extract scene-level features from a BEV representation. To address the mode collapse issue in multimodal trajectory prediction, Cheng et al. [408] proposed GATraj, an attention-based graph model. It utilizes attention mechanisms for modelling spatial-temporal dynamics, a graph convolutional network for capturing agent interactions, and a Laplacian mixture decoder to ensure diverse and reliable multimodal predictions.
3.4.6 Attention and Transformers
Table 13: Summary of reviewed DL-based Models relying on Transformers (part 2)
Reference Name Description Chen et al. [458] (2023) VNAGT Combines a variational non-autoregressive graph transformer with a unified graph attention module to capture social and temporal interactions. Mozaffari et al. [459] (2023) - Decoder integrates maneuver-specific information via cross-attention to predict multimodal trajectories and maneuver sequences. Feng et al. [460] (2023) MacFormer Two key modules: the Coupled Map integrates map constraints and Reference Extractor to identify and select the most relevant map features. Huang et al. [461] (2023) GameFormer Uses self-attention mechanisms to extract spatial and temporal context and decoder refines interactions based on level-k game theory principles. Shi et al. [462] (2023) - Mode-level encoding and social-level decoding. Shi et al. [463] (2024) MTR++ Employs a symmetric scene context encoding and a mutually-guided intention querying module to enhance interaction-aware predictions. Liu et al. [464] (2024) - Uses GCNs to aggregate social interactions and refine adjacency matrices for spatio-temporal graphs and Transformers for temporal dependencies. Shi et al. [465] (2024) MTR Motion Transformer uses motion query pairs to combine global intention localization and local movement refinement.
Attention mechanisms enhance prediction models by allowing them to focus on the most relevant spatial and temporal features in complex environments. Spatial attention mechanisms assign importance weights to neighboring agents computed at each timestamp, based on features such as relative distance, velocity, and heading, which influence the target agent’s trajectory. Temporal attention processes sequences of historical states, assigning weights to time steps that have the most significant impact on future predictions. This is often integrated with recurrent architectures, where the hidden states from prior time steps are passed through an attention module to refine the context vector for trajectory prediction. Advanced implementations combine spatial and temporal attention in multi-head architectures, enabling the model to process multiple attention patterns in parallel, effectively handling heterogeneous agents in dense environments. Table 11 provides a summary of reviewed models relying on attention mechanisms.
Transformers extend attention mechanisms by unifying spatial and temporal dependencies across agents and time steps in a global framework. In trajectory prediction, they learn from past trajectories by employing self-attention layers within an encoder-decoder architecture. The encoder processes input trajectories combined with positional encodings to preserve temporal order while capturing long-range dependencies across time steps. Self-attention enables the model to filter out less relevant variations and identify meaningful motion patterns. The decoder generates future trajectories by leveraging cross-attention, which refines predictions based on learned spatio-temporal dependencies, ensuring consistency with past motion patterns. Tables 12, 13 provide a summary of Transformer-based prediction models.
Among SOTA Transformers, Liu et al. [442] proposed mmTransformer, a multimodal trajectory prediction framework utilizing stacked Transformers to hierarchically process information. It integrates a motion extractor for past trajectories, a map aggregator for road topology, and a social constructor to model agent interactions. The model employs a region-based training strategy, grouping trajectory proposals into spatial regions to optimize predictions for each region. Ngiam et al. [445] introduced the Scene Transformer, a factorized Transformer architecture leveraging self-attention layers to model inter-agent interactions. It incorporates road graph information via cross-attention mechanisms, enabling agents to account for both static and dynamic road features. Chen et al. [2] proposed a Spatio-Temporal Transformer for heterogeneous traffic prediction. The model integrates a spatial self-attention sub-layer to capture agent-agent interactions and combines these with a temporal Transformer decoder to refine predictions autoregressively, jointly modelling spatio-temporal dependencies within the encoder. Zhou et al. [452] presented the Hierarchical Vector Transformer (HiVT), a two-stage framework for multi-agent motion prediction. The first stage focuses on local context extraction, aggregating spatial and temporal features in localized agent-centric regions, while the second stage captures long-range dependencies across the scene through a global interaction module. This hierarchical approach ensures computational efficiency. Finally, a series of works [437, 451, 456] explore graph Transformers, which apply Transformer-based attention mechanisms to graph representations. Here, nodes denote agents and edges represent interactions. These models capture spatial dependencies through graph-based attention and integrate temporal dependencies using sequential attention mechanisms or positional encodings, enabling the simultaneous capture of local and global context.
3.5 Reinforcement Learning-based Methods
Reinforcement Learning-based methods generate predictions by interpreting agent behaviour as a high-dimensional policy optimization problem. RL methods learn through a reward system to generate trajectories that maximize long-term objectives. This section reviews two main categories of methods existing in the literature: Inverse Reinforcement Learning (IRL) and Imitation Learning.
Inverse Reinforcement Learning (IRL) methods aim to learn a reward function from expert demonstrations by observing trajectories and inferring the underlying structure that drives the demonstrated behaviours. Markov Decision Process (MDP) is a common framework used in IRL for trajectory prediction [466, 467, 468]. To obtain an estimate of the reward function, expert demonstrations are first analyzed to provide an initial estimate, capturing the implicit motivations behind the observed behaviour. Using the MDP framework, this reward function is iteratively refined by ensuring that the expert’s observed actions align with the optimal policy derived from the reward function. Once the reward function is fully inferred, it can be used to model decision-making and predict socially-aware trajectories in dynamic environments. For estimating the reward function, Maximum Margin IRL methods [467, 469] aim to minimize the feature expectation difference between expert demonstrations and predictions. Choi et al. [467] employ Maximum Margin IRL to estimate the reward function while enforcing a margin-based separation between expert and non-expert trajectories. By alternating between reward optimization and trajectory prediction, this approach ensures realistic, context-aware predictions that align with observed behaviors. Another approach, Maximum Entropy IRL methods [470, 466] aim to estimate the reward function by maximizing the entropy of the policy distribution under the inferred reward, ensuring robustness in cases with suboptimal or ambiguous expert trajectories. González et al. [466] proposed a model that infers a cost function as a linear combination of static and dynamic features from expert demonstrations. Their approach further incorporates Maximum Entropy IRL to resolve ambiguities in cost estimation and integrates a heuristic policy model for risk-aware decision-making.
On hierarchical inverse reinforcement learning, Sun et al. [471] proposed a two-layer model: the first layer handles discrete decisions, such as yielding or merging, which partition the trajectory space into distinct classes; and the second layer focuses on continuous decisions, such as velocity adjustments and smoothness optimization, within each class. This framework learns probabilistic models for both layers from human driving demonstrations using Maximum Entropy IRL. Lastly, on the integration of IRL and Transformers, Geng et al. [472] proposed a framework leveraging a spatial-temporal Transformer network to generate trajectory candidates. IRL is then applied to infer a reward function from expert-like behavior, evaluating these trajectories to select the optimal one.
Imitation Learning models [473] bypass the stage of learning explicit reward functions and instead extract an optimal policy directly from the data, generating future trajectories similar to expert demonstrations. Rhinehart et al. [474] introduced the Re-parametrized Pushforward Policy (R2P2) model, which combines imitation learning and probabilistic modelling. While the model implicitly learns a policy mapping states to future trajectories through a neural network, it models uncertainty and multimodality in future trajectories using a latent variable framework. A latent variable is sampled from a fixed distribution to capture variability, and the re-parametrized pushforward mechanism transforms these latent variables into trajectory predictions via a neural network. Hu et al. [475] propose a model that leverages a probabilistic latent dynamics framework to generate future states and actions based on offline expert data, employing BEV representations to enhance spatial reasoning. Generative Adversarial Imitation Learning (GAIL) combines imitation learning with adversarial training, where a generator produces trajectories and a discriminator differentiates between expert and generated trajectories to optimize the policy. Kuefler et al. [476] extend GAIL by incorporating recurrent policies using GRUs, enabling the model to capture temporal dependencies in driver behavior. Li [477] enhances GAIL with a latent decision-making process, where latent decisions inferred from historical trajectories guide multimodal path generation. The generator-discriminator framework is further improved by modelling these latent decisions as distributions and optimizing mutual information for greater diversity and realism in predictions. Choi et al. [478] integrate GAIL with a Partially Observable Markov Decision Process (POMDP), allowing the generator to account for both observable past states and unobservable latent states when generating trajectories, while the discriminator ensures alignment with real data.
4 Incorporating Driving Knowledge
To enhance performance and increase the explainability of trajectory prediction frameworks, recent studies [479, 480, 481, 482] have incorporated driving knowledge. The underlying intuition is that integrating external information to constrain motion reduces the set of possible future trajectories. In most scenarios, vehicles operate within designated lanes and adhere to traffic rules. Including information such as lane geometry, routes, road boundaries, and traffic regulations at intersections and roundabouts plays a crucial role in accurate predictions. Another type of driving knowledge that can improve prediction accuracy includes vehicle motion constraints, such as turning radius and braking distance, as well as pedestrian mobility limitations. Additionally, accounting for different driving styles and behavioral priors can improve model generalization and performance. For instance, in some regions, it is common to turn without signaling or to take a turn from a neighboring lane instead of the designated one to avoid traffic congestion. While such actions may constitute traffic rule violations, they can also be viewed as region-specific behavior patterns, reflecting common trends rather than individual choices. This section explores different types of driving knowledge and the ways in which this knowledge is incorporated into prediction frameworks. For a more detailed discussion, the reader is encouraged to refer to the survey paper [9].
4.1 Types of Driving Knowledge
We categorize driving knowledge into three groups: i) map-related information encoding the static environment, such as road topology, pedestrian crossings, and stop lines; ii) agent-related information about the kinematics of vehicles and the physical constraints of pedestrians; and iii) traffic-related knowledge, which indicate any additional information related to traffic rules, such as driving styles and interaction patterns between road users.
4.1.1 Map-related knowledge
HD maps provide information about road geometry, traffic lights, stop lines, pedestrian crosswalks, and traffic signs. Section 2.2 discusses rasterizing map data into grid images and encoding them using CNNs. While effective for extracting spatial features, this method often results in information loss, as the structured and geometric details of the map are rasterized, reducing precision. Recent advancements use GNNs to encode HD maps, preserving structural and semantic details while remaining scalable and effective.
One spotlight example is VectorNet, proposed by Gao et al. [410], which represents HD map elements and agent trajectories in a vectorized format. Each map element is broken into polylines composed of individual vectors, forming local graphs where nodes represent vectors and edges capture spatial or sequential relationships. These local graphs are aggregated into a global graph to encode lane-to-lane or agent-to-lane interactions using a hierarchical GNN. This approach enables interaction-aware feature encoding for downstream tasks. Next, Liang et al. [377] introduced LaneGCN, which leverages HD maps by constructing a lane graph where each lane segment is a node, and edges denote connections like predecessors, successors, and neighboring lanes. The model dynamically encodes the scene by capturing the topology and dependencies of the lane graph. By integrating actor-map interactions through a fusion network, it combines real-time traffic information with map-based context for prediction. Deo et al. [483] proposed a complementary approach, which focuses on conditioning trajectory predictions on specific paths within the lane graph. The proposed method uses discrete policy rollouts to select plausible routes, reducing computational complexity by targeting relevant map regions. By decoupling lateral variability (e.g., route choices) and longitudinal variability (e.g., speed, acceleration) and focusing predictions on sampled paths, the model enhances multimodal diversity, scene compliance, and trajectory prediction.
4.1.2 Agent-related knowledge
Agent-related knowledge enables prediction models to account for behavior and motion constraints. A vehicle’s kinematics determine its turning radius, braking distance, and acceleration capacity — all factors that significantly impact its potential future trajectory [221, 484]. Another aspect of agent-related driving knowledge involves understanding different driving styles. By analyzing a target vehicle’s historical trajectories, models can draw on established principles of driving behavior to identify patterns that align with known driving styles. Observing maneuvers and typical trajectory choices enables the model to estimate the driver’s style based on these predefined categories. For instance, an aggressive driving style might suggest a higher likelihood of selecting risky trajectories, such as overtaking with insufficient gaps or increasing the risk of collisions by forcing following vehicles to brake abruptly.
4.1.3 Traffic-related knowledge
Traffic-related knowledge encompasses additional driving information beyond agent- and map-related categories. Traffic rules provide a foundation for constraining vehicle behaviour and narrowing the set of plausible trajectories. Incorporating explicit traffic rules, such as speed limits and right-of-way regulations, into prediction models helps ensure safety and compliance with real-world driving norms. Kumar et al. [485] discuss the impact of traffic rules on intersection interactions, where vehicles may be required to stop, yield, or proceed depending on traffic signals or signs. In roundabouts, Hu et al. [486] highlight how right-of-way rules help define permissible behaviors and generate conflict-free trajectories for all participants. Similarly, Ju et al. [487] note that vehicle acceleration often reflects adherence to traffic regulations, such as slowing down at crosswalks or maintaining appropriate speeds in speed-restricted zones. By encoding traffic rules into models, predictions can be better aligned with expected behaviors in regulated environments.
Another type of traffic-related knowledge is interaction knowledge, which helps participants coexist in dynamic environments. In the study [485], pairwise interactions at intersections are categorized into discrete types such as ignoring, going, or yielding, based on the behavior of the target vehicle. Ban et al. [488] extend this approach by introducing follow and yield interactions, requiring models to explicitly predict these labels to better capture traffic dynamics. For more complex scenarios, such as roundabouts, Hu et al. [486] define interaction types using reference paths of interacting vehicles, enabling the generation of conflict-free trajectories. Ju et al. [487] link interactions to acceleration patterns, proposing a sub-network that predicts vehicle acceleration to infer interaction-driven trajectories. Additionally, social norms and informal yielding practices can also be of use, e.g., pedestrians giving way to those with trolleys, or a driver signaling a pedestrian to pass. These implicit forms of interaction knowledge can be formalized and incorporated into models to better predict complex and multimodal interactions in dense traffic scenarios.
Table 14: Summary of reviewed Models leveraging Driving Knowledge
Reference Type of Knowledge Description Lee et al. [64] (2017) map, traffic-related For map knowledge, utilizes static semantic information from maps, such as road layouts, drivable areas, and road boundaries; for traffic knowledge: uses predefined cost function evaluating trajectories based on collision avoidance, road compliance, and interaction dynamics. Xie et al. [221] (2018) agent-related Leverages vehicle kinematic parameters such acceleration and velocity. Cho et al. [312] (2019) traffic-related Integrates traffic rules as Signal Temporal Logic (STL) and filters out invalid trajectories. Greer et al. [489] (2020) traffic-related Uses a designed auxiliary loss function, which penalizes trajectories deviating from the expected lane heading. Yoon et al. [490] (2020) map-related Incorporates curvilinear coordinates, lane boundaries, and feasible maneuvers from HD maps. Cui et al. [491] (2020) agent-related Embeds a kinematic model of two-axle vehicles, constraining turning radius, acceleration, and heading to align predictions. Kawasaki and Seki [492] (2020) map, agent-related Uses a vector map to define lane geometry and encodes the geometric relationship between a vehicle’s state and surrounding lanes. Incorporates vehicle-motion constraints. Deo et al. [483] (2021) map-related Captures geometric and semantic details like lane centerlines, stop lines, crosswalks, and traffic flow directions. Bahari et al. [493] (2021) map, agent-related For map-knowledge, extracts from HD map lane geometry, centerlines, and boundaries; for agent-relates knowledge, integrates kinematic constraints through Model Predictive Control (MPC), such as turning radius and speed limits. Wang et al. [494] (2022) map-related Uses lane segments extracted from HD map, encoding geometric and semantic information, as proposals for predictions. Tang et al. [484] (2022) agent-related Incorporating vehicle dynamics constraints through Model Predictive Control (MPC).
4.2 Injecting Driving Knowledge
Formalizing and incorporating driving knowledge into prediction models can enhance AVs reasoning and adaptability to uncertain or unfamiliar scenarios. Human drivers accumulate this knowledge through learning and experience. Similarly, AVs can benefit from a combination of injected driving knowledge and learned patterns from data. Table 14 provides a brief summary of learning-based models, which incorporate driving knowledge elaborating on how the knowledge is injected and how it is formalized.
For agent-related knowledge, several works focus on integrating kinematic parameters to ensure physically feasible trajectory predictions. Xie et al. [221] proposed the Interactive Multiple Model (IMM) method, which leverages vehicle kinematic parameters such as acceleration and velocity in a physics-based model. Cui et al. [491] embedded a kinematic model of two-axle vehicles into the prediction pipeline, constraining outputs like turning radius, acceleration, and heading to align with vehicle dynamics. Additionally, works [493, 484] incorporated vehicle dynamics and kinematics constraints through Model Predictive Control (MPC), ensuring that predicted trajectories comply with speed limits, acceleration, and other physical constraints.
In terms of traffic-related knowledge, DESIRE [64] incorporates this knowledge through a predefined cost function within its Inverse Optimal Control (IOC)-based ranking module. The cost function evaluates trajectory predictions based on criteria such as collision avoidance, road compliance, and interaction dynamics, assigning lower costs to socially and contextually plausible trajectories to ensure alignment with expected traffic norms. Greer et al. [489] proposed the YawLoss framework, which integrates traffic-related knowledge by enforcing vehicle motion constraints aligned with lane directionality. This is achieved through a designed auxiliary loss function, YawLoss, which penalizes trajectories deviating from the expected lane heading, thereby ensuring adherence to legal and social driving norms, such as lane-following behavior. Cho et al. [312] integrates traffic-related knowledge by representing predefined traffic rules as Signal Temporal Logic (STL) formulas. The model jointly predicts future trajectories and the degree of satisfaction (robustness slackness) for these rules, filtering out invalid trajectory predictions for surrounding vehicles.
Map-related knowledge represents the largest class of models integrating driving knowledge. Wang et al. [494] proposed a lane-based prediction model that represents lane segments as vectorized elements, encoding geometric and semantic information such as centerlines, lane types, and heading angles, enabling interaction modelling between agents and lane segments. Deo et al. [483] integrates map-related knowledge by representing HD maps as lane graphs, capturing geometric and semantic details like lane centerlines, stop lines, crosswalks, and traffic flow directions. Nodes represent lane segments, while edges encode successor relationships and potential lane changes, ensuring trajectories align with legal routes. Yoon et al. [490] proposed a road-aware prediction model that incorporates curvilinear coordinates, lane boundaries, and feasible maneuvers from HD maps, using this information as prior knowledge to constrain trajectory predictions. Kawasaki and Seki [492], in the LAMP-Net architecture, use a vector map to define lane geometry and introduces a novel lane feature that encodes the geometric relationship between a vehicle’s state and surrounding lanes, enabling lane-aware trajectory predictions. Similarly, Bahari et al. [493] extract lane geometry, centerlines, and boundaries from HD maps to generate scene-compliant initial trajectory predictions. Lastly, DESIRE [64] also integrates map-related knowledge through the Scene Context Fusion (SCF) layer, which utilizes static semantic information from maps, such as road layouts, drivable areas, and road boundaries, to provide contextual guidance for trajectory generation.
5 Accounting for Uncertainty
Figure 11: Types of Uncertainty: the left side shows the categorization followed in the section and the right part demonstrates the flow of the prediction model and propagation of perception and tracking modules’ errors.
Motion prediction in complex traffic scenarios is fraught with high degrees of uncertainty coming from both external environment and internal system. Uncertain intentions of traffic participants, randomness of human behavior, and violations of traffic rules can be considered external sources of uncertainty. Non-obedience to traffic rules can tangibly disrupt the predictability of motion, as traffic rules are meant to constrain behavior and reduce the set of possible trajectories. As for internal systems, uncertainty arises primarily from the limitations of the prediction model itself. These limitations include biases in the training data, lack of robustness to unseen scenarios, and an inability to generalize effectively across diverse driving conditions.
For the detect-track-predict paradigm, there is a notion of propagated uncertainty. Failures in the perception module, such as inaccurate bounding boxes, missed detections, or false detections, can directly degrade tracking accuracy. Errors in the tracking system, such as incorrect object associations or ghost tracklets, propagate to the prediction module, which relies on these tracklets to generate future trajectory forecasts. These cascading errors compound the uncertainty in motion forecasting. The majority of existing prediction models assume error-free detectors and trackers, typically taking ground-truth past trajectories as their input during development and evaluation. However, in real-world deployment, such assumptions are unrealistic. Each subsequent system in the pipeline should incorporate mechanisms to mitigate the effects of upstream errors and quantify uncertainties. This includes developing models that explicitly account for propagated uncertainty, ensuring the prediction module remains robust to imperfections in the detection and tracking stages.
To address these challenges, uncertainty can be categorized into aleatoric and epistemic. Aleatoric uncertainty arises from the intrinsic randomness in data. In the context of autonomous driving, it is often linked to sensors’ characteristics, such as noise in sensor measurements, limited range, calibration issues, or low resolution. These factors introduce noise and ambiguities that pose inherent challenges to prediction models. Aleatoric uncertainty reflects the limitations in fully capturing the environment. For instance, dynamic scenarios, such as predicting whether a pedestrian will cross a crosswalk or continue along a sidewalk, inherently involve aleatoric uncertainty due to their multimodal nature. This uncertainty can be further categorized as heteroscedastic (varying across samples, e.g., lighting conditions affecting image quality) or homoscedastic (uniform across samples, e.g., consistently miscalibrated sensors). Epistemic uncertainty arises from the lack of comprehensive data representing the diverse scenarios that occur in real-world traffic. It reflects the limitations in the model’s knowledge and can be reduced through improved data collection, better modelling techniques, and increased training diversity. For example, cross-dataset domain adaptation is critical for addressing epistemic uncertainty. Many motion prediction models assume that training, testing, and deployment data are identically distributed, but this assumption rarely holds in real-world applications. This domain gap can lead to degraded performance in operational settings, particularly when encountering rare or unseen scenarios. Addressing epistemic uncertainty is vital to ensuring that models generalize well and perform reliably in diverse, real-world settings.
In general, accounting for different sources of uncertainty in prediction models and autonomous driving, as well as quantifying and mitigating the associated risks, is crucial for the correct operation of autonomous vehicles in unfamiliar environments. This is especially important in the presence of abnormal behavior, sensor failures, partial occlusions, and complex traffic scenarios. Figure 11 summarizes the discussed categorization of uncertainty based on its sources, nature, and the phenomenon of cascaded uncertainty. For a more detailed examination of uncertainty sources and mitigation approaches in autonomous driving, we refer the reader to the review [495].
Table 15: Summary of reviewed Prediction Models quantifying Uncertainty
Reference Type of Uncertainty Description Bhattacharyya et al. [51] (2017) epistemic, aleatoric Captures aleatoric uncertainty by modelling the distribution of bounding box predictions and epistemic through posterior distributions over model parameters. Huang et al. [496] (2019) epistemic, aleatoric Uses a mixture-of-experts architecture, with confidence estimations determining the best predictor for specific scenarios. Djuric et al. [497] (2020) epistemic, aleatoric Quantifies aleatoric uncertainty by modelling displacement errors and epistemic uncertainty through dropout-based Bayesian analysis. Casas et al. [354] (2020) epistemic, aleatoric Captures scene dynamics, actor interactions, and unobservable variables in a distributed latent space using GNNs, leveraging the Evidence Lower Bound (ELBO) framework to balance reconstruction accuracy and latent variable expressivity. Pustynnikov and Eremeev [498] (2021) epistemic, aleatoric Quantifies uncertainty using a spectral-normalized Gaussian process (SNGP), which estimates posterior variance to measure scene uncertainty and detect out-of-distribution scenarios. Ivanovic et al. [499] (2021) epistemic, aleatoric Quantifies uncertainty by incorporating class probabilities from perception systems into a CVAE framework, where these probabilities influence the output distributions of predictions. Tang et al. [484] (2022) epistemic, aleatoric Incorporates quantified prediction uncertainty into MPC framework with a potential field that accounts for vehicle dynamics, road constraints, and prediction uncertainty. Tang et al. [500] (2023) interaction-driven Quantifies individual uncertainty and interaction-driven collaborative uncertainty using multivariate Gaussian and Laplace distributions. Liu et al. [501] (2023) aleatoric Maps trajectories into bivariate Gaussian distributions, separating intrinsic uncertainty, and employs diffusion processes for trajectory sampling. Gu et al. [502] (2024) map uncertainty Extends HD map estimation models to output Laplace-distributed location uncertainties for map elements and confidences for road elements, capturing spatial ambiguity.
5.1 Mitigating Uncertainty
The effectiveness of data-driven prediction models heavily depends on the quality and diversity of their training datasets. Out-of-distribution (OOD) scenarios, where traffic conditions differ substantially from the training data, lead to distributional shifts that challenge prediction models during real-world operations. To mitigate this uncertainty, it is crucial to ensure that the training data provides substantial coverage of diverse scenarios and conditions. However, defining what constitutes “substantial coverage” remains an open research question. Laurent et al. [503] proposed a criterion for comprehensive testing, requiring scenarios to cover key parameters of autonomous driving systems that affect decision-making, while employing statistical methods to manage uncertainties during simulation runs.
To address uncertainties from incomplete inputs, Liao et al. [504] proposed a physics-informed model employing a Wavelet Reconstruction network to recover missing historical data and a Kinematic Bicycle model to enforce physically plausible predictions. A Risk-Aware module further integrates real-time risk factors using probabilistic attention, enabling adaptation to underrepresented scenarios. Bhattacharyya et al. [505] mitigated aleatoric and epistemic uncertainty by employing synthetic likelihoods, which encourage diverse model predictions and better capture multi-modal data. Synthetic likelihoods relax the requirement to explain every data point, enabling the model to generate diverse hypotheses for uncertain and multi-modal scene predictions. Next, Zhang et al. [506] introduced the Out-of-Sight Prediction framework (OOSTraj), which mitigates uncertainty in predictions for out-of-sight agents by utilizing a mobile denoising encoder and a camera parameters estimator to address sensor noise and map sensor-based to visual trajectories. The framework employs a Transformer-based decoder to predict trajectories even in the absence of direct visual observations.
Pourkeshavarz et al. [507] proposed a causal disentanglement approach to mitigate uncertainty caused by distribution shifts. Using a Transformer-GNN with disentangled attention blocks, the model separates causal and spurious factors in driving scenes, simulates potential shifts, and incorporates an invariance objective to focus on causal features while suppressing spurious ones for enhanced robustness. Finally, Luo et al. [40] addressed cascaded uncertainty by designing an end-to-end framework, “Fast and Furious” (FaF), which integrates 3D detection, tracking, and motion forecasting into a single convolutional network. By jointly optimizing these tasks, FaF implicitly mitigates cascading errors across the detection and prediction pipeline.
5.2 Quantifying Uncertainty
To minimize driving risks, it is essential to quantify motion prediction uncertainty and integrate this information into the decision-making module for safer and more robust planning. A common approach is to generate multi-modal predictions for each agent [251, 486, 122, 61, 508], rather than relying on a single deterministic trajectory. Probabilistic predictions address the uncertainty caused by unknown intentions of traffic participants, where a single past trajectory can unfold into multiple plausible futures. Some approaches estimate distribution parameters to provide a probabilistic representation of potential outcomes, such as modelling future trajectories using Gaussian distributions [509, 116, 510, 511] or mixture models [61, 125]. Table 15 summarizes deep learning-based models for quantifying uncertainty.
Bhattacharyya et al. [51] proposed a Bayesian LSTM framework for uncertainty quantification, capturing aleatoric uncertainty by modelling pedestrian bounding box distributions and epistemic uncertainty through posterior distributions over model parameters. Final predictions are represented as probabilistic distributions, providing a measure of reliability. Huang et al. [496] introduced a variational neural network that evaluates the value of information from each input and integrates with other prediction models to create a composite expert system, generating confidence estimates across different time horizons to effectively manage uncertainty. Similarly, Djuric et al. [497] quantified aleatoric uncertainty by modelling displacement errors and epistemic uncertainty through dropout-based Bayesian analysis, combining prediction variances from multiple stochastic forward passes to derive final confidence estimates. Building on these models, Li et al. [512] designed UQnet, which employs a 2D histogram-based deep learning model with deep ensemble techniques, using entropy-based metrics to quantify both types of uncertainty. Lastly, Gilles et al. [513] quantified uncertainty by calculating the variance of predicted spatial probability distributions from heatmap outputs, providing confidence measures for trajectory predictions and dynamically adjusting the sampling radius based on estimated uncertainty.
6 Planning-conditioned Trajectory Prediction
Planning-conditioned trajectory prediction methods rely on the assumption that agents act rationally and that their future trajectories are guided by a motion planner. Recent SOTA models [514, 515, 516, 517] effectively adopt this paradigm by introducing planning-prediction-coupled frameworks, where predictions are conditioned on the observed agents’ motion plans or intentions. These methods go beyond solely relying on past motion, contextual, and interaction features; by modelling the notion that agents follow potentially planned trajectories or have specific destinations to achieve. This approach significantly enhances prediction accuracy, particularly in interactive and complex scenarios.
To review planning-conditioned approaches, we follow two categorizations. First, these approaches are divided into model-based planning, learning-based planning, and combined methods, which are reviewed in Section 6.1. The second categorization focuses on anchor-based prediction methods, where trajectories are generated based on predefined anchors or guidance points derived from the environment or learned from data. These anchors typically represent potential future goals, paths, or key waypoints, that influence trajectory prediction. Anchor-based methods can be further divided into goal anchor-based prediction, where tentative goals for the observed agent are estimated online, followed by the generation of paths toward those goals, and trajectory-anchor-based prediction, where trajectories are directly constructed based on environmental or physical constraints. This categorization is further explored in Section 6.2.
6.1 Planning Methods for Trajectory Prediction
Motion planning literature has extensively developed methods for generating safe and efficient trajectories for ego vehicles. This research has inspired the application of motion planning algorithms to predict the trajectories of other detected agents, under the assumption that their future paths align with specific planning objectives. As a result, planning-conditioned trajectory prediction approaches adopt a categorization similar to that of motion planning literature, namely, model-based and learning-based methods [518]. Model-based approaches rely on explicit rules, physical constraints, and optimization techniques to simulate rational agent behavior. These methods are interpretable and often incorporate principles like kinematics or dynamics to ensure physically feasible predictions. On the other hand, learning-based approaches infer plans or intentions directly from observed data, enabling adaptability to complex and diverse scenarios. Combined methods integrate the strengths of both paradigms. Reviewed works of these approaches in trajectory prediction are summarized in Table 16.
Table 16: Summary of Model-based, Learning-based, and Combined planning methods for Trajectory Prediction
Type Reference Description Model-based Gong et al. [519] (2011) Optimizes a handcrafted cost function to plan paths for the observed agents. Yan et al. [520] (2014) Introduces a probabilistic motion model towards specific goals in the environment. Vasishta et al. [289] (2017) Uses A* algorithm with the potential cost function for trajectory prediction. Cui et al. [491] (2020) Accounts for planning kinematic constraints while predicting trajectories. Learning-based Previtali et al. [104] (2016) Uses IRL for pedestrian trajectories by modelling their interactions with the environment, such as avoiding collisions or heading toward crosswalks. Choi et al. [478] (2020) GAIL for matching long-term distributions over states and actions. Zhou et al. [182] (2023) Uses reusable scene encoding and a query-based decoder with adaptive refinement to handle multimodal future behaviours effectively. Chen et al. [521] (2024) Uses GNN for trajectory predictions conditioned on guided-goals. Combined Approaches Song et al. [514] (2021) (PRIME) a prediction framework combining model-based trajectory generation with learning-based evaluation to ensure accurate, feasibility-guaranteed multimodal outputs. Chen et al. [522] (2024) Combines model predictive control with learned prediction models to optimize joint behavior.
6.1.1 Model-based planning approaches
Model-based planning for trajectory prediction forecasts agents’ future paths by leveraging models that simulate dynamics, kinematic constraints, decision-making processes, and interactions with the environment. This approach integrates physical motion principles, behavioral insights, control objectives, and road structure information to predict trajectories in a structured and interpretable manner. By explicitly modelling system dynamics, these methods produce feasible trajectories that reflect agents’ capabilities and account for their interactions with the environment [523, 524]. The main components of model-based planning for trajectory prediction include dynamics models, which simulate the evolution of an agent’s state over time using kinematic or physics-based representations [525]; cost function defining optimization objective, such as minimizing energy consumption, travel time, or collision risks [519]; planner solving the optimization problem and generating trajectories that minimize the cost function while satisfying constraints.
6.1.2 Learning-based planning approaches
Learning-based planning for trajectory prediction employs data-driven models to generate efficient and adaptable trajectories for observed dynamic agents [514]. These methods utilize techniques such as reinforcement learning and imitation learning to infer planning strategies directly from data, enabling agents to navigate complex and interactive environments effectively. Training on large datasets allows learning-based approaches to capture both high-level decision-making processes (e.g., navigating traffic junctions or performing lane changes) and low-level control actions (e.g., reacting to obstacles). Unlike traditional model-based methods that rely on explicit motion models, learning-based planning adapts to diverse and unstructured scenarios by learning implicit patterns and behaviors from data. Additionally, these methods excel in trajectory prediction by modelling interactions among agents within dynamic environments, offering robust and versatile predictions [61, 411, 64].
Chen et al. [522] proposed the Interactive Joint Planning (IJP) framework, which optimizes the behaviors of the ego vehicle and surrounding agents through deep-learning-based prediction models trained on agent interaction data. This bidirectional integration ensures that the ego vehicle’s planning influences the prediction of other agents’ behaviors while incorporating the impact of their anticipated responses. The IJP framework employs Model Predictive Control (MPC) to generate precise motion plans by integrating deep-learned prediction models with gradient-based trajectory optimization techniques, enabling the vehicle to adjust its trajectory based on predicted interactions with surrounding agents. Previtali et al. [104] employ IRL to model agent behavior by inferring cost functions from observed trajectories. IRL is used to predict pedestrian trajectories by modelling their interactions with the environment, such as avoiding collisions or heading toward crosswalks. The approach involves learning a cost function that represents human decision-making processes and using a planning algorithm to generate plausible future paths based on these inferred behaviors.
The work by Deo and Trivedi [251] assumes that each surrounding agent can perform one of several possible maneuvers, such as lane changes or braking, which are learned using LSTMs without explicitly identifying their final goals. The model assigns probabilities to these potential maneuvers and predicts the corresponding trajectories by considering the interactions between the ego vehicle and surrounding agents. Next, Ziebart et al. [526] proposed a model that assumes pedestrians behave as rational planners, efficiently moving toward their destinations. The prediction model learns a cost function that explains observed pedestrian trajectories and uses it to predict future positions. The ego vehicle’s planning module integrates this predictive model to generate a trajectory that avoids pedestrian obstacles while pursuing its own objectives. The intention-interaction graph method proposed by Li et al. [527] jointly models pedestrians’ intentions (goals) and their social interactions. The graph represents each pedestrian and their intended destination as nodes, enabling the model to predict trajectories by inferring goals based on observed behaviors and interactions, without requiring explicit knowledge of the final destinations.
6.2 Anchor-based Trajectory Prediction
Figure 12: A general pipeline for anchor-based prediction models: (left) - goal anchor-based and (right) - trajectory anchor-based.
Anchor-based trajectory prediction methods assume that future trajectories are conditioned on planning objectives, such as goal destinations or potential paths. These methods use predefined spatial or trajectory “anchors” to guide the prediction process. Figure 12 illustrates the general pipeline for the coupled planning and prediction framework commonly found in the literature. The pipeline begins by encoding past trajectories and scene context into a context vector that captures essential features. This context vector is then utilized in the anchor-based decoding process to predict possible future trajectories, which can be guided by two types of anchors: goal anchors or trajectory anchors.
Goal anchor-based methods focus on predicting a set of feasible goal points that serve as anchors for trajectory generation [528]. These methods predict potential goals based on the current context and agent behavior [528, 529]. Once the goals are identified, trajectories are generated by conditioning on these anchors, often through a path-planning or motion-generation module. Examples include predicting key destinations, such as parking spots or intersections, and using goal-conditioned policies to create paths leading to these predicted goals. A significant advantage of goal anchor-based methods is their natural alignment with goal-driven tasks, such as reaching specific destinations. Trajectory anchor-based methods, on the other hand, rely on predefined or learned trajectory templates, referred to as anchors, which represent potential future paths [3]. These methods involve selecting or predicting one or more trajectory anchors as the basis for possible future motions. The anchors provide a coarse representation of the motion, which can either be refined or directly used for trajectory prediction. Examples of trajectory anchors include predefined templates for straight, curved, or turning paths. Table 17 provides a summary of reviewed anchor-based prediction methods.
Table 17: Summary of Anchor-based Trajectory Prediction methods
Type Reference Description Goal Anchor Based Zhang et al. [530] (2020) Selects potential goal points based on observed trajectories and contextual features and ranks candidate goals based on learned scoring function. Tran et al. [531] (2020) Dual-channel framework: a goal channel predicts potential destinations, and a trajectory channel generates detailed paths conditioned on these goals. Mangalam et al. [532] (2020) Generates goals by sampling plausible future endpoints directly from the scene context, leveraging map features, and observed trajectories. Mangalam et al. [533] (2021) Hierarchically models goals, waypoints, and paths to address epistemic and aleatoric uncertainty, generating scene-compliant multimodal predictions. Zhao and Wildes [534] (2021) Leverages Dynamic Time Warping (DTW) to retrieve pseudo goal positions from an expert repository of previously observed trajectories. Yao et al. [179] (2021) Uses a recognition network to encode the relationship between past trajectories and target goals, and a prior network for sampling latent variables. Wang et al. [535] (2022) Generates coarse stepwise goals iteratively, incorporates them into the encoder and decoder using an attention-based aggregator. Chiara et al. [285] (2022) U-Net-based architecture to predict future goals as probability distributions based on scene semantics. Wang et al. [536] (2023) Uses a ”GoICrop” module to aggregate map features around potential goal areas, incorporating spatial interactions between actors and maps. Dong et al. [537] (2023) The Generates goals by learning sparse representative instances of future trajectories using a mask generated from observed trajectories. Bae and Jeon [538] (2023) Predicts a sequence of control points to divide a pedestrian’s trajectory into three sections, capturing intermediate decisions. Wang et al. [528] (2023) Generates goal anchors via self-attention mechanism on the unified agent-centric scene representation, embedding rich contextual information. Chen et al. [521] (2024) Goals are predicted for each agent using a goal-guided state refinement (SSR) module. Trajectory Anchor Based Song et al. [514] (2021) Samples dynamic reference paths and imposes explicit kinematic and environmental constraints to produce feasibility-guaranteed trajectories. Varadarajan et al. [3] (2022) Learns latent anchor embeddings to guide context-aware trajectory distributions, enhancing diversity and prediction accuracy by dynamically adapting to the scene context. Zhou et al. [529] (2023) Uses a proposal module to generate adaptive trajectory anchors recurrently, which are subsequently refined by an anchor-based module, incorporating context-specific biases.
6.2.1 Goal Anchor-Based Prediction Methods
To predict plausible goal points as anchors for trajectory generation, models utilize approaches that consider the current context and agent behavior. Once these goals are identified, trajectories are conditioned on them, typically using a path-planning or motion-generation module. Examples include forecasting key destinations, such as parking spots or intersections, and employing goal-conditioned policies to create paths toward these objectives. A key advantage of this approach is its natural alignment with goal-driven tasks, like reaching a specific destination. Goals are often estimated based on static environmental features, including road topology, lane structures, or sidewalks. In road networks, goals can be defined as lane endpoints, intersections, or turn exits, providing a structured basis for trajectory prediction.
Among SOTA data-driven trajectory prediction methods, a set of studies integrate goal-directed principles from planning to infer potential goals and generate goal-conditioned trajectories [539, 81, 532, 540]. By estimating plausible goals, these approaches significantly enhance trajectory prediction accuracy. The BiTraP model [179] utilizes a bidirectional decoder conditioned on predicted goals to reduce accumulated errors. The SGNet framework [535] introduces a stepwise goal estimator that infers and leverages goals at multiple temporal scales. Similarly, PECNet [532] focuses on predicting distant goals to improve the accuracy of long-term trajectory predictions.
6.2.2 Trajectory Anchor-Based Prediction
Trajectory anchor-based methods focus on predicting potential paths (trajectories) that a vehicle might take to reach its goals [541, 542]. TPNet model [297] generates feasible future trajectories using polynomial curves. Inspired by the sampling-based paradigm in vehicle motion planning [543, 544], PRIME combines planning-based approaches with learning-based prediction by directly searching for reachable paths and generating potential trajectory anchors. This hybrid approach handles complex agent-map interactions while satisfying environmental and kinematic constraints [514], leveraging the strengths of both model-based and learning-based methods for optimal performance in complex scenarios. However, while much focus has been placed on improving point-level prediction accuracy, data-driven frameworks often struggle to enforce constraints during trajectory generation. To address this, Li et al. [545] proposed the Planning-inspired Hierarchical (PiH) framework, which employs a hierarchical lateral and longitudinal decomposition to generate path and goal intentions for agents. This framework directly influences the ego vehicle’s trajectory planning, ensuring alignment with predicted path intentions and promoting safe and efficient navigation.
6.3 Summary and Discussion
This section reviewed planning-conditioned trajectory prediction methods, focusing on their classifications, methodologies, and contributions to autonomous driving. These methods leverage motion planning principles to predict future trajectories while ensuring constraint compliance and contextual awareness. The reviewed approaches encompass model-based, learning-based, and hybrid techniques, along with anchor-based methods such as goal and trajectory anchors, each presenting distinct advantages and limitations in trajectory prediction.
Model-based approaches employ explicit mathematical representations of agent dynamics and environmental constraints to produce interpretable and physically feasible trajectories. Although reliable and constraint-compliant, these methods often lack adaptability to unstructured or highly dynamic environments. In contrast, learning-based methods excel in capturing complex interactions and multimodal behaviors by employing data-driven techniques, but their reliance on large datasets and inability to explicitly enforce safety constraints remain significant challenges. Anchor-based methods, including goal and trajectory anchors, offer a structured approach to trajectory prediction by aligning with real-world driving objectives. While effective in generating diverse trajectory options, these methods may encounter scalability and computational efficiency issues in dense and highly interactive scenarios.
Future research in planning-conditioned trajectory prediction should address these challenges by adopting hybrid frameworks that combine the robustness and interpretability of model-based methods with the adaptability and scalability of learning-based approaches. For instance, methods like PRIME [514], which integrate dynamic trajectory anchors with data-driven evaluation, exemplify the potential of such hybrid approaches to deliver multimodal, constraint-compliant predictions. Additionally, achieving computational efficiency remains critical for real-time applications, particularly on resource-constrained edge devices. Lightweight architectures and parameter-efficient designs can enhance scalability without compromising prediction accuracy. Planning-conditioned trajectory prediction is a crucial area of research that blends strategic motion planning with predictive accuracy. Overcoming current challenges by leveraging hybrid, interaction-aware, and computationally efficient methodologies will be key to advancing autonomous systems.
7 Vision-Language Models for Prediction
Vision-Language Models (VLMs) are seen as promising technology that could enable vehicles and systems to deeply understand their surroundings, marking a significant advancement in the era of intelligent and explainable transportation. VLMs [546, 547, 548, 549, 550, 551, 552, 553, 554, 555] break down the barriers between textual and visual information to effectively connect multimodal data. In the context of autonomous driving, VLMs can be roughly categorized into: Multimodal-to-Text (M2T) [556, 557, 558, 559] models, which take an image and text as input and generate text as output; Multimodal-to-Vision (M2V) [560, 561] models, where the input is an image and text pair and the output is an image; and Vision-to-Text (V2T) [562, 563] models, which take an image as input and generate text as output. This section mainly focuses on M2T models in the context of perception and prediction tasks. Readers are encouraged to look into the surveys [564, 565] for further details on other VLM configurations and downstream tasks. Figure 13 shows the general flow of a VLM-based prediction model. The model processes two main inputs: visual and textual data. While the image encoder processes visual inputs, e.g., single-camera images, multi-camera feeds, or LiDAR data, extracting spatial and object-level features; the text encoder interprets contextual information, which can include detailed instructions, scene descriptions, specific object inquiries, navigation commands, etc. The encoded features are then passed to a cross-modal attention mechanism [566, 567, 568], which captures the relationships between visual cues and textual instructions to generate joint representations. This representation is further processed by an autonomous driving task-specific network, e.g., as for the prediction task, by a context understanding layer and decoder. The generated future trajectory considers both visual observations and semantic guidance. By leveraging VLMs for autonomous driving tasks, high-level instructions can be aligned with visual perception, resulting in more flexible and human-like driving responses.
Figure 13: VLM-based trajectory prediction model: encoders for visual and textual inputs, a cross-modal attention mechanism to create a joint representation and context understanding layer tailored for a specific downstream task.
This section further reviews VLM-based models designed for trajectory prediction and traffic scene understanding. Given the limited number of prediction models in the literature, we have included a survey of traffic scene understanding models, considering their potential integration into prediction pipelines and their relevance to readers. Next, we provide a list of available language-enhanced AD datasets with brief descriptions. To conclude the section, we discuss the challenges associated with integrating VLMs into the autonomous driving stack.
7.1 Trajectory Prediction
Table 18 summarizes reviewed VLM and LLM-based models for trajectory prediction. Yang et al. [569] proposed a two-stage evaluation framework. In the first stage, a VLM model analyzes images from six camera views and a BEV visualization with labelled agents to rank the prediction difficulty of agents. To improve the model’s response accuracy, the VLM is provided with several examples of rankings constructed based on displacement errors from the ground truth. In the second stage, the VLM’s output is used to create a subset of training data based on the ranking, resulting in a highly representative and challenging training dataset that enhances the training efficiency of a conventional prediction model. Moon et al. [570] introduced VisionTrap that leverages human gazes and gestures, vehicle turn signals, and road conditions to enhance motion prediction accuracy. The model comprises four main components: a per-agent state encoder that generates agent-specific context features, a visual semantic encoder that encodes images combined with road segments, a text-driven guidance block that guides learning and understanding of detailed visual semantics, and a trajectory decoder. The study includes an inference-time analysis demonstrating the model’s suitability for real-time predictions.
Table 18: Summary of reviewed VLM and LLM-based models for trajectory prediction
Reference Name Dataset Description Keysan et al. [568] (2023) - NuScenes The prompt describes the target agent and its environment. Yang et al. [569] (2024) - NuScenes Uses VLMs to create a smaller and representative subset of training data to improve the training efficiency of conventional models. Moon et al. [570] (2024) VisionTrap nuScenes-Text Utilizes human gazes and gestures, vehicle turn signals, and road conditions to enhance motion prediction accuracy. Peng et al. [571] (2024) LC-LLM highD Uses LLM as an explainable lane change prediction model with refined Chain-of-Thought (CoT) reasoning. Zheng et al. [572] (2024) TC-Map LLM WOMD Generates Transportation Context Maps as visual prompts and corresponding text prompts to extract situational insights and pass them to a prediction pipeline.
Keysan et al. [568] expanded the previously proposed CoverNet model [69] by incorporating a text-based scene representation. The prompt accepted by the model contains structured information about the target agent and its environment, constructed from polyline and Bézier curve lane representations. Peng et al. [571] introduced an explainable lane change prediction model (LC-LLM). The study reformulates the problem of lane change prediction as a language modeling task, processing various driving scenario data as natural language prompts. To enhance transparency and prediction reliability, the model’s Chain-of-Thought (CoT) reasoning is refined to include explanatory elements in the prompts during the inference phase. Zheng et al. [572] leveraged LLMs to improve the global understanding of traffic context for motion prediction. The study transforms complex traffic scenarios and agents’ past trajectories into visual prompts (Transportation Context Maps) along with corresponding textual prompts. TC-Maps, combined with text prompts, enable the model to generate comprehensive situational insights, extracting details such as agent intentions and scenario types. The extracted information is then encoded and integrated into a conventional motion prediction pipeline.
Table 19: Summary of reviewed VLM and LLM-based models for scene understanding
Reference Name Dataset Description Guo et al. [573] (2024) VLM-Auto Custom simulated dataset Uses Qwen-VL model to analyze visual inputs for identifying environmental conditions and generating control parameters to guide driving behaviour. Gopalkrishnan et al. [574] (2024) EM-VLM4AD DriveLM A lightweight multi-frame model for VQA tasks in autonomous driving. Dona et al. [575] (2024) - Waymo PREPER CITY Evaluates hallucinations in pedestrian detection performed by GPT-4V and LLaVA models. Mo et al. [576] (2024) NexusAD CODA-LM A framework based on InternVL2-26B for understanding driving corner cases and generate perception and driving recommendations. Zhou et al. [577] (2025) ELM nuScenes Ego4D Embodied language model (ELM) to enhance scene comprehension using space-aware pre-training and time-aware token selection. Tian et al. [578] (2024) DriveVLM - VLM for planning, comprising modules for scene description scene analysis, and hierarchical planning. Chen and Lu [579] (2024) - 100 images from BDD100k Uses CogVLM model to handle sensors data and employs a Chain-of-Thought (CoT) method to structure reasoning into scene understanding, prediction, and decision-making.
Table 20: Language-Enhanced Autonomous Driving Datasets: types of sensors (Camera, LiDAR), source, year of release and tasks for which datasets are annotated (Single Object Referring (SOR), Action Explanation (AE), Visual Question Answering (VQA), Decision Making (DM), Image Captioning (IC), Visual Spatial Reasoning (VSR), Multi-object Tracking (MOT)).
Dataset Source Dataset Camera data LiDAR data Video Tasks Talk2Car [580] nuScenes - SOR BDD-OIA [581] BDD100K - AE NuPrompt [582] nuScenes - MOT Refer-KITTI [583] KITTI - MOT DRAMA [584] - - IC, VQA Talk2BEV [585] nuScenes - VQA Reason2Drive [586] nuScenes, Waymo, ONCE - VQA Rank2Tell [587] - VSR NuScenes-QA [588] nuScenes VQA
7.2 Traffic Scene Understanding
Table 19 summarizes reviewed VLM and LLM-based models for scene understanding. Choudhary et al. [559] introduced Talk2BEV, a VLM designed for BEV maps in autonomous driving. This model handles diverse tasks, including predicting traffic participants’ intentions, visual and spatial reasoning, and decision-making based on visual information. Romero et al. [553] argue that using VLMs to query images via natural language addresses issues of expressivity and the need for multiple single-purpose query models, but it often returns redundant and low-quality results. To improve this, they introduced Zelda, a system that adds discriminator and synonym terms to refine user queries into natural language prompts, enhancing accuracy by filtering out low-quality frames. Mo et al. [576] designed a framework for understanding driving corner cases, built upon the foundation model InternVL2-26B and consisting of four phases. For visual perception, the framework uses DINO for 2D detection and DepthAnything for depth estimation. This detection information, along with the original image, is processed by a scene-aware retrieval module to generate contextually relevant information. The Driving Prompt Optimization module then uses this information to guide the model, suppressing hallucinations and focusing on critical scene aspects, ultimately producing responses for perception and driving recommendations. Zhou et al. [577] introduced the Embodied Language Model (ELM) to enhance traditional VLMs’ scene comprehension, supporting tasks such as object localization, detailed scene description, prediction, and event memorization. ELM incorporates space-aware pre-training for robust spatial localization and time-aware token selection to effectively process long-term video data and retrieve relevant past events. Finally, Tian et al. [578] proposed the DriveVLM system for planning in complex environments, with modules for scene description, scene analysis, and hierarchical planning. The scene description module generates linguistic representations of driving conditions and passes these to the scene analysis module, which examines objects’ motion states and behaviours that may influence vehicle actions.
7.3 Benchmark Datasets
To advance research in VLM-based solutions for autonomous driving, the community is increasingly releasing language-enhanced datasets. Table 20 provides a summary of existing datasets, specifying the source datasets they are built upon and the tasks for which they are suited, such as object tracking, action reasoning, spatial understanding, and question answering in autonomous driving contexts.
Talk2Car dataset [580], designed based on nuScenes dataset, focuses on task of single object referring by pairing natural language commands with specific objects in driving scenes. Similarly, the BDD-OIA dataset [581], based on BDD100K, contain annotations enabling models to interpret the reasoning behind driving actions. Next, the NuPrompt dataset [582] expanded upon nuScenes, enabling models to track multiple objects based on descriptive language prompts. NuScenes-QA [588] introduced question-answer pairs, challenging models to interpret and respond to questions about driving environments. Talk2BEV [585], another dataset built upon nuScenes, supports VQA with an emphasis on BEV maps, which help models to understand spatial layouts from an overhead perspective. For a broader variety of environments, Reason2Drive [586] integrates data from three datasets nuScenes, Waymo, and ONCE, supporting diverse question-answer sets across varied driving conditions.
7.4 Summary and Discussion
VLMs offer significant potential to enhance trajectory prediction models by incorporating common knowledge and contextual understanding essential for nuanced interpretations of surroundings and environmental factors. This section briefly reviewed foundations of VLM-based models for trajectory prediction, outlined recent advancements in forecasting and scene comprehension models, and provided an overview of current language-embedded datasets for autonomous driving. Despite the notable capabilities that VLMs introduce, seamlessly integrating these models into autonomous driving systems still demands further, in-depth exploration.
VLMs are large-scale models, and the inference time of the VLM-based prediction pipeline is a primary concern when considering integration and computational demands. From an application standpoint, this may necessitate equipping the vehicle with a fully operational computer if inference is conducted locally. Model quantization [589] is a technique employed to reduce model size, offering faster inference at the cost of some accuracy. Other techniques, such as pruning redundant weights [590] and knowledge distillation [591], serve similar purposes. Further experimentation is needed to explore the trade-off between reasoning capability, model size, and inference speed. Hardware-dependent factors, such as parallelism and processor type (TPU, GPU, or CPU), also play a role in optimizing inference time. If not hosted locally, cloud solutions are a possibility. A set of available VLMs [592, 593] offer cloud hosting capabilities, implying that an AV would need to request information via an Application Programming Interface (API). This reliance can lead to latency or may require a shift in the type of computer added to the vehicle, favouring a communication-efficient setup over a computation-heavy station to minimize delays and ensure data transmission with minimal losses.
Another challenge is prompt “engineering”. The alignment of visual and textual modalities within VLMs is facilitated through prompts, leading to the development of prompt engineering —– the process of refining input prompts to guide the model toward a desired behavior or specific output. Providing context to VLMs, alongside the task they need to perform, has a considerable impact on output. For instance, specifying whether the model acts as a “driver” or a “pedestrian” yields different responses in traffic scene scenarios. Additionally, prompts can include instructions on desired output length and format to ensure that only essential information is generated. While the prompt alone may suffice for specific output formatting, few-shot learning is an alternative approach where an example of the expected output format or length is included in the prompt. More advanced techniques, such as Chain-of-Thought prompting, as described in certain models, guide the model through a logical sequence to enhance response quality. An efficient prompt, with reduced redundancies and a clearly specified output format, results in fewer required tokens and, consequently, faster inference times.
Lastly, the non-deterministic nature of some VLMs [594, 595], meaning they can produce varied outputs for the same image-text pair across different runs, requires specific strategies for effective integration. Techniques such as fine-tuning the temperature parameter to balance the model’s creativity, applying majority voting, or averaging results among the top-$k$ outputs across multiple runs are some of the strategies that can be employed. Balancing consistency and flexibility is essential, as VLMs are particularly valuable for addressing edge cases where conventional pipelines may encounter out-of-distribution scenarios. Forcing the model to consistently generate the same output for identical inputs enhances stability and predictability, but it may also limit the model’s ability to adapt to uncertain situations. Therefore, careful evaluation of generative versus deterministic approaches in the context of trajectory prediction is critical. While VLMs show promise for enhancing long-term trajectory predictions, determining their overall utility and optimal configuration within the autonomous driving stack remains an area for ongoing exploration and refinement.
8 Collaboration between AVs for Prediction
Collaborative perception enhances the capabilities of individual vehicle sensors by utilizing multiple data sources from various angles. By aggregating diverse sensor inputs, it allows for more accurate detection of distant and obscured objects. Depending on the collaboration mode, the existing SOTA models can be grouped into early, intermediate, and late collaboration. Another line of categorization is based on the collaboration task, e.g., prediction [596], planning [597, 598], task-invariant [599]. A substantial part of models in the literature aims to find a balance between increasing the prediction performance and limited bandwidth along with mitigating the impact of communication delays, which can lead to misalignment of the features. Under these challenges, we discuss in this section models, which address transmission delays, communication redundancy and pose misalignment. Lastly, we review the models aiming to collaborate through a unified fusion space addressing the variations in utilized sensors suits and different architectures of prediction modules. This section reviews collaborative models aiming to improve performance not only for prediction but rather various downstream tasks, e.g., detection, segmentation and tracking. Due to the novelty of the field in terms of application to trajectory prediction, reviewing only collaborative prediction models will not provide the reader with the proper overview of the field and available effective fusion mechanism that can be utilized or adapted for collaborative prediction. For more details and challenges on collaborative perception, we refer the reader to comprehensive survey papers [600, 601, 602, 603, 604].
8.1 Type of Fusion
Collaborative perception can be achieved by combining data and features at different stages of the perception pipeline. Currently, the literature distinguishes three fusion modes: early, intermediate, and late fusion (please refer to Figure 14).
In early fusion, raw data is aggregated at the beginning of the perception pipeline. This approach allows the ego-vehicle to address challenges like occlusions and distance limitations. However, transferring raw data requires significant bandwidth and necessitates efficient suppression techniques to transmit only complementary data, reducing message sizes. Chen et al. [605] proposed an unsupervised 3D representation learning approach (CO3), leveraging point clouds from vehicle and infrastructure side views to generate contrasting yet semantically aligned representations.
In intermediate collaboration, deep semantic features are exchanged between agents, and the ego-vehicle integrates these features to generate final predictions. Compared to early fusion, intermediate fusion involves more information loss and requires optimal feature selection and fusion strategies. Xu et al. [606] proposed V2X-ViT, a vehicle-to-infrastructure cooperative perception system based on a Vision Transformer model. This model employs heterogeneous multi-agent self-attention and multi-scale window self-attention to integrate data, addressing asynchronous data exchange and positioning inaccuracies. Zhou et al. [607] introduced a GNN-based framework for multi-robot perception, evaluated under challenging conditions like camera occlusion and heavy noise. Xu et al. [608] developed CoBEVT, a simple yet high-performing framework for collaborative BEV map creation. Its fused axial attention module (FAX) integrates camera features from various viewpoints and agents, capturing both dense local and broad spatial relationships. Wei et al. [609] introduced CoBEVFlow, a robust framework for collaborative perception using a BEV flow approach to manage non-uniform collaboration messages and transmit perceptual data without generating new features. Yang et al. [610] proposed SCOPE, a framework designed to unify spatio-temporal awareness across agents for effective information exchange. Wang et al. Wang et al. [611] introduced CORE, a bandwidth-efficient framework for multi-agent cooperative perception that combines a compression mechanism, attention-based integration of messages, and a reconstruction module. Lastly, Qiao and Zulkernine [612] developed an adaptive spatial-wise feature fusion model with a trainable feature selection module for efficient information aggregation.
Figure 14: Collaboration between AVs: (a) - scenario for collaborative perception, vehicle $iv_{j}$ can pass information about vulnerable traffic agents, such as pedestrians at a crossing, to vehicle $iv_{0}$. Additionally, both vehicles can refine each other’s predictions by sharing and complementing observations perceived from different angles. (b, c, d) schematically show different collaboration schemes: (b) - early collaboration, (c) - intermediate collaboration, and (d) - late collaboration.
In late fusion, predictions are generated independently by agents and shared at the end of the pipeline. The ego-vehicle combines these outputs to refine its predictions. While this mode is bandwidth-efficient, it can be affected by noise and incomplete data in the individual outputs. Lu et al. [613] proposed a hybrid collaboration framework (CoAlign) to mitigate pose estimation inaccuracies, using an agent-object pose graph to improve alignment and integrating features at multiple spatial resolutions. Yu et al. [614] introduced the DAIR-V2X dataset, a multi-modal and multi-view dataset for vehicle-to-infrastructure collaborative perception. It includes VIC3D, a benchmark model for jointly detecting and classifying 3D objects using vehicular and infrastructural sensors, addressing synchronization issues and transfer costs. Glaser et al. [615] tackled bandwidth limitations and obstructions by introducing the Multi-Agent Spatial Handshaking network, which integrates varying and overlapping raw sensor data from different viewpoints under strict bandwidth constraints.
8.2 Challenges
Current SOTA models not only focus on designing efficient fusion layers but also tackle critical challenges in collaborative perception, such as transmission delays, communication redundancy, and pose misalignment. Communication redundancy manifests in two ways: determining which surrounding agents should share data to complement the ego-vehicle’s observations and designing efficient compression mechanisms to minimize message sizes while maintaining essential information, thereby addressing bandwidth constraints. Transmission latency is another significant challenge, as limited communication speeds and large message sizes can lead to delays and asynchronous data fusion across different timestamps. This issue further complicates data alignment, which is already affected by inaccuracies in localization systems and transformation processes. Effective collaborative frameworks must incorporate mechanisms to handle these delays and misalignments. Lastly, pose misalignment, whether caused by transmission delays or inconsistencies in feature maps, remains a persistent issue. The following sub-sections categorize SOTA approaches based on the specific challenges they address.
Communication Redundancy. Liu et al. [616] proposed a three-stage handshake communication protocol for collaborative perception, addressing the challenge of determining which agents to communicate with. In the first stage, the ego-agent broadcasts a request message to nearby agents. In response, other agents calculate matching scores based on the received request and their own observations. The ego-agent then connects with the agents offering the highest matching scores to obtain the requested feature maps. Several works have focused on addressing bandwidth constraints by designing architectures that fuse the sparsest yet most informative and complementary features [617, 618, 619]. Li et al. [620] introduced a distilled collaboration graph (DiscoGraph), leveraging a teacher-student training methodology. The teacher model integrates early collaboration with comprehensive inputs, while the student model focuses on partial inputs at intermediate stages of collaboration. Hu et al. [621] introduced a spatial confidence map to address the uneven spatial distribution of perceptual data, enabling agents to share only sparse, critical information to enhance complementary perception. Lastly, Chen et al. [622] proposed the TransIFF framework, incorporating a Cross-Domain Adaptation module to align features and a Feature Magnet (FM) module to dynamically fuse instance-level features.
Transmission Delays. Lei et al. [623] highlight the practical challenges of latency in communication systems, which can disrupt collaborative perception. To mitigate these issues, the authors proposed a latency-aware collaborative perception framework. This framework includes SyncNet, a synchronization component that aligns asynchronous sensory data from multiple agents at the feature level. SyncNet employs attention mechanisms and temporal adjustment strategies to synchronize data into a unified timeframe despite delays. Yu et al. [624] introduced the Feature Flow Net (FFNet), which leverages temporal consistency in consecutive frames from infrastructure cameras. FFNet transmits static feature maps through feature-flow fusion, ensuring robust communication and minimizing the impact of latency on collaborative perception.
Pose Misalignment. Vadivelu et al. [625] proposed a model based on V2VNet [596] to address pose noise and its impact on the reliability of spatially dependent communications. To mitigate localization errors, the authors introduced a neural reasoning model capable of facilitating communication, identifying inaccuracies, and collaboratively resolving pose inconsistencies. Gu et al. [626] addressed the challenge of noisy poses and observation misalignments among multiple agents with FeaCo, a system designed to achieve robust feature-level consensus. FeaCo incorporates a Pose-error Rectification Module (PRM) to align feature maps from multiple vehicles and a multi-scale Cross-level Attention Module (CAM) to enhance information aggregation across agents.
Unified Fusion Space. The literature explores challenges arising from variations in sensor suites, model architectures, and the integration of collaborative components within existing systems. Xu et al. [627] addressed discrepancies in neural network architectures across collaborating agents by proposing a trainable feature resizer to harmonize feature dimensions and a sparse cross-domain transformer for domain adaptation. In the related study [628], the authors introduced a confidence calibrator to adjust for biases in prediction confidence scores across agents. Xiang et al. [629] tackled the challenge of heterogeneous sensor modalities with a model for hetero-modal cooperative perception, enabling collaboration among agents using different sensor suites. By removing the assumption of uniform sensor setups, the model supports new, diverse agent types in collaborative perception. Lu et al. [630] proposed the HEAL framework, which constructs a unified feature space using a pyramid fusion network and aligns new agent types to this space via backward alignment. Li et al. [631] addressed discrepancies in LiDAR data with a model comprising three components: a Domain-Mixing Instance Augmentation module to ensure uniform data representation, a Progressive Domain-Invariant Distillation module for domain-neutral feature adaptation, and a Domain-Adaptive Fusion module with calibration-aware attention mechanisms. Finally, Ma et al. [632] introduced the MACP framework, equipping single-agent models with cooperative capabilities by freezing the existing prediction model and integrating trainable lightweight modules.
8.3 Benchmark Datasets
Collaborative perception is advancing significantly with the release of new datasets that replicate complex urban environments and diverse driving scenarios, supporting the development of algorithms across multiple applications. Table 21 outlines the details of available simulated and real-world datasets. This summary includes datasets annotated not only for motion forecasting but also for tracking, which can be adapted for use in prediction tasks.
Table 21: Large-scale Collaborative Perception Datasets: types of sensors (Camera, LiDAR), number of agents interacting, types of communication (V2V - vehicle to vehilce, V2I - vehicle to infrastructure), number of frames and tasks for which the dataset are annotated (Object Detection (OD), Object Tracking (OT), Semantic Segmentation ($S$), Trajectory Prediction (TP), Motion Prediction (MP), Domain Adaptation (DA)).
Dataset Source Frame V2V V2I Camera LiDAR Agents Tasks V2V-Sim [596] Simulation 51K $\checkmark$ - - $\checkmark$ 1-7 OD, MP V2X-Sim [633] Simulation 10K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1-5 OD, SS, OT OPV2V [634] Simulation 11K $\checkmark$ - $\checkmark$ $\checkmark$ 1-7 OD, OT, MP LUCOOP [635] Real 54K $\checkmark$ - - $\checkmark$ 3 OD, OT V2V4Real [636] Real 20K $\checkmark$ - $\checkmark$ $\checkmark$ 2 OD, OT, DA V2X-Seq [637] Real 15K - $\checkmark$ $\checkmark$ $\checkmark$ 2 OD, OT, TP DeepAccident [638] Simulation 57K $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1-5 OD, OT, SS, MP, DA TumTraf-V2X [639] Real 7.5K - $\checkmark$ $\checkmark$ $\checkmark$ 2 OD, OT
Table 22: SOTA Collaborative Prediction Models.
Model Venue Dataset Task Input Results V2VNet Wang et al. [596] (2020) ECCV V2V-Sim Detection, Forecasting Point Cloud - V2VNet Vadivelu et al. [625] (2021) CoRL V2V-Sim Detection, Forecasting Point Cloud - V2X-Graph Ruan et al. [640] (2023) - DAIR-V2X-Traj Forecasting Vector Map - V2X-Graph Ruan et al. [640] (2023) CVPR V2X-Seq Forecasting Point Cloud minADE : 1.17, minFDE : 2.03, MR : 0.27 V2XFormer Wang et al. [638] (2025) AAAI DeepAccident Detection, Motion Prediction RGB Image mIOU: 56.2, VPQ: 44 CMP Wu et al. [641] (2024) - OPV2V Motion Prediction Point Cloud MinADE_6@5s: 1.74, MinFDE_6@5s: 4.05
In simulated datasets, V2X-Sim 1.0 [596] is a V2V dataset created using CARLA and SUMO, focusing on dense traffic scenarios and using 32-channel LiDAR for 3D perception. V2X-Sim 2.0 [633] extends this dataset by adding V2I elements and enhancing vehicle sensor setups. OPV2V [634] broadens the scope with varied road types and detailed vehicle annotations, supporting collaborative tasks such as object detection, tracking, and motion forecasting. DeepAccident [638] focuses on diverse collision scenarios in autonomous driving, with multi-modal data from vehicles and infrastructure, realistic accident simulations, and support for both perception and prediction tasks.
Among real-world datasets, LUCOOP [635] is a multi-vehicle dataset aimed at urban navigation, object detection, and connected autonomous vehicle research. It includes sensor data from multiple vehicles over a 4 km route, designed to simulate cooperative traffic using both static and dynamic sensors in an urban setting. The V2V4Real dataset [636], designed for V2V perception, covers various driving scenarios with annotated 3D bounding boxes and high-definition maps, collected over extensive urban and highway routes in Columbus, Ohio. This dataset includes five object classes and supports detection, tracking, and behavior prediction. V2X-Seq [637] explores the impact of infrastructure on perception and trajectory forecasting with 15,000 frames from vehicle and infrastructure perspectives across diverse urban scenarios. Lastly, TUMTraf V2X [639] captures complex traffic scenarios in day and night conditions, with nine sensors providing 360-degree views at intersections to address cooperative perception challenges like pose estimation errors, latency, and synchronization, making it highly applicable for real-world use cases.
8.4 SOTA Prediction Models
For benchmarking and evaluation of new collaborative prediction models, Table 22 summarizes existing collaborative prediction models detailing the type of fusion and dataset on which the model was evaluated.
V2VNet [596] is a network designed for joint perception and prediction. Its architecture comprises three components: first, a CNN block processes raw sensor inputs to generate a compressible intermediate form; next, a cross-vehicle integration mechanism combines the received data with the ego-vehicle’s data; and finally, the fused intermediate output is used to generate the final predictions. Robust V2VNet [625], an extension of V2VNet, focuses on addressing localization errors for improved performance. V2X-Graph [640] is a late fusion framework that facilitates collaborative trajectory forecasting using graph-based feature fusion. The framework constructs a graph where agents’ motion features and map elements are represented as nodes, and their spatial-temporal relationships as edges. It employs three interpretable sub-graphs: the Agent-Lane sub-graph (ALG) for map-based interactions, the Motion Fusion sub-graph (MFG) for motion features, and the Cooperative Interaction sub-graph (CIG) for inter-agent relationships. These sub-graphs aggregate heterogeneous features across views to enhance trajectory prediction accuracy. The V2XFormer model [638] is a V2X-based framework that utilizes a Swin Transformer as its backbone to extract BEV (Bird’s Eye View) features from multi-view camera data. These BEV features, generated by intelligent vehicles and infrastructure, are spatially aligned to the ego-vehicle’s coordinate system and fused through CoBEVT [608] to produce aggregated representations. The fused BEV features are subsequently used for tasks such as motion prediction, 3D object detection, and accident prediction. Lastly, the CMP framework [641] enhances trajectory prediction by fusing data at multiple stages. During perception, BEV features are compressed and shared for fusion using the FuseBEVT module [608], improving object detection and tracking. In the prediction aggregation stage, vehicles share their predicted trajectories, which are further refined through a multi-head attention mechanism for enhanced trajectory prediction accuracy.
8.5 Summary and Discussion
Collaborative perception overcomes limitations in individual perception systems, such as restricted sensing ranges, occlusions in dense environments, and sensor noise. However, its implementation poses significant challenges for collaborative trajectory prediction, including pose misalignment, communication redundancy, and time delays, which remain open issues. This section examines different fusion approaches, reviews SOTA models addressing these challenges, and summarizes collaborative datasets alongside descriptions of trajectory prediction SOTA models.
The concept of multiple vehicles sharing complementary information assumes a perfectly functioning, fully connected collaboration network. Most existing models operate under the assumption that the ego-vehicle aggregates information while all other nodes act as transmitters. A key research direction is exploring diverse communication topologies, transmitting frequencies and naturally occurring time delays within such networks. Current models primarily address bandwidth limitations but often overlook the impact of transmission delays. In dynamic tasks like trajectory prediction, data transmitted even 0.5 seconds late may become irrelevant for the ego-vehicle, underscoring the critical need to account for these delays. Another research direction addresses the assumption in existing models that the ego-vehicle performs fusion and decodes trajectories. However, none of the current studies provide an inference time analysis. In trajectory prediction, timely delivery is critical, as planning must receive predictions within a specific time window. Providing fused features could significantly slow down inference. Beyond time complexity analysis, it is essential to develop a smart strategy to determine when fusion is necessary and when to rely solely on existing individual predictions. In a collaborative mode, the assumption is that the system operates as a network of intelligent agents, each capable of autonomous operation and fusing information only when required.
Lastly, in a network of intelligent vehicles, the reliability and trustworthiness of shared information must be thoroughly evaluated before fusion. Autonomous vehicles perceive their environment using both on-board sensors and data received from other vehicles within the network. However, the accuracy of this perception depends not only on the performance of an individual vehicle’s sensors but also on the quality of the shared data. A transmitting vehicle may have malfunctioning or degraded sensors, introducing errors into the cooperative system. To address this, a verification step should be implemented to assess the credibility of incoming data, ensuring that only trustworthy information contributes to the fused representation. This process is critical to maintaining the integrity and effectiveness of the collaborative perception system.
9 Discussion
In recent years, both academic and industrial research communities have made significant progress in advancing trajectory prediction methods, characterized by diverse approaches and innovative solutions. This section summarizes the topics reviewed and addresses three key questions posed at the outset: (Q1) What are the existing solutions for trajectory prediction? Section 9.1 revisits available modelling approaches in the literature to answer this question. (Q2) What are the emerging research trends in trajectory forecasting realm? Section 9.2 discusses the primary focus areas of researchers, the challenges being tackled, and emerging trends. (Q3) Is the trajectory prediction problem solved, or is there still room for improvement? Section 9.3 talks about existing challenges and gaps in current studies. Finally, Section 9.4 elaborates more on open challenges and proposes potential future research directions.
9.1 On Research Question 1
Answering question 1, trajectory prediction has evolved through several distinct phases of modelling approaches, driven by advances in deep learning and machine learning techniques. Early deep-learning methods were heavily reliant on recurrent models, such as RNNs and LSTMs, which effectively captured temporal dependencies in sequential data. Around 2017-2018, these models gained significant attention for their ability to process dynamic motion patterns, often through hierarchical or multi-layer designs. As the field progressed, modeling more complex agent interactions led to the emergence of attention-based mechanisms and graph-based approaches.
The current trend shifts towards Transformer architectures and Graph Neural Networks (GNNs), which have emerged as SOTA since 2020. Transformers leverage self-attention to model both spatial and temporal dependencies across agents, offering adaptability in heterogeneous traffic environments. GNNs, on the other hand, excel in capturing agent-to-agent interactions through graph representations, enhancing their effectiveness in complex, multimodal scenarios.
SOTA models commonly integrate various architectures to address the complexities of trajectory prediction. These include CNNs for scene understanding, attention mechanisms for feature weighting, and encoder-decoder setups for trajectory generation. Examples of such combinations include graph-enhanced Transformers and CNN-LSTM pipelines. Additionally, hybrid solutions that combine physics-based modelling with deep learning approaches are gaining attention for their ability to integrate domain knowledge with data-driven adaptability; like, models incorporating kinematic constraints alongside neural networks enhance prediction accuracy while maintaining physical plausibility.
9.2 On Research Question 2
Answering question 2, recent research trends in the prediction domain focus on enhancing the robustness and generalization of prediction models. As discussed in Section 5, a significant research gap exists in quantifying and mitigating uncertainty within prediction pipelines. One approach to addressing aleatoric uncertainty is through collaborative perception, which compensates for individual perception system limitations, such as restricted sensing capabilities, occlusions, and sensor noise. Collaborative perception enables multiple vehicles to share information, creating a more comprehensive environmental view. This area, detailed in Section 8, is investigated in detection and tracking and is gradually expanding into prediction, underscoring the need for further investigation and the development of collaborative prediction models.
In improving robustness and generalization, formalizing and incorporating driving knowledge remains an active area of research, as outlined in Section 4. Planning-conditioned trajectory prediction methods, as discussed in Section 6, represent a promising direction by incorporating motion planning principles into prediction frameworks. These methods assume rational agent behavior and condition predictions on agents’ motion plans or intended goals. By modelling planned trajectories and potential destinations, planning-conditioned approaches provide interpretable and constraint-compliant predictions, particularly in interactive and complex scenarios. Additionally, Vision-Language Models (VLMs), as discussed in Section 7, offer common knowledge and contextual understanding necessary for navigating complex and uncertain environments. While VLMs hold promise as anchors in uncertain conditions, their real-world application in autonomous driving is still under evaluation. Practical challenges, such as hosting these large models on the cloud, can introduce latency issues that may compromise real-time predictions. The integration of VLMs into the autonomous driving stack requires further exploration to determine their effectiveness and deployment configurations.
These trends reflect a shift in trajectory prediction research towards developing robust, collaborative, context-aware, and planning-informed methods that prioritize both accuracy and real-time operation in dynamic environments.
9.3 On Research Question 3
On question 3, in the overview of the prediction pipeline, Sections 2.6 and 2.7 outline the metrics used for evaluating prediction models and list benchmark datasets with leaderboard solutions. A challenge in comparing different methods from the literature (apart from leaderboard rankings) lies in inconsistencies in the prediction horizon across the same dataset. Additionally, comparing methods based on performance across multiple datasets is difficult, as most methods are evaluated on only one or, at most, two datasets. Few studies address the critical issue of robustness, which is essential to ensure reliable predictions despite detection errors, tracking deficiencies, localization inaccuracies, or map changes. Furthermore, no research has systematically compared methods under such real-world conditions.
Despite the extensive range of methods and refined solutions, it is evident that the rate of improvement in prediction accuracy is diminishing. On leaderboards, the performance gaps between top solutions show only marginal gains. In practical scenarios, extremely precise trajectory forecasts may not significantly enhance the effectiveness of prediction modules. Human drivers, while less precise than autonomous vehicles, navigate effectively across varying traffic densities. Thus, understanding how such high precision contributes to the overall success of an autonomous stack, including collision avoidance, is crucial. While prediction models achieve higher precision in most scenarios, failures in specific situations could lead to severe consequences. Current studies lack qualitative evaluations to identify scenarios where predictions falter, and there is a noticeable gap in testing for generalization. Specifically, cross-dataset evaluations, where models trained on one dataset are tested on another, are underexplored, leaving performance in out-of-distribution scenarios unaddressed. Consequently, while existing models demonstrate strong performance using reported metrics, these evaluations often fail to comprehensively reflect real-world operational needs, requiring further testing and investigation.
9.4 Future Directions
This section outlines the existing research gaps and proposes future directions for further investigation, based on the comprehensive review and discussion presented.
9.4.1 Real-time Evaluation
For autonomous vehicles to navigate complex traffic environments safely, they require a deep semantic understanding of both static and dynamic scene elements. Achieving this understanding necessitates advanced deep learning architectures with multiple processing and reasoning layers to interpret contextual features, predict agent behavior, and plan appropriate responses. However, the incorporation of semantic layers to enhance comprehension inevitably increases inference time, presenting a critical trade-off between model complexity and processing speed.
Lin et al. [642] emphasize that real-time autonomous vehicle systems must process and react to traffic conditions faster than human drivers to ensure safety. While the fastest human reaction times range from 100–150ms [643, 644], autonomous systems are required to achieve processing latencies below 100ms and maintain a frame rate exceeding 10 frames per second. This dual requirement enables both rapid responses to individual events and continuous awareness of dynamically changing traffic conditions. These stringent latency demands have significant implications for the design of autonomous driving pipelines. In a complete pipeline encompassing detection, tracking, prediction, and planning, the combined latency must remain under 100ms, establishing a strict upper bound for the prediction module. Despite this, much of the current research focuses on maximizing prediction accuracy while neglecting real-world inference-time evaluation, creating a gap between theoretical advancements and practical deployment. Real-world applications demand a balance between accuracy and speed to enable seamless integration of prediction systems with planning modules, underscoring a critical area for further development.
When designing prediction systems, it is essential to prioritize architectures that can be effectively parallelized and fully leverage available hardware to minimize inference time. Attention-based models, in particular, offer significant advantages in this regard. Their parallel computation capabilities enable faster inference times on modern GPUs and TPUs optimized for such operations. While factors such as model size, sequence length, and memory requirements affect absolute performance, the parallel nature of attention mechanisms allows for further optimization through techniques like multi-head parallelization and batch processing. This advantage has been demonstrated in studies such as [645, 245], where optimized embedding systems combined with hardware-aware implementations significantly improved inference speed without sacrificing prediction accuracy.
Model selection should be guided by the hardware capabilities of the deployment platform and the system’s operational requirements. In some scenarios, a lighter, faster model (e.g., attention-based architectures with optimized implementations) that provides near-real-time predictions may be more practical than a more accurate but slower alternative. While such models may not achieve state-of-the-art accuracy on benchmarks, their ability to deliver rapid updates enables more frequent error correction and quicker system responses. In dynamic traffic scenarios, where predictions must be continuously refined, the capability to quickly adapt to changes is often more critical for safe navigation than marginally higher accuracy achieved at the cost of computational speed.
9.4.2 Enhancing Robustness
Future research must address key challenges to ensure the effective deployment of trajectory prediction systems in real-world autonomous driving environments. Most existing methods are tailored for specific tasks, scenarios, or motion types, performing well under controlled conditions with clear motion patterns, well-defined spatial structures, and predictable objectives. However, these models often falter in dynamic and unpredictable scenarios, highlighting the need for comprehensive evaluation across diverse and realistic conditions.
To achieve robustness, prediction systems must dynamically adapt to environmental changes and variations in agent behavior. Current models are predominantly trained on datasets from limited driving scenarios, which fail to capture the breadth of real-world diversity in driving styles and configurations. Enhancing adaptability to novel and unseen situations remains a critical challenge. Incorporating driving knowledge, such as traffic rules, road layouts, and interaction dynamics, can provide models with a strong foundation for understanding traffic systems, thereby improving their robustness and adaptability. However, managing rare and unexpected behaviors requires a different approach. Drawing inspiration from human cognition, autonomous vehicles (AVs) could adopt a dual-process strategy: relying on predefined, optimized algorithms in familiar scenarios while leveraging generalized reasoning mechanisms for unfamiliar or uncertain situations. Vision-language models (VLMs) show promise in enabling AVs to handle complex, unstructured edge-cases, and research in this area is rapidly advancing, warranting further exploration.
Another critical area gaining attention is the mitigation and quantification of uncertainties, specifically aleatoric (inherent randomness) and epistemic (model-related) uncertainties. These uncertainties must be explicitly addressed and conveyed to downstream planning systems, allowing them to make informed decisions. Despite progress in some models, effectively mitigating uncertainty remains a significant gap in current prediction pipelines. Many existing methodologies assume error-free detection and tracking, often overlooking the impact of perception errors on performance. Evaluation metrics like average displacement error (ADE) are commonly used, reflecting the accuracy of prediction model, but is not sufficient to comprehensively capture the nuances of failure cases or scenarios influenced by perception errors. This narrow focus neglects qualitative insights into critical edge cases where prediction failures could result in severe consequences, such as collisions. Future evaluation frameworks should emphasize a balance between robustness and accuracy. While minor displacement errors can be mitigated through conservative planning strategies, complete prediction failures pose a substantial safety risk. Robustness under perception errors must be prioritized, with evaluation strategies focusing on failure case analysis to build more resilient and reliable systems capable of navigating real-world complexities.
Additionally, most current models specialize in either short- or long-horizon predictions, neglecting the benefits of integrating both temporal scales. Long-horizon predictions are essential for strategic planning, enabling proactive trajectory optimization and efficient navigation. However, their inherent uncertainty increases with the prediction horizon. Conversely, short-horizon predictions offer precise immediate forecasts for rapid reactive measures but lack the broader context required for long-term planning. A more effective solution involves developing integrated prediction systems that combine the strengths of both temporal scales. In such systems, long-horizon predictions establish a strategic plan, continuously refined by short-horizon predictions in real-time. This iterative refinement creates a robust and adaptive framework capable of both proactive decision-making and reactive safety responses, addressing the complexities of real-world driving while ensuring safety and efficiency.
Finally, expanding a model’s capacity to handle novel scenarios beyond its initial training distribution is essential for safety. This issue can be addressed through simulation-to-reality (Sim2Real) adaptation and continual learning. Sim2Real techniques involve generating synthetic scenes mimicking real-world scenarios and conditions, allowing models to learn from controllable environments, where different levels of diversity can be gradually introduced before deploying the model in actual settings. For continual learning, during trial deployments or pre-deployment phases, a small dataset specific to the target city can be collected to adapt the model to local conditions. This continual learning phase ensures the model is fine-tuned to the deployment context while retaining prior knowledge. Although some works, such as [646, 647, 648], explore this direction, further research is needed to fully realize its potential in autonomous systems.
9.4.3 Heterogeneous Trajectory Prediction
In autonomous driving, trajectory prediction must account for all traffic agents, including vehicles of varying sizes (e.g., trucks, vans, sports cars), pedestrians, bicycles, and motorcycles. Each of these agents exhibits unique kinematics, significantly influencing their motion patterns. For vehicles, kinematic constraints, such as steering limits and acceleration bounds, heavily restrict possible trajectories. In contrast, pedestrian motion is far less constrained, introducing greater randomness and variability. This distinction makes pedestrian trajectory forecasting particularly challenging, as it is influenced by a broader range of factors, including individual intent, environmental context, and social interactions. A critical challenge lies in designing a unified prediction framework that accommodates the diverse characteristics of all agent types. One approach is to use a single model trained on extensive data to learn the dynamics of each agent class implicitly. While this enables scalability and uniformity, it often limits the model’s ability to explicitly integrate kinematic constraints, especially for agents with well-defined physical dynamics, such as vehicles. On the other hand, a specialized approach involves using separate prediction models tailored to specific agent classes, enabling the explicit incorporation of agent-specific constraints and behaviors. However, this method introduces complexity, including the need to accurately classify agents and manage potential detection and classification errors. In both approaches, the reliance on object classification and detection performance introduces dependencies that must be carefully addressed. Misclassification or detection errors can propagate through the system, degrading prediction accuracy. Therefore, robust mechanisms for error handling and uncertainty modelling are essential. Designing a comprehensive prediction framework that effectively integrates driving knowledge, accounts for the diverse nature of traffic agents, and mitigates detection-related challenges remains an open area of research. Such a framework must strike a balance between generalization and specialization, ensuring safety, accuracy, and scalability in real-world autonomous driving scenarios.
An additional challenge in human trajectory prediction lies in balancing individual and group-aware predictions. For individual motion, it is crucial to incorporate appearance-based features, posture detection, individual characteristics, and elements of behavioral analysis to enhance prediction accuracy. In contrast, group-aware predictions require modelling group formations, relationships, and dynamics to understand collective motion patterns effectively. Previous studies have highlighted the importance of these elements in understanding pedestrian group behavior and dynamics [649, 650, 651, 652, 653, 654, 655, 656, 657, 658]. Developing hybrid models that integrate crowd modelling with individual predictions represents a promising direction, especially in urban environments with dense pedestrian crossings. If a model can accurately recognize crowd formations and adapt by predicting the collective flow and motion of the crowd, computational resources can be significantly conserved. In such scenarios, the detector would not need to identify each individual or run predictions for every agent separately. Conversely, for people who leave the crowd or are not part of it, the system should seamlessly switch to individual-level predictions. This approach requires further investigation, particularly in the development of diverse and representative crowd datasets for autonomous driving applications. Existing datasets [159], serve as a foundation but must be expanded to include more complex and dynamic scenarios. Advancing this area of research will enhance the ability of autonomous systems to operate efficiently and safely in environments characterized by dense and dynamic pedestrian activity.
9.4.4 Collaborative Perception
Collaborative perception enhances the sensing capabilities of individual vehicles by integrating data from multiple sources and viewpoints. This integration allows for more accurate detection of distant or obscured objects and supports robust decision-making in dynamic environments. However, several critical challenges remain to be addressed. One key research direction is ensuring the feasibility of collaborative models in terms of inference time and computational resources. Another crucial aspect is the generalization of collaborative models. Learning-driven approaches excel at deriving comprehensive representations from complementary views, but their ability to generalize to new datasets and scenarios requires thorough investigation and evaluation.
Different collaboration modes introduce unique challenges that require tailored solutions. Vehicle-to-Vehicle (V2V) communication is crucial for enabling rapid and dynamic collaboration by facilitating the sharing of localized sensor data among vehicles. Future research should focus on optimizing data-sharing efficiency under bandwidth constraints and high-speed conditions. Developing context-aware protocols that dynamically determine which information to aggregate and transmit based on real-time conditions is a promising direction. Vehicle-to-Infrastructure (V2I) communication adds a centralized component to collaborative perception by integrating data from fixed infrastructure sensors, such as roadside units and traffic cameras. These sensors offer a broader sensing range and a more holistic perspective on the driving environment. Adapting V2V algorithms for V2I scenarios could improve performance, especially in diverse and complex urban settings. Additionally, research should explore the scalability and resilience of V2I systems, ensuring their reliability despite sensor failures or communication interruptions. Vehicle-to-Everything (V2X) communication, which combines V2V and V2I systems, introduces added complexity due to challenges such as ensuring data authenticity, establishing trustworthiness, and synchronizing information across multiple, diverse entities. To validate the feasibility of V2X systems, future research must extend beyond simulated environments and emphasize testing with real-world data. Experimentation in live traffic scenarios is essential to address these challenges, ensuring the reliability, scalability, and practical applicability of V2X communication in real-world traffic systems.
Finally, designing a robust testing platform is crucial for enabling fast prototyping and evaluation of cooperative driving algorithms. One notable solution is the open-source OpenCDA ecosystem, introduced by Xu et al. [659]. This platform integrates a range of SOTA models for collaborative detection, offering comprehensive support for testing in various cooperative scenarios. While OpenCDA provides tools for benchmarking cooperative perception and driving automation, its extensibility can further be enhanced to address emerging challenges. These include testing benchmarks for communication robustness, V2I collaboration, and diverse real-world scenarios to ensure broader applicability in autonomous driving research.
10 Conclusion
In this work, we have conducted an in-depth review of trajectory prediction approaches for autonomous driving, providing a comprehensive taxonomy to classify existing methodologies. This taxonomy distinguishes between learning-based and non-learning methods, with learning-based techniques further categorized into machine learning, deep learning, and reinforcement learning approaches. Additionally, we broadened the scope of the review to include a detailed overview of the prediction pipeline, covering input and output modalities, benchmark datasets used for evaluation, and the performance metrics critical for assessing prediction models.
Beyond the taxonomy and evaluation, this study highlights the pressing challenges and evolving trends in the field. Sections on robustness and generalizability emphasize the need for uncertainty quantification and mitigation, while collaborative perception is discussed as a promising solution for addressing occlusions, sensor noise, and restricted sensing capabilities. The review also explores planning-conditioned trajectory prediction, which incorporates motion planning principles to model rational agent behavior. Methods such as goal-anchor-based prediction play a pivotal role by estimating potential destinations and generating trajectories aligned with those goals, ensuring more interpretable and goal-driven outputs. Additionally, the integration of VLMs into prediction pipelines offers an exciting frontier, providing common-sense knowledge and contextual understanding to enhance decision-making in uncertain conditions. However, challenges like latency and scalability of these models require further investigation. Furthermore, the study underscores the importance of designing systems that generalize across diverse datasets and scenarios, addressing the existing gap in cross-dataset evaluations and qualitative assessments.
We have addressed the posed research questions, providing a comprehensive summary of the state of the art while identifying potential future research directions. Despite significant advancements over the past two decades, trajectory prediction remains an open and challenging area of research. The field has evolved from early physics-based models to Graph Transformers. The shift toward combining domain knowledge with data-driven methods, such as physics-based hybrid models, represents a crucial step forward in balancing interpretability with predictive capacity. This review not only captures the evolution of trajectory prediction methodologies but also sheds light on key research gaps, such as robust evaluation under real-world conditions and generalization to out-of-distribution scenarios. We believe this work serves as a valuable resource for researchers and practitioners, fostering innovation and inspiring further exploration in trajectory prediction domain for autonomous driving.